A Complete Bibliography of The Bell System Technical Journal, 1980–1983, AT&T Bell Laboratories Technical Journal, 1984, and AT&T Technical Journal, 1985–1989

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

04 September 2023 Version 1.13

Title word cross-reference

16 [NMR^{+}84]. _{2} [KKLS83, SCWB84]. k [RWJ86]. L [San81a]. M [Kav84, Yue80b, Kav85a]. μ [GAB^{+}80, LTT^{+}83, SB80, TZD^{+}83, HFE85, RCS86]. μm^{2} [NMR^{+}84]. N [Sal80].

-ary [Kav84, Yue80b]. -Asymptotically [San81a]. -means [RWJ86]. -quadrature [Kav85a]. -Way [Sal80].

1 [BMD84, Dos83, Fle84, Fre82, GLM82, Jag82a, Jag85b, PR86, SJ85, WC85]. 1.2- [LTT^{+}83]. 1.3- [HFE85, TZD^{+}83]. 1.38- [SB80]. 1.5- [LTT^{+}83]. 1/1A [GLM82]. 10-Year [LBP80]. 100-GHz [Sem85]. 105D [Lie82]. 10A [ACSW82, AFS82, AFHS82, BDL82b, BDL^{+}83, BNS82, JKT82, KMM82, NS82]. 12-GHz [Aca81b]. 16-ary [Pra80]. 16-QAM [FPQ^{+}84]. 1982/83 [CCD^{+}84, HLL^{+}84]. 1A [BHS82, BKV82, CS82, GKW82, GLM82, Nus82, Wor82]. 1B [ADRW83, AM80, BW83a, BDV83, CHTW83, CS83, DVW83, DM83, GKS83, HHS83a, SC83].

2 [CR82]. 20-Mb [Wan83]. 20-Mb/s [Wan83]. 200-Hz [WCS81]. 255 [McN80].

30-MHz [WCS81]. 32-kb [BBDS86]. 32-kb/s [BBDS86]. 370 [FMM84]. 3B20 [EC83]. 3B20D [BS83a, BW83b, EC83, GCD83, HP83a, HPW83, HK83, KAM83, KM83, KBF83, MBD^{+}83, QEG83, RBH83, RW83, Sca83, TG83, YMNW83].

4-GHz [Wan83]. 4-Wire [WPG87]. 47A [Lon80]. 48-Bit [Rob82].

56-kb [DGH^{+}82]. 56-kb/s [DGH^{+}82]. 5ESS [BHK^{+}85, CCG^{+}85, CP86, DFOS85, DHK89, FGNS85, HLRW85, HPS86, MMTV84].

630 [BM87].

75 [AK88, BBB^{+}85].

800 [BBM87, CMN84]. 800-MHz [CMN83, CMN84].

9.6-kb [CCJS82]. 9.6-kb/s [CCJS82]. 90-Mb [FPQ^{+}84]. 90-Mb/s [FPQ^{+}84]. 96 [AS84a, BH84a, BBB^{+}84, Can81, COW82, LOS84, MMTV84, MSO84].

Abbreviating [SAT83]. Abbreviations [Ano81a, Ano81b, Ano82f, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e]. ability [Wil85]. above [Ber84]. Absenteeism [Var81]. absorption [SCWB84]. abstraction [Str84]. accelerated [JLN^{+}85]. Access [ANSW82, Hey82, LDGF82, Pfe80, WA88, WPG87, Yue80a, BCW86, DS84, Kam89b, Kam89a]. accuracy [You84]. Accurate [LMM81]. Achieving [Pet86, CC88]. Acoustic [Fla80, MF82, SBD82]. Acquisition [BGGW83, HLL^{+}84]. Acronyms [Ano81a, Ano81b, Ano82f, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e]. Active [AFPP82, Tow83]. activities [MM85]. Activity [Smi83]. Adaptive [BJM^{+}81, FPQ^{+}84, Feu83, GL81, GMW82, JC80, Jay81a, LM81b, Mue81a, Pha82, Pir82, SS83, SA81b, SB83, GY85, RJ84]. Added [GNS^{+}81]. Adding [IST^{+}83]. Address [Ein80, Rei81]. Addressable [Roo82]. Adjacent [Yue80b]. Adjustment [Mue81b]. Administration [BDL^{+}82a, CF88, Nea80, ST82, WH82, YMNW83, FGNS85]. Administrative [BM82b, HSS82]. admission [RS85b]. Admittance [WCS81]. ADPCM [BBDS86, Jay81a, Jay83, RJ84]. Advances [Li87]. Affecting [Rab82]. after [Log84d, Log84b, Log84c]. age [Ber84]. aging [JLN^{+}85]. Aided [MS81]. Aids [War80]. Airline [LS80c]. AKM [SS86]. Al [DJ80]. Algebraic [Lee83, Yoo85]. Algorithm [GMW82, Hen81, IG82, MR83, MS82, Mor82, Str83, Sze82, Aga89, LF84, Tur85, WRM84]. Algorithms [GM87, Hon83, Kar83, Mue81a, Mue81b, MR81, NS85, Rab82, San80, San81b, HHL89, KLSW89]. Allocation [Elk80, GL80, SA81a, CV85]. Almost [SA85]. Almost-periodic [SA85]. altering [Fos85]. Alternative [Sun83a]. Alternatives [Koo80, FW88]. America [ACL88]. American [Rei85]. Amplifier [BCHR83, HHS83b, LFG82, Tow83]. Amplifiers [SS83, Min84]. Amplitude [Aca81a, Kar83, Log83a, SW80, Kav85a]. Amplitude-Equalizing [Kar83]. amplitude-modulated [Kav85a]. Analog [CT83, CBB^{+}87, Fla80, JCMQ83, LM81b, CCD^{+}84]. Analog/Digital [LM81b]. Analysis [Aca81a, AK88, BH84b, BB87, BS83b, BS83c, CP83a, CB81, Dos85, FH88, FH80, FLGD83, GH82, Gre85, Hey82, Hey83, HTB83, Hon84, JCV80, JCV83, Kum83, Kur80, Lon80, MR83, Maz80, Mor80a, MY83, RWA81, RD82, Sal82, WM83, Yum80, BC85, Ben84, BGGN^{+}86, Fle84, Fre86, Lo87, MS84, WM88]. Analysis/simulation [Gre85]. Analytical [Ram88, Sha85]. Analyzer [Whi83b, Whi83c]. Analyzing [Gri83, RM82]. Angle [TA82, Sil84]. Announcement [ABC^{+}81, FKSW81]. Announcements [RB82]. Answers [ND81]. Antenna [But81, Sem80, Sem83a, Sem83b, Sha85, Sil84]. Antennas [Dra81a]. Anti [Mea80]. Anti-Sidetone [Mea80]. Antireflection [Cla83, Eis84]. Aperture [JC80, Kum80]. Application [ALS80, BGK^{+}86, Elk80, LRS83, MZ82, RLS83, Rub82, Str83, WC85, Yum80, CK88, Jag85a, JLR^{+}86, Leu86, Lo87, Rab84, RWT84, RSL84, Sha85]. Applications [AT&T86, Bod81, BO85, DKW88, Dut89, EH82, HTB83, Hol83, LBHB83, MBD^{+}83, MS82, NS87, Nuc81, RCS86, RS88, Var81, BLT89, Hin87, PR86, SCS85, SP89]. Applied [Han83, Koh83]. Approach [AKH87, HG81, RW81, Eck85, SRJR87, YG85, Yoo85]. approaches [MZCW84]. Approximate [Coy82, Jag82a, Kum83, Fle84, San85b]. Approximation [Fre80b, GC81, Lee82, Yue80b]. Approximations [Fre80a, Fre82, KW84, Whi84b, Whi84c, Whi84a]. AR6A [ADG^{+}83, BCHR83, BMM^{+}83, DRS83, FS83a, GMMR83, GK83, HP83b, HHS83b, HCH^{+}83, KP83, Mar83]. arbitrary [GGV84]. architectural [CCG^{+}85]. Architecture [AFS82, BKP88, BDE^{+}81, CS82, GLM82, GKS83, KAM83, Mar82, NS82, Roc86, ST82, TG83, WH82, BBD^{+}89, BBB^{+}85, BCH^{+}86, Cam87, Far89, LFLB85, NS86, PR86]. architectures [Jon88]. Area [LF82, SKP84, Ran84]. Arranged [Ben81a]. Arrays [DG80]. ARSB [GH82]. Articulatory [MF82]. artificial [BH88]. ary [Kav84, Pra80, Yue80b]. aspects [Pet84]. ASPEN [HLL^{+}84]. ASQIC [BBM87]. Assembler [Sem81]. Assemblies [BB87]. assembly [Str88]. Assignment [Ein80, FH80, MZ82, Str83, BH84b, LF84]. Assignments [SA81a]. Assistance [ALS80, RRW80]. assistant [Kow88]. Assisted [BJW88]. astigmatic [Sem85]. astronomical [Row84b]. Asymptotic [BS83c, MMR81, MM82]. Asymptotically [San81a]. AT&T [AT&T86, AAJ^{+}86, BKP88, Cal88, CHL^{+}89, CKP88, Eck86, EEF^{+}86, Far89, Fow86, God86, HHL89, NS86, SKP84, Sur86]. AT&Tt [Bul87]. Atmospheric [KFS83]. Attenuation [LBP80, CMN84]. Audio [MDSL82]. auditing [KC88]. Authors [Ano80k, Ano80l, Ano80m, Ano80n, Ano80o, Ano80p, Ano80q, Ano80r, Ano80s, Ano80t, Ano81o, Ano81p, Ano81q, Ano81r, Ano81s, Ano81t, Ano81u, Ano81v, Ano81w, Ano81x, Ano82v, Ano82w, Ano82x, Ano82y, Ano82z, Ano82-27, Ano82-28, Ano82-29, Ano82-30, Ano82-31, Ano83s, Ano83t, Ano83u, Ano83v, Ano83w, Ano83x, Ano83y, Ano83z, Ano83-27, Ano83-28]. Automated [BR82a, BBOV82, BM82b, CHW82, DRT82, EGKS81, FV82, GH82, GB82, Hol82, LZ82, Mar82, Ove82, Ray88, RD82, Rub82, SB82, Uhr82]. Automatic [ALS80, Hol83, Lec80, LRS83]. automatically [Wil85]. Automating [Pul88]. automation [Cam87, Kow88]. Avalanche [SF82]. Average [LMOS83, SSW80]. Avoidance [Alb83].

B.S.T.J.

[Aha80, Ben80, FD82, Hen81, JCMS82, JC80, Lec80, PLR82, SB80, TA82, Wit80]. balancing [HJ88]. Band [EY81, Cro81, Wer85]. Band-Limiting [EY81]. Bandpass [Tow83]. Bandwidth [CMSP81, EH82, JCMS82, Log83a, Log85, SA81b, LL88]. Bandwidth-Conserving [Log83a]. Bandwidth-error [Log85]. Bank [AFPP82, AFHK82, ANSW82, AGS82, BM82a, CL82, COW82, CGG^{+}82, DRM83, DGH^{+}82, GLMS82, LOS84, Row84a, ACL88].

Base [BDL^{+}82a, FV82, MBD^{+}83, ST82, Gla84]. Baseband [Kav85a]. Based [BR82b, DRM83, JCMQ83, Jay83, KBG^{+}88, Lec80, LS80b, RWA81, Ves88, Yav82, AO85, BTOA84, DeT84, Jay87, LL88, RS85b, Tol84]. Bases [MY83]. Basic [WPG87]. Basics [Bra88]. basilar [SA85]. Batch [Hal83, Whi83a]. be [Cox81, San81a]. Beam [Kum80, Pan89]. Behavior [Dra81b, KKLS83, Lin80, Lin81, Liu80, Nuc81]. Behavioral [Ano83n, Ano83q, Ano83o, Ano83p, CHM83, CK83, DK83, Ega83, Eig83, Fla83, Fra83, FLGD83, Gin83, Han83, Hol83, JB83, KRS83, Koh83, Loe83, Mac83, Pau83, Ros83, SAT83, Sum83]. behaviour [Ack84]. Bell [Aho82, Ano80k, Ano80l, Ano80m, Ano80n, Ano80o, Ano80p, Ano80q, Ano80r, Ano80s, Ano80t, Ano81o, Ano81p, Ano81q, Ano81r, Ano81s, Ano81t, Ano81u, Ano81v, Ano81w, Ano81x, Ano82v, Ano82w, Ano82x, Ano82y, Ano82z, Ano82-27, Ano82-28, Ano82-29, Ano82-30, Ano82-31, Ano83s, Ano83t, Ano83u, Ano83v, Ano83w, Ano83x, Ano83y, Ano83z, Ano83-27, Ano83-28, Bat84, Ber84, Gig80, Han83, Hol83, Rei85, SBD82]. BER [Gla84]. Between [CK83, Inf80]. Biased [FT83]. biases [CW85]. Biasing [SW83]. Bidirectional [MDSL82]. Billing [LW83]. Binary [WW81, Col88, Fos84, Maz85]. Biphase [LFG82]. Biquad [Tow83]. birefringence [SS84]. Birth [Nuc81]. Birth-Death [Nuc81]. Bit [GS83a, KSB88, Rob82]. Blit [Car84, Pik84]. Block [Aha80, Chu80b]. Blocking [Ben81a, Bur82, DH82, Fre80b, GGV84, KRS82, Nea80, Whi85, BH84b, Whi84a]. board [AAB^{+}86]. Bombarded [HHG83]. Bombardment [DJ80]. both [WRM84]. boundaries [Tho87]. Bounds [AS83, CMS83, Tur85]. Breakdown [Wit80]. Breaking [SV81]. Bridge [WCS81]. Brief [Han83]. Briefs [Aha80, Ben80, FD82, Hen81, JCMS82, JC80, Lec80, PLR82, SB80, TA82, Wit80]. Broadband [CE83, DL84, HB85]. Broadcast [EHS83]. Broadcast-Quality [EHS83]. Bubble [NWB^{+}80, RB82, Wyn81, NMR^{+}84]. buckling [Len85]. buffer [Ack84]. Build [EP84]. Building [And88, Gri83, CK88]. Buildings [CMN83, Wal83]. Bulk [CEGM89]. bumps [Ray88]. Bundle [Gri82]. Bureau [BR82a, BBOV82, BM82b, CHW82, DRT82, FV82, GH82, GB82, Hol82, LZ82, Mar82, Ove82, RD82, Rub82, SB82, Uhr82]. Buried [FLRT81, Ung80]. Bursts [JS82]. Bursty [BS83c]. Bus [SA81a]. Bus-Allocation [SA81a]. buses [DeT84]. business [GE87, Rei85]. Busy [ARW80, Yum80]. Busy-Period [Yum80]. BX.25 [Kur85].

C

[Car84, CDK88, FF88, Rei81, Ros84b, Str84]. Cabinets [Coy82]. Cable [BM82b, FLRT81, GPP87, Gri82, Noe81, Ung80, CC85, HFE85]. Cables [Chu83]. Calculation [Blu80b, Len83, PCPF81, KD84]. calculations [BH84c, Jag84, Min84]. California [ACL88]. Call [BS88, Liu80, BBM87]. Calling [BBP^{+}82, CRT82, EY82, Wor82]. Calls [Inf80, Tor81]. Canceler [DC80, Fan81, Wer85]. Cancelers [Sal83, CW85]. Cancellation [GL81, Gre85, Kav85a, KS85]. Capabilities [GLM82, FGNS85, HLRW85]. Capability [FKSW81, HSS82, SW82]. Capacitated [MS82]. Capacitor [LFG82]. Capacity [Aca80, CS83, Dat80, HG81, Nuc81, Rai81, Wit84]. Card [BBP^{+}82, CRT82, EY82]. Carolina [Car80]. Carrier [BKL^{+}82, DR81, GK83, Hey82, KLRS81, AS84a, BH84a, BBB^{+}84, LOS84, MMTV84, MSO84]. Carrier-Sense [Hey82]. Carrying [Rai81].

Case [Wit81, Pha86, Row84b]. Casper [BS88]. Cassegrainian [Dra81a]. Cathodoluminescence [CTM81]. CATV [HB85]. Caused [CC80, Oga81, Gla84]. CCIS [DFS82, FM82]. CCITT [Gad81, GN83, JN81]. CD [DeT84, Max82]. Cell [Sun83a, NMR^{+}84]. Cells [Blu80b]. Cellular [Sun83b, Wal83]. Center [BDV83, GM87, GGV84, SMNV89]. Center-Tap [GM87]. Centering [Eic80]. Central [GFGB83, RBH83, Ros83, War80]. certifiable [Sha88]. certification [Eck85]. CEST [FF88]. Chains [Maz82, Jua85]. Change [Aha82a]. Changes [CB83]. Channel [AFPP82, AFHK82, ANSW82, AGS82, BM82a, CDH81, CL82, COW82, CGG^{+}82, DH82, DGH^{+}82, Fan81, GLMS82, GS83a, GC80, Jay81b, MDSL82, OOB83, Rum82, Wit81, Kar86, KM85, Log85, LOS84, OW84, Row84a, Tay84, Wer85]. channel-bank [Row84a]. Channels [FS83b, Jay81b, Oza80, AS84b, Kav84, Kav85a, KS85, RWRH85, Sal85b, SSW84, SWS84a, SWS84b]. Characteristics [Ano83n, BDL^{+}82a, BBP^{+}82, But81, CMLF80, Inf80, RCS86, Tay84]. Characterization [CDH81, DR81, EGKS81, Hef80, KLRS81, Pre81, CCD^{+}84]. Characterize [Chu83]. Cheap [Wei84]. Check [Fan81]. Chip [DC80, WPG87, AO85, GOP87, Rai84]. Chirp [EH81]. Chromatic [LTT^{+}83]. CIMA [BCH^{+}86]. Circuit [AFS82, AM80, CB83, NS82, PKSG83, Smi83, WA88, BN84, Kum88]. Circuits [Aha82a, Alb83, BBF81b, GLMS82, Wes81, AB89, Col88, MZCW84]. Circulators [Kne80]. Clad [JCMS82]. Cladding [GNS^{+}81, AG84]. Class [Aha82a, Fre82, Hef80, Hef82, MMR81, Mea80, Sal80]. Cleveland [Car80]. Click [Log84a]. clipping [Log84d, Log84b, Log84c]. Clock [MRW81]. clocked [Ack85, Dos85]. Closed [MMR81, MM82, Whi84d]. coated [KD84, Len85]. Coating [SF81]. Coatings [Cla83, Eic80, Eis84]. Code [AFPP82, BJM^{+}81, CB83, GS83b, Joh81, Rei81, SV81, CS85]. Code-Breaking [SV81]. Coded [Ein80, NB82]. Coder [CCJS82, Pir82, Jua84b]. Coders [Cox81, RN82]. Codes [Aha80, CMS83, JN81]. Coding [Aha80, BJM^{+}81, Cro81, GN83, GS83a, JC80, Jay83, JSC83, NR80, Oza80, SNR80, WW81, WWZ80, WWKZ84, BBDS86, Gha84, Kav85b, KM85]. coefficient [KD84]. Coefficients [RSL83]. Coherent [Sal85a, Sun82, Maz85]. Collapsing [CB81]. Collision [Alb83]. Collision-Avoidance [Alb83]. Color [Has81]. Colorization [PC82]. combinatorial [LF84]. Combined [GS83a, MF82]. Combining [Kar83, RSL84, YG85]. Come [SA81a]. Command [RW84, SAT83]. commercial [BN88]. Commitment [Dat80]. Common [Fan81]. Common-Channel [Fan81]. Communication [Aca80, GAB^{+}80, HTB83, MDSL82, Str83, KM85, Maz85]. Communications [FS83b, LS80b, LF82, MBJ81, SF82, BBB^{+}85, Col88, Kar86, Kav85b, Sal85a]. Companded [EH82, SSW84, SWS84b]. Companding [Log83b]. Company [Pfe83]. Comparative [MR81, Pre81]. Comparing [Whi83a]. Comparison [Cox81, DJ80, LM81a, Pau83, PCP80, Sch83b, DeT84]. Comparisons [DH82, Min84]. Compatible [GN83, Rze83]. Compensated [NR80, RN82]. Compensation [SNR80]. Compiling [Rei81]. Completion [Liu80]. Complex [BB87, Dra81a, NS82]. Complexity [Ben83, KM85]. component [Dec88]. Components [BR82a, Lin81, SY82]. Compound [DSMBB89, CEGM89]. Compressed [EY81]. Compression [EY81, EH81, EH82, EH83b, EHS83]. Computation [Sal80, Ack84, Ack85].

Computational [Gri83]. Computer [But84, Fre80a, Gri83, LMP82, PW88, SS86, WCS81, Yav82, RS85c]. Computer-Based [Yav82]. Computer-Operated [WCS81]. computers [BGS^{+}89, FHR84]. Computing [Mey82, Rai84, Gav85]. COMSTAR [TA82]. Concentration [WW80]. Concentrator [Can81]. Concentricity [SF81]. concept [ABC89, Leu86]. concepts [Hei89]. Conditional [Gha84, SJ85]. Conducting [ABS82]. Conductivity [HM80]. Conductors [Rai81]. cones [SS86]. Configuration [CHTW83]. Configurations [Sun83a]. configuring [Gav85]. Confinement [PPC81a, PLR82]. Conformal [Dra81a]. Congestion [DN88, Fre80b]. connect [MS84]. Connected [MR81, MRR81, RBW82, Rab84, RWT84, RWJ86]. Connected-Digit [RBW82]. Connected-Word [MR81]. Connecting [Ben81a, Ben83]. Connection [CCD^{+}84, HLL^{+}84]. connectionist [Jon88]. Connections [CHN80]. Connector [Fla83, AFM87, BMD84]. Conserving [Log83a]. Consideration [CMSP81]. Considerations [Oga82, Tow83, Yav82]. Constrained [AS83]. constraints [GP84, KW84, WRM84]. construction [EP84, Lin84]. constructive [LF84]. Consumption [Gri83]. Contained [Jay81b]. containment [GP84]. Content [Roo82]. Content-Addressable [Roo82]. Context [Hol82]. Context-Sensitive [Hol82]. Continuing [Lie82]. Continuity [Fan81]. Continuous [Kur80, Pau83, Sun82, Sun83b, RJLS85]. Continuous-Groove [Pau83]. Continuous-Phase [Sun82]. continuum [WC85]. contours [Rab84, RPS84]. Contrasting [Fos84]. Contributors [Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano81n, Ano82g, Ano82h, Ano82i, Ano82j, Ano82k, Ano82l, Ano82m, Ano82n, Ano82o, Ano82p, Ano82q, Ano82r, Ano82s, Ano82t, Ano83f, Ano83g, Ano83h, Ano83i, Ano83j, Ano83k, Ano83l]. Control [AKK81, BDV83, DHM^{+}86, GFGB83, Gri83, JCV83, Kar83, KP83, NS82, PKSG83, QEG83, Ros83, SF81, BBB^{+}85, Cam87, DN88, LL88]. Control-Complex [NS82]. Controlled [AM82, BDL^{+}82a, BBP^{+}82, CRT82, DFS82, EY82, FM82, GLM82, HSS82, HMSY82, LLS82, RWR80, ST82, SW82, Bro85]. Controller [Kar80b, Lec80]. Convergence [Maz80, Mue81b]. Conversant [PR86]. Conversational [LS80c]. Conversational-Mode [LS80c]. Conversion [Fla80]. convexity [Jag85b]. Convolutional [Kav85b]. Coordinates [Dra81a]. Copper [HM80]. Core [GNS^{+}81, Ung80]. Core-Shield [Ung80]. Cores [Phi82]. Corporate [DKW88]. Correction [GS83b, LS80b]. corrugated [DL84]. Cost [MS82, She88, KS86]. counting [Wei84]. Coupled [Tow83, Sal85b]. Coupled-Single-Amplifier-Biquad [Tow83]. couplers [LG84]. Covariance [Hef80, Hon83]. Cover [Sem80]. COZY [WCS81]. Craft [BS83a]. Criteria [San81a, San85a, SF80]. Critical [GL80]. Cross [KS85, Gre85, Kav84, Kav85a, MS84, Sal85b]. cross-connect [MS84]. cross-coupled [Sal85b]. Cross-polarization [KS85, Gre85, Kav85a]. cross-polarized [Kav84]. crosscorrelation [Ben84]. Crosstalk [Lin80, Lin81]. crypt [RW84]. Cryptographic [Hen81]. Crystal [JCV83, CEGM89]. Crystals [JCV80]. CSMA [DeT84, Max82]. CSMA/CD [DeT84, Max82]. Cubic [Gra83]. Current [Blu80b, Dix80, Rai81]. Current-Carrying [Rai81]. Custom [GLMS82, Wor82]. Custom-Integrated [GLMS82]. Customer [CM80, Fre80a, Hal83, Liu80, She88, Whi83a, WR83, Wil85]. Customer-Viewed [Fre80a]. Cutoff [Tor81]. Cyclic [Aha80, WC85]. Cycling [GNS^{+}81]. cyclostationary [Ack84]. cyclotomic [HKMP85].

D [SCWB84]. D4 [AFPP82, AFHK82, ANSW82, AGS82, BM82a, CL82, COW82, CGG^{+}82, DGH^{+}82, GLMS82]. Data [AE83, AMS82, AGS82, BO83, BDL^{+}82a, BPRS83, BP83, BC83, BKP88, BGGW83, DKZ80, DKW88, DGH^{+}82, EH83a, FV82, FGP^{+}83, GL81, GS88, GK88, GFGB83, HKRS83, HM83, Hef80, Hey83, IST^{+}83, KBN88, LDGF82, LM81b, Max82, MY83, MS81, Pfe83, ST82, Sch83a, SV81, SKP84, Str84, Wel87, WSX82, Yum80, ACL88, AS84b, BC85, BBM87, But84, CCD^{+}84, CW85, DS84, Fos85, FM84, HB85, HLL^{+}84, Hon84, Kum88, ML84, Ran84, Row84a, Row84b, Sha88, WM88, Wer85, Yan85]. Data-Aided [MS81]. data-driven [CW85, Wer85]. data-flow [Sha88]. Data-Handling [AMS82]. Data-Transport [Hey83]. Database [Aho82, BBN82, CR82, Coh82, Gog82, Gol82, LDGF82, Lin82, Roc82, Roo82, WH82, Wei82, CDG^{+}89]. Databases [Lee83, Wei82]. Dataport [ANSW82, AGS82, DGH^{+}82]. Datastream [Swa85]. Day [KRS82]. Day-to-Day [KRS82]. DBM [CR82]. DBM-2 [CR82]. Death [Nuc81]. Debugging [Car84]. Decision [Maz80, SF80, SWS84a]. Decision-Directed [Maz80]. Decoding [Tim80, Tim81]. decomposition [WC85]. Decreases [Bur82]. Decryption [Hen81]. Deep [DJ80]. Definition [Rze83]. Degradation [Wit80]. degradations [Gla84]. degrees [Ber84]. Delay [Aha82a, AS83]. Delays [Fre80a, Hal83, Whi83a, Ack85]. Delivery [GB82]. Delta [SA81b, Ste82]. Demand [Inf80, IG82, Sze80, Sze82, BH84b, You84].

Demands [Ano83q]. demodulation [SWS84a]. Density [Fla83, Son83]. Dependent [WW80]. Depolarization [Chu80a]. depressed [AG84]. Derivative [JCV83]. Description [BDL^{+}82a, BBP^{+}82, DM83, LF82, SBD82]. Descriptions [WW81, WWZ80]. Design [AFS82, Aha82b, ACL88, AFHK82, AG84, Ano83o, ACM81, BC85, CR82, CS82, DRT82, Dec88, GMMR83, GH82, Gog82, HP83b, HPU81, JKT82, LS80a, LZ82, Lin82, Mac83, NWB^{+}80, NS82, PKSG83, Pha86, PK84, ST82, Sem81, Sha88, Tow83, WH82, WA88, Wes81, Wil82, BH84a, BGS^{+}89, Eis84, FT87, KS86, Kow88, LFLB85, MZCW84, SMNV89]. designed [Log84d, Log84b, Log84c]. Designer [BF80, MT80, O’N80, Tho80]. Designing [KRS83, Aga89]. Designs [Chu80b]. Detecting [Ben81b, Fos85]. Detection [EH81, Feu81, Gad81, HPW83, JS82, Pra80, Maz85, WRM84]. Detector [Wyn81]. Determination [CC85, HFE85, SA85]. Deuterium [SB80]. Deuterium-Hydrogen [SB80]. Develoment [GB82]. Developing [War80]. Development [And83, BGK^{+}86, Dix80, Geh81, HHS83a, KC87, KR80, RW83, BHK^{+}85, BBW88, KPW85, Pet86, RS86]. Device [BJW88, Bod81, Dix80, Dut89, NWB^{+}80, DD87, MS89, NMR^{+}84, SP89, Sha88]. Devices [DSMBB89, DJ80, JDW89, LMM81, AB89, JLN^{+}85, KD84, Wit84]. Diagnostic [QEG83]. Diagnostics [PC82]. diagrams [AG84]. Dialer [RWR80]. Dialing [Liu80]. Diameter [GNS^{+}81, JCV83]. Diffeomorphisms [San80, San81b]. Difference [Sch83b]. Differences [Ega83]. Differential [BJM^{+}81, GS83b, GH81, Wyn81, Maz85]. Differentials [EH82]. Difficulty [CK83]. Diffusion [Phi82, WW80]. Digit [RBW82]. Digital [Aag81, AFPP82, Aha82b, Alb82, AFHK82, Ang81, ANSW82, AGS82, BBF^{+}81a, BM82a, BBF81b, Bod81, BDE^{+}81, BJM^{+}81, BST81, BGK^{+}86, BKL^{+}82, BKO81, Can81, CC80, CL82, COW82, Cox81, Cro81, CGG^{+}82, DFN81, DGH^{+}82, EG81, Els80, Fav81, FYC81, Fla80, FS83b, Gad81, GW81, GAB^{+}80, GLMS82, HG81, HMPV81, Kar80b, LM81b, MBJ81, McN80, SS83, Sal83, Sal85b, Sem81, Smi82, Sun83a, Sun83b, Wan83, AS84b, BH84c, Cho86, FPQ^{+}84, Fos85, Gla84, Gre85, GY85, KS85, YG85]. Digital-to-Analog [Fla80]. Digitally [GMW82, CW85]. Digitizers [PS81a]. Digits [WR83]. Dijkstra [Sha81]. Dimension [KSM81]. Dimensional [But81, WW80, CS85, NS85, Row84b]. Dimensioning [Els80]. Diodes [KBZ80]. Direct [Fla80, FM82, Liu80]. Direct-Signaling [FM82]. Directed [BK83, Gol82, Maz80, Str88]. Direction [San80, San81b]. directional [LG84]. Directions [Dix80, GS88, AR86, BBW88]. Directory [ALS80, RRW80, RWT84]. Discipline [Dos83, FS88, Fle84]. disciplines [FM84]. Discrete [FH80]. Discrete-Event [FH80]. Dislocation [JCV80]. Dispersion [LTT^{+}83, Nas81, PLR82, RCS86]. Dispersion-Shifted [RCS86]. Dispersionless [PPC81a, PPC81b]. Displacement [SNR80]. Displays [MF82]. Distance [ALS80, CMS83, Liu80, RS85a]. Distortion [Ben84, Wit81]. Distributed [Coh82, LMP82, LMP83, Lin82, Rub82, DS84, Gav85, HJ88, Leu86]. Distribution [Fre82, Mey82, PCP80, SY82, SF80, Kli85, RS85c]. Distributions [Nea80, Whi84b, Whi84c]. Dithered [JN81, Joh81]. Diversity [Kar83, Rum82, Sun82, Gla84, GY85, KM85, YG85]. Dividers [Sal80]. DMERT [BS83a, BW83b, EC83, GKS83, GCD83, HP83a, HPW83, HK83, KAM83, KM83, KBF83, MBD^{+}83, QEG83, RBH83, RW83, Sca83, TG83, YMNW83]. Document [Ker89]. Documentation [GB82, HPU81]. Documents [CK83]. Domain [Rum80]. Dominant [Lin81]. Doped [PCPF81]. Doppler [LS80b]. Doppler/Fading [LS80b]. Double [CMLF80, DJ80, JCMS82]. Double-Clad [JCMS82]. Double-Heterostructure [DJ80]. Double-Mode [CMLF80]. Downlinks [Aca81b]. DPCM [Jay81b, NB82, Pir82]. DPCM-Coded [NB82]. driven [CW85, Wer85]. DSP [CCJS82]. DSP32 [BGK^{+}86]. DTWP [BTOA84]. Dual [AFPP82, Toy87, Gre85, Len85]. dual-coated [Len85]. dual-polarization [Gre85]. dually [AS84b, KS85]. Duplex [CT83]. During [JS82, But84]. Dynamic [ACM81, AKK81, HJ88, MR81, MRR81, Rab82, BTOA84, Jua84a, Wei84].

EADAS [BGGW83]. Earth

[Chu80a, LBP80]. Earth-Satellite [LBP80]. Echo [CHN80, DC80, Fan81, Sal83, CW85, Wer85]. Economic [Koo80, Ove82]. Editing [Max80]. Editor [Ano82u, Ano83r, Hol87]. Editors [BK83]. Effect [Fan81, GNS^{+}81, LL88, SWS84a]. Effective [Jay87, RS86, KS86]. Effects [CHN80, CW85, Con80, DRM83, EY81, Hea81, KRS82, Koh83, Kum80, RWA81, TA82, Ung80, Wer85]. Efficiency [Pra80, SW81]. Efficient [Blu80b, CGH81]. eight [CS85]. eight-state [CS85]. EISPACK [Len83]. EL2 [KR80]. Electret [KR80]. Electrical [Pet84]. Electromagnetic [Ber82, Blu80a, PPC81a]. Electromechanical [BB87]. Electronic [BB87, DSMBB89, HJG88, Wil82, MS89, SP89]. Electronics [CD80, Coy82]. Electroplating [Blu80b]. Element [Kne80, Len83, CC88]. Elementary [Row82, Pet84]. Elements [JB83, NWB^{+}80, Smi82]. Elevation [TA82].

Eliminating [Sem80]. Embedded [GS83b, RWT84]. emergence [BH88]. Emitting [KBZ80]. Empirical [CP83a]. Employing [NWB^{+}80]. End [CHW82, Yav82, CCD^{+}84, HLL^{+}84]. Energy [BR82b, Gri83, PPC81a, Rab84, RPS84, RSL84]. engineered [Aki84]. Engineering [BKV82, DK83, HM83, Yav82]. Enhancement [RJ84]. entry [But84]. Envelope [EH81]. Environment [BPRS83, BF80, DKW88, MT80, PK84, Tho80, Wal83, Yue80a, BHK^{+}85, DS84, KPW85]. Epitaxy [Pan89]. Equalization [BMM^{+}83, GW81, Kas82, LM81b, MS81, AS84b, GY85, KS85]. equalizaton [FPQ^{+}84]. Equalizer [GW81, GMW82, Maz80, Mue81b]. Equalizers [Mue81a]. Equalizing [Fos85, Kar83]. equation [DS85]. Equations [Blu80a, GH81, PCPB80]. Equilibrium [Kum83]. Equipment [DRS83, FS83a, FGP^{+}83, GFGB83, Tor81, War80]. Equivalent [Kum83]. Error [GS83b, Hea81, JS82, Sun82, Log85]. Errors [GS83b, SSW84, SWS84a]. especially [WC85]. ESS [ABC^{+}81, BBN82, BDL^{+}83, CH81, DG81, FKSW81, GLM82, GY83, HMPV81, IST^{+}83, LW83, Mar81, MRW81]. Estimates [Reu83]. Estimation [Aca81a, Hea81, LSY88, Shi82, Jua85, NS85]. Estimations [Toy87]. Estimator [Ben80]. Evaluating [KRS83]. Evaluation [CP83a, CS83, KSB88, KK82, Koo80, Ove82, Pha82, RD82, CLMS87, Jua84b]. Event [Ben81b, FH80]. Evolution [BBOV82, CH81, Fed84, GY83, Lie82, Rit84a, Rog87, Ros84b, Ald86, God86]. Evolving [CDK88, CF86]. Exact [KD84, PCPB80]. Example [Sha81]. Examples [GH82]. excellence [CC88]. Exchange [SB80, Log85]. Exhibiting [PLR82]. existence [San85a]. Existing [IST^{+}83]. exit [SS86]. Expansion [Nuc81]. Expansions [MMR81, MM82, San82a, San82b, San85a].

Experience [DG81]. experiences

[BHK^{+}84]. Experimental

[BT83, CE83, HAK83, JCMS82, KFS83, Max80, NMR^{+}84, PKSG83, Sil84, Wan83]. Experiments [CHM83, LBP80]. Expert [Cal88, Mar88, FF88, Kow88]. Explicit [Jay83]. Exponential [Feu83, Whi84c]. Extending [Tho87]. Extension [JN81]. Extensions [LM81a]. Extremal [Maz82, Whi84b].

Fabricated [DJ80]. Fabrication [DSMBB89, PLR82, PKSG83]. facets [Eis84]. Facilities [Bod81, Str83, Whi85]. Facility [Max80, RWRH85]. Facsimile [JN81]. Factors [Ano83n, Ano83q, Ano83o, Ano83p, CHM83, CK83, DK83, Ega83, EY82, Eig83, Fla83, Fra83, FLGD83, Gin83, Han83, Hol83, JB83, KRS83, Koh83, LDGF82, Loe83, Mac83, Pau83, Rab82, Ros83, SAT83, Sum83]. Factory [DTV85, Cam87, Pul88]. Fading [BMM^{+}83, FS83b, GC80, GC81, LS80b, Rum80, Rum82, Sun82, Vig81, Yue80a, AS84b, Kav84, Kav85a, KS85, Log85, RWRH85, SSW84, SWS84a, SWS84b]. Fail [Alb82]. Fail-Safe [Alb82]. Fair [Hen84]. Family [AFPP82, AFHK82, ANSW82, AGS82, BM82a, CL82, COW82, CGG^{+}82, DGH^{+}82, GLMS82, BMD84]. Fasnet [Hey83, LF82, ML84]. Fast [Hen81, KBZ80, MR83, Shi82, And88]. faster [Fos84]. Fault [CB81, EG81, HPW83, LM81a, LMP83, Wil82]. Fault-Collapsing [CB81]. Fault-Simulation [LM81a]. Fault-Tolerant [Wil82]. Feature [ARW80, KSM81, AO85, DHK89]. Features [BK83]. Feed [CE83, DL84]. Feeder [BKL^{+}82, Dat80, Elk80, Koo80, Noe81]. Feeder-Cable [Noe81]. Females [Koh83]. Fiber [BS83b, CP83a, CP83b, CMSP81, Eic80, GNS^{+}81, JPW87, JCMS82, KC87, Kas82, Len83, Oga82, Pau83, PLR82, Phi82, SF81, BBB^{+}84, CC85, Len85, TNG88].

Fiber-Coating [SF81]. Fiber-Optic [BS83b, Pau83]. Fibers [CMLF80, LTT^{+}83, PS81a, Pre81, SB80, AG84, SS84, SCWB84]. Field [ABS82, Eig83, EC83, Gin83, KLRS81, PPC81a, Pau83, YMNW83]. Field-Measurement [KLRS81]. Field-Repair [Pau83]. Field-Tracking [ABS82]. Fields [PPC81a]. File [HP83a, RW84, Roc82, ASG85]. files [Swa85]. Film [AFPP82]. Filter [AFPP82, DRM83, EH81, PW82]. Filter-Bank-Based [DRM83]. Filtering [Ang81, DG80, Jay81a, Pha82, Row84b]. Filters [LFG82, LS80a, Tow83, BH84c, HKMP85]. Finding [MS82]. Fine [And83, Rai81]. Fine-Line [And83, Rai81]. Finite [Dra81b, LS80a, Len83, Ack84]. FIR [LS80a, MR83]. Firestopped [Gri82]. Firmware [AFHS82, Bre81]. First [SA81a, Son83]. First-Come-First-Serve [SA81a]. First-Order [Son83]. Fixed [Bur82, Hon83]. Fixed-Order [Hon83]. Flexible [LFHV88]. Flow [Blu80b, CGH81, PPC81a, LL88, Sha88]. Flows [MS82]. FM [Ben81b]. Focal [DG80]. force [Str88]. force-directed [Str88]. Forces [Koh83]. Forecast [Noe81]. Forecasting [Feu83, Mor82, PW82, Sze82, You84]. Foreword [Aho84]. Fork [BS88]. form [Yoo85]. Formulae [Hef82]. Formulas [Ste82]. Forward [Eic80, GS83b]. Foundations [HM83]. Four [CS85]. Four-dimensional [CS85]. Fourier [Log84c]. Fractional [Aha82a]. Fractionally [GW81, GMW82]. Fractionally-Spaced [GW81]. frame [Col88, DN88, Kum88, RS85a]. frame-relay [DN88, Kum88]. framing [FM84]. Frequencies [HM80]. Frequency [Aca80, Ben80, BBF81b, Dra81b, Ein80, Feu81, GHP80, HHG83, Has81, JCMQ83, KP83, Lec80, MF81, PS82, Rum80, Sal80, SW80, Sch83b, Tim80, Tim81, Tim82, Yue80a]. Frequency-Domain [Rum80]. Frequency-Hopped [GHP80, PS82, Tim80, Tim81, Tim82]. Frequency-Hopping [Yue80a]. Frequency-Reuse [Aca80]. Frequency-Shift [Ben80]. Front [CHW82]. Front-End [CHW82]. FSK [GHP80, Tim80, Tim81, Wan83]. Full [CT83]. Fully [Kar80b]. Function [Hef80, SY82, FF88]. functional [San85a]. Functions [Ano83p, HSS82, Lee82, LRS83]. Fundamental [HHL^{+}87, JB83]. future [Hei89, Ros84b].

G

[Dos83, Fle84, Fre82, Jag82a, Jag85b, WC85]. Ga [DJ80]. GaAs [Dix80, JCV80, JCV83, Oga81]. Gain [JCV83, SNR80]. Gate [RS88]. Gates [LMM81]. Gauss [LMOS83]. Gaussian [Mey82, Row82, SSW84, Ste82, Sun82, SWS84b]. GE [Rei85]. General [Aha82a, AAB^{+}86]. general-purpose [AAB^{+}86]. Generalizable [IABB^{+}89]. Generalizations [MMR81, Log84d]. Generation [BK83, Fav81, JCV80, Rub82]. Generator [Rob82]. generators [CK88]. Generic [LLS82, MZCW84]. Germania [PCPF81]. Germania-Doped [PCPF81]. GHz [Aca81b, Sem85, Wan83]. GI [Fre82, Jag82a]. GI/G/1 [Fre82, Jag82a]. gigabit [Rog87]. global [San85a]. goal [LF84]. goal-programming [LF84]. Good [Rob82]. government [BN88]. Graded [PPC81a]. Graded-Index [PPC81a]. Gradient [Feu83]. Gradients [Sch81]. Graph [MZ82]. Graphic [MF82]. Graphics [BM87, Pik84]. Graphs [CDH81, CHM83, DH82]. Grating [DG80]. gray [Gha84]. gray-level [Gha84]. Gregorian [Dra81a]. GRIN [LG84]. GRIN-lens [LG84]. gripper [Bro85]. Groove [Pau83]. Group [KRS82]. Groups [Els80]. Growth [Ben83, JDW89, KKLS83, CEGM89].

H [KKLS83, SCWB84]. Half [CT83, Jay81b]. Half-Rate [Jay81b]. Hand [Koh83]. Hand-Turning [Koh83]. Handling [AMS82]. Hardware [CHTW83, CRT82, HAK83, SKP84]. Haul [DKZ80, GPP87]. Heat [Gri82]. Heavy [Whi84a]. Heavy-traffic [Whi84a]. HEQS [DS85]. Hermitian [Ber82]. Heterojunction [KBZ80]. Heterostructure [DJ80]. Hidden [RLS83, Jua84a, RLS84, RJLS85]. hierarchical [Aki84, DS85, KV84]. High [DKZ80, Dra81b, Fla83, GAB^{+}80, HHG83, HG81, LBHB83, RS88, Rze83, SF81, Dos85, Kav85b, SS84]. high-birefringence [SS84]. High-Capacity [HG81]. High-Definition [Rze83]. High-Density [Fla83]. High-Frequency [Dra81b, HHG83]. High-Power [LBHB83]. High-Quality [RS88]. High-Speed [DKZ80, GAB^{+}80, SF81, Kav85b]. History [Ano83m, Han83, Nus82, God86]. Homenet [HB85]. Homogeneous [Ber82]. Hopped [GHP80, PS82, Tim80, Tim81, Tim82]. Hopping [Yue80a]. Horn [But81, Sem80, Sem83b, Sha85, Sil84]. Horn-Reflector [But81, Sem80, Sem83b, Sha85, Sil84]. Host [BDL82b]. Hosting [BDL^{+}83]. Household [Inf80]. houses [CMN84]. Human [Ano83n, Ano83q, Ano83o, Ano83p, CHM83, CK83, DK83, Ega83, EY82, Eig83, Fla83, Fra83, FLGD83, GH82, Gin83, Han83, Hol83, JB83, KRS83, Koh83, LDGF82, LZ82, Loe83, Mac83, Pau83, Pfe80, Ros83, SAT83, Sum83, Yav82]. Hybrid [Dos83, Sem83a]. Hybrid-Mode [Sem83a]. Hydrogen [SB80]. Hydrothermal [KKLS83]. Hyperbolic [PCP80]. Hypergraph [Gol82]. Hz [WCS81].

IC [BGGN^{+}86]. ICS [HHL^{+}87]. Identification [MR83]. II [Chu80b, GH81, JC80, KW84, Log84c, OW84, WW81]. III [AB89, CEGM89, DSMBB89, Dut89, JDW89, Log84d, SP89, Whi84c]. Imaging [CE83, DG80]. Impact [Noe81]. Impacts [She88]. Impairment [LS80b]. impairments [Wer85]. Impedance [HHG83, WCS81, Bro85]. Impedance/Admittance [WCS81]. Impedances [Dra81b]. Implanted [NWB^{+}80, NMR^{+}84]. Implementation [BBF^{+}81a, CR82, Coh82, CRT82, HSS82, Kar80b, AO85, FMM84, KV84, Leu86, SKP84]. Implemented [Cox81, GMW82, CW85]. Implementing [Rob82]. Implications [JCV83]. Improved [GW81, NB82, RBW82, Tim80, TRW82, WRM84]. Improvement [Aca81b, AH86, FHH^{+}87]. Improving [FS88, SSR81, KS86]. impulse [Ben84]. in-band [Wer85]. In-process [Fow86]. incorporating [WRM84]. Independent [Log83a, RLS83, RRW80, RLS84]. Index [PPC81a, PPC81b, PCP80, WM83]. indoor [KM85]. Induced [BS83b]. inductive [Rai84]. industrial [Rei85]. inequalities [Jag85a]. infinite [Log84b]. Information [ALS88, Feu81, FLGD83, LS80c, MF82, Ber84, Hei89, NS86]. Infrared [Nas81]. InGaAs [SCS85]. InGaAsP [LTT^{+}83, TZD^{+}83]. Initial [Mue81b]. Injection [HHG83, LTT^{+}83]. Input [BW83b, Kur80, LS80c, Sal82, Ste82, Rit84b, San85a, San85b]. input-output [Rit84b, San85b]. Input/Output [BW83b, San85a]. Inputs [Row82]. INS [Man89]. Insensitive [LFG82]. insertion [Dec88]. insertions [Str88]. inspection [Ray88, dlRR86]. inspections [Fow86]. instruction [Wei84]. Instructions [KRS83]. Integral [Blu80a, MMR81, MM82]. Integrated [GLMS82, ML84, Max82, PKSG83, Wes81, AB89, BCH^{+}86].

Integration [ADRW83, BBF^{+}81a, GH81, KM83, MMTV84]. intelligence [BH88]. Intelligent [Roo82, And88]. Interactive [Wes81]. Interconnection [PK87, CLMS87, HHL^{+}87, HHC87, Man89]. Interface [Ano83o, BS83a, BDV83, HMPV81, Pfe80, BGS^{+}89, Far89, GOP87, IABB^{+}89]. Interfaces [Wel87]. Interference [CC80, GL81, Yue80b, Kav85a]. Interframe [Pir82, SNR80, MS84]. International [CDG^{+}89]. Interoffice [Fan81]. Interpolation [SB83]. Interpretation [KBG^{+}88]. Intersymbol [CC80, GL81]. Interval [Con80]. Intra [Pir82]. Intra-Interframe [Pir82]. Introduction [Ang81, FH88, LRS83, MM85, Sch83a, Sum83]. Inventory [Bre81]. Inversion [Jag82b]. investigation [Sil84]. Ion [NWB^{+}80, NMR^{+}84]. Ion-Implanted [NWB^{+}80, NMR^{+}84]. ISDN [Ald86, CP86, DKW88, DN88, GOP87, HPS86, Kum88, NS86, Roc86, WPG87]. Isolated [BR82b, DRM83, RWA81, RW81, RRWK82, Rab82, RLS83, TRW82, WR83, BTOA84, RPS84, RLS84, RSL84]. Issue [Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano81n, Ano82g, Ano82h, Ano82i, Ano82j, Ano82k, Ano82l, Ano82m, Ano82n, Ano82o, Ano82p, Ano82q, Ano82r, Ano82s, Ano82t, Ano83f, Ano83g, Ano83h, Ano83i, Ano83j, Ano83k, Ano83l, GK88]. Issues [Lin82, ND81, WH82, HHL^{+}87]. Itakura [Jua84b].

Job [Hef82]. Josephson [FD82]. Judgements [CHM83].

Kalman [PW82, Pha82]. Karmarkar [KLSW89]. Karmarkar-type [KLSW89]. kb/s [BBDS86, CCJS82, DGH^{+}82].

Kendall [Nuc81]. Key [FLGD83, CC88]. Key-Word [FLGD83]. Keyboards [Loe83]. Keying [PS81b, Yue80a, Yue80b]. keys [Jay87]. Kinds [FT83]. Knapsack [Hen81]. Knobs [Koh83]. Knowledge [KBG^{+}88, Bra88]. Knowledge-Based [KBG^{+}88]. KORBX [CHL^{+}89, HHL89].

labeling [FW88]. Laboratories [Aho82, Ano80k, Ano80l, Ano80m, Ano80n, Ano80o, Ano80p, Ano80q, Ano80r, Ano80s, Ano80t, Ano81o, Ano81p, Ano81q, Ano81r, Ano81s, Ano81t, Ano81u, Ano81v, Ano81w, Ano81x, Ano82v, Ano82w, Ano82x, Ano82y, Ano82z, Ano82-27, Ano82-28, Ano82-29, Ano82-30, Ano82-31, Ano83s, Ano83t, Ano83u, Ano83v, Ano83w, Ano83x, Ano83y, Ano83z, Ano83-27, Ano83-28, Han83, SBD82]. laboratory [RWRH85]. Labs [Ber84]. Language [CDK88, FF88, HBH88, Swa85]. LANs [WC85]. Laplace [Jag82b, Jag85a]. Large [BBF^{+}81a, Mor80b, MW82, RRWK82, WR83, Yav82, BBW88, Gav85, Kli85, LMS^{+}89, Swa85, Wil85]. large-scale [BBW88, Gav85, Kli85]. Laser [Dix80, Eck85, Eis84, HFE85, dlRR86]. Lasers [DJ80, HHG83, LTT^{+}83, LBHB83, SW83]. Lattice [Shi82]. Launch [Kum80]. launchers [Sem85]. layer [Eis84, Kur85]. Layout [Wes81]. Leakage [GMW82]. Least [Ben80, Hon83, MR83, Mue81a, Mue81b]. Least-Squares [Ben80, Hon83, MR83, Mue81a, Mue81b]. LEDs [DKZ80, GAB^{+}80, TZD^{+}83]. length [Gha84]. lens [LG84]. Letter [Ano82u, Ano83r, RWT84]. Level [JSC83, Gha84]. Levinson [Ack84]. library [FF88]. Light [CD80, KBZ80]. Light-Emitting [KBZ80]. Lightguide [Alb82, Kum80, MDSL82, NS87, PCP80, PCPF81, CC85]. Lightguides [PPC81a, PPC81b, PCPB80, RCS86]. Lightning [Ung80]. Lightwave [AKH87, AFM87, BJW88, DD87, Dut89, GAB^{+}80, San87, Wel87, Li87, Maz85, Rog87, Sal85a, Tho87]. Likelihood [Aca81a, Jua85]. Limit [Bur82]. Limiting [EY81]. Limits [Smi82]. Line [And83, ARW80, BH86, HP83b, PKSG83, Rai81, Rum80, Sch81, Sch83b, She88, Col88, GB82]. Line-of-Sight [Rum80, Sch81]. Linear [AS84b, LS80a, Lee82, Sal85b]. Linear-Phase [LS80a]. Linearization [SS83]. Lines [Aha82b, BGGN^{+}86, Pet84]. Link [Yum80, Kur85]. Links [DKZ80, IST^{+}83, KBZ80, Wel87]. Liquid [JDW89]. Listener [CHN80]. listing [RWT84]. Lithographic [And83]. Load [Bur82, Con80, KRS82, Mor82, HJ88]. loading [LMS^{+}89]. Lobes [DG80]. Local [Gol82, Inf80, LF82, Max82, McN80, Ran84, DS84, Lav87, SKP84]. Local-Area [LF82]. Localization [FLRT81, Log84b]. Locating [Fla83]. Location [AE83, HKMP85, Kli85]. Log [SY82]. Log-Normal [SY82]. logarithmic [SWS84a]. Logic [CB81, EG81, LMM81, LMP82, Smi82]. Logical [Coh82]. Long [BW83a, CMSP81, GPP87, JS82, OOB83, SF82]. Long-Haul [GPP87]. Long-Range [BW83a]. Long-Wavelength [OOB83, SF82]. Loop [BKL^{+}82, CM80, Car80, CD80, DRT82, Dat80, DK83, Elk80, FH80, GPP87, GL80, Gol82, Hol82, Koo80, NS87, Rub82, Uhr82, AS84a, BH84a, Bat84, BBB^{+}84, LOS84, MMTV84, MSO84, TNG88]. Loops [WPG87]. Loss [BS83b, CC80, GNS^{+}81, Kum80, PLR82, Reu83, WM83]. Lossless [Mea80]. Low [PLR82, TA82, KM85]. low-complexity [KM85]. Low-Elevation-Angle [TA82]. Lower [AS83]. LPC [AO85, BR82b, BTOA84, RWA81, RSL83, RSL84]. LPC-Based [BR82b, RWA81, AO85, BTOA84]. LT [BMD84]. LT-1 [BMD84]. Lumped [Kne80]. Lumped-Element [Kne80].

M [SJ85, GAB^{+}80, HFE85, LTT^{+}83, RCS86, SB80, TZD^{+}83, Dos83, Fle84, Jag85b, SJ85, WC85]. M/G/1 [Dos83, Fle84, Jag85b, WC85]. M/M/1 [SJ85]. Machine [BJW88, Kar80a, FHK^{+}84]. Machine-Vision-Assisted [BJW88]. Made [Ben81a]. Magnetic [FLRT81, RB82]. Maintenance [AM80, BM82a, BKV82, Hol82, KMM82, Cal88, FGNS85, HLRW85, MSO84]. Making [Wei82, Rei85]. Management [BBN82, BP83, CR82, Coh82, CKP88, EH83a, Lin82, LL88, RS86, Yan85]. Manager [Roo82]. Managerial [Var81]. mandatory [FW88]. manual [Pul88]. Manufacture [BJW88]. Manufacturing [JPW87, KC87, LFHV88, BGGN^{+}86, CC88, FHH^{+}87, Fre86, KS86, McC86, DHM^{+}86]. Mapping [Dra81a]. Maps [CHM83, San85a, San85b]. Margin [Aca81b]. Markov [Jua84a, Jua85, LRS83, LMOS83, Maz82, RLS83, RLS84, RJLS85]. Markovian [MMR81, MM82, RM82]. Mask [And83]. Mass [ABC^{+}81, FKSW81]. master [BBM87]. Match [CK83]. Matching [JSC83, DL84]. Material [LBHB83, Nas81]. Material-Processing [LBHB83]. Materials [JDW89, Nas81, Pet84]. Mathematical [Elk80]. Matrix [Sal82]. Maximal [Koh83]. Maximum [Aca81a, Jua85, Kar83]. Maximum-Likelihood [Aca81a, Jua85]. Maximum-Power [Kar83]. Maxwell [PCPB80]. Mb/s [FPQ^{+}84, Wan83]. McDonald [Sha81]. MCVD [CP83a, CP83b]. Mean [AS83, Jag82a]. means [RWJ86]. Measured [JCV83, CMN84]. Measurement [AE83, EGKS81, Hoa81, KLRS81, SF81, Sze80, AO85]. Measurement/Trouble [AE83]. Measurements [Car80, CMSP81, Con80, CMN83, FYC81, LTT^{+}83, LW83, Phi82, Sem83b, Min84, Sem85]. measures [Jua84b, RS85a]. Measuring [Yum80, Ben84]. Mechanical [Chu83, Str88]. Mechanisms [LS80b]. Mechanized [DRT82, Rub82, Uhr82]. Mechanizing [GB82, MY83]. Media [Ber82]. Medium [HPU81, RLS84]. medium-size [RLS84]. Membrane [Loe83, SA85]. Memories [RB82]. Memory [AS83, HP83a, HK83, Kar80b, Sal82, Wyn81, NMR^{+}84, WWKZ84]. Memory-Constrained [AS83]. Memoryless [Row82]. MESFETs [Oga81]. Messagesin [FM82]. Metallic [CMN83]. Method [Chu83, Gri83, Len83, Sha81, Ack84, KS86]. Methodology [JLN^{+}85, PW88]. Methods [Ano83m, BKL^{+}82, Eig83, GL80, Jag84, LM81a, AH86]. metrics [Ing86]. Metropolitan [DFN81]. MFSK [Tim82]. MHz [CMN83, CMN84, WCS81]. Microprocessor [Bre81, Lec80, Rub82]. Microprocessor-Based [Lec80]. Microscope [SBD82, dlRR86]. Microwave [ADG^{+}83, BCHR83, BMM^{+}83, Chu80a, DRS83, FS83a, GMMR83, GK83, HP83b, HHS83b, HCH^{+}83, KP83, Kne80, Mar83, Rum80, Sch81, Vig81, GY85, Kav85b, RWRH85, YG85]. Mildly [Sal82]. Military [Lie82]. Millimeter [HM80]. Millimeter-Wave [HM80]. minicomputer [Tol84]. minicomputer-based [Tol84]. Minimal [Wit80]. Minimizing [Wit81]. Minimum [Aha82a, ALS80, CMS83, MS82, Str83]. Minimum-Cost [MS82]. Minimum-Delay [Aha82a]. Minimum-Distance [ALS80]. Minimum-Weight [Str83]. Misclassification [Hea81]. Mixed [Hef82]. Mixing [WM83]. mixture [FM84, Jua85]. Mixtures [Whi84c]. MLS [FW88]. Mobile [GHP80, HG81, LS80b, PS82, Reu83, Sun83a, Sun83b, Tim82, Gla84, Kar86]. Mobile-Satellite [Reu83]. Modal [CP83a, CP83b, WM83]. Mode [CMLF80, CMSP81, GPP87, JCMS82, KC87, LS80c, LTT^{+}83, Oga82, PPC81a, PPC81b, PLR82, Sem83a, WM83, AG84, SS84]. Model [Bre81, BS83c, CHN80, Coy82, Gol82, GC80, Gri82, Ram88, Rum82, Smi83, TRW82, Jua84a, Kli85, LF84, Log85, RJLS85, Ros84a, WC85]. Modeling [FH88, GC81, Fre86, PK87]. Models [Gav85, KSM81, LMM81, PW82, RLS83, Col88, Pet84, RLS84, SA85, WC85, Whi84d]. Modernization [BDL^{+}83, GY83, IST^{+}83, LW83]. Modes [Len83, PCP80]. Modulated [CB83, Kav85a]. Modulation [Aca81a, AFPP82, Ben80, BJM^{+}81, GS83b, Log83a, PS82, SW80, SA81b, Ste82, Sun82, Sun83b, CS85, Log84a]. modules [ABeB88]. Molecular [Pan89]. Molecular-Beam [Pan89]. Moment [BH84c, Hef82]. Moments [SY82]. Monitoring [Eic80]. MOS [And83, LMM81]. MOSFET [OOB83]. MOSFETs [Tay84]. Motion [NR80, RN82, NS85]. Motion-Compensated [NR80, RN82]. Movable [Max82]. Movement [Ros83]. Moving [SSW80]. Moving-Average [SSW80]. MPCS [DHM^{+}86]. MSK [Sun82]. MSK-Type [Sun82]. Mu [McN80]. Mu-255 [McN80]. MULGA [Wes81]. Multi [Hef82]. Multi-Job-Type [Hef82]. multiclass [RS85b]. Multihop [AKH87]. Multilevel [GN83, GHP80, Tim80]. Multimode [Kas82, WM83]. multiobjective [LF84]. Multipair [Lin80]. Multipath [BMM^{+}83, GC80, GC81, Rum80, Rum82, FPQ^{+}84, Kav85a, KS85, RWRH85]. Multiple [AMS82, Dat80, Hey82, Sal82, WW81, WWZ80, Yue80a, Kam89b, Kam89a]. Multiple-Access [Hey82, Yue80a, Kam89b, Kam89a]. Multiple-Input [Sal82]. Multiple-Output [Sal82]. Multiplex [DRS83]. multiplexed [Pik84]. Multiplexer [WA88]. multiplexers [Cho86]. Multiplexing [EY81, EH83b, EHS83, Has81, Sch83b, Kam89a]. Multiplier [FD82]. multipoint [Col88]. Multiprocessor [Bac84]. Multiprogrammed [Fre80a, RS85c]. Multistage [Ste82, Tim81]. Multitasking [BM87]. Multitone [GC81, Tim82]. Multivariate [Lee82, Jua85].

Names [RRW80, SAT83]. NaOH [KKLS83]. National [BP83]. Natural [HBH88]. Near [BGGW83, RCS86, Sem83b]. Near-In [Sem83b]. Near-Real-Time [BGGW83]. Network [Alb83, AE83, AM82, BO83, BDL^{+}82a, BPRS83, BP83, BBP^{+}82, BC83, BGGW83, CGH81, CKP88, CRT82, DFS82, EH83a, EY82, FH80, FM82, FHR84, FGP^{+}83, GLM82, GOP87, GFGB83, HSS82, HKRS83, HM83, HMSY82, KBN88, LLS82, LF82, Man89, Mar88, MRW81, Nea80, Nus82, PW82, Pfe83, ST82, Sch83a, SF80, SW82, WA88, Whi83b, Whi83c, WPG87, ACL88, Cal88, CCD^{+}84, Eck86, HB85, Hon84, LMS^{+}89, MZCW84, Sur86, TNG88, WM88, Yoo85]. Network-interface [GOP87]. Networking [BKP88, DKW88, GS88, GK88]. Networks [AKH87, ADG^{+}83, Alb82, AS83, ACM81, AKK81, Ben81a, Ben83, CB81, Chu80b, CD80, Els80, GPP87, Hef82, KSB88, KK82, Kum83, Kur80, Max82, MMR81, MM82, Mea80, MS82, Mor81a, PW88, RM82, Wit80, Aga89, Aki84, Col88, DeT84, DN88, Hei89, HJG88, Kam89b, Kam89a, KV84, Kum88, Whi84d]. neural [HJG88]. Newton [San80, San81b]. Newton-Direction [San80, San81b]. Nmake [CF88]. No [ACSW82, AFS82, ADRW83, ABC^{+}81, AFHS82, ARW80, BBN82, BW83a, BDV83, BDL82b, BDL^{+}83, BNS82, CH81, CHTW83, CS83, DVW83, DG81, DM83, FKSW81, Gad81, GLM82, GKS83, HHS83a, HMPV81, JKT82, KMM82, Mar81, MRW81, NS82, SC83, Wil82]. Nodes [Alb82]. Noise [Ben80, Ben81b, Jay83, Oga81, Ste82, Sun82, Bat84, Rai84]. Noisy [SSR81]. Noncolocated [EH83b]. nondispersive [Kav84]. nonhierarchical [Aki84]. Nonlinear [BB87, LF84, MZ82, Sal82, San81a, San82a, San82b, San85b, Min84]. Nonlinearities [Row82]. nonproduct [Yoo85]. Normal [MM82, SY82]. Note [Rab82, RSL83]. NTSC [Has81]. Number [Rob82, Jay87]. Numerical [Ack85, GH81, Kum80, PCPB80, PCPF81].

O [KKLS83]. Objectives

[BPRS83, DG81, SC83]. observations [Jua85]. Obstruction [Vig81]. Occurrence [Ben81b]. Off [PKSG83]. Off-Line [PKSG83]. Offered [Bur82]. Office [BO83, BKV82, CCD^{+}84, GFGB83, HLL^{+}84, McN80, War80]. offices [Lav87]. Offset [Sem83a]. OH [Phi82]. On-line [GB82]. once [WWKZ84]. One [LFG82, LF84, Row84b, Sha88]. one-dimensional [Row84b]. one-way [Sha88]. Open [Whi84d]. Operated [WCS81]. Operating [BS83a, BW83b, EC83, GKS83, GCD83, HP83a, HPW83, HK83, KAM83, KM83, KBF83, MBD^{+}83, Pfe83, QEG83, Roc82, RBH83, RW83, Sca83, TG83, Wei82, YMNW83, Bac84, BHK^{+}84, GM84]. Operation [GMW82, Min84, MSO84]. Operational [BDL^{+}82a, BBP^{+}82, LFG82, DFOS85]. Operations [FH80, Hol82, McC86, FGNS85, Tol84]. Operator [DHK89, BLT89]. Operators [Var81]. Optic [BS83b, Pau83, TNG88]. Optical [CP83a, CP83b, CC80, DKZ80, Eic80, GNS^{+}81, JPW87, Kas82, KBZ80, KFS83, Len83, LBHB83, Nas81, Oga81, OOB83, PLR82, PS81a, Phi82, Pre81, PC82, RCS86, Smi82, SF82, SB80, TZD^{+}83, KD84, SCWB84]. Optical-Fiber [Phi82].

Optically [MDSL82]. Optimal [CV85, MZ82]. Optimally [Lav87]. Optimization [ACM81, CMSP81, Pha86, SMNV89]. Optimum [GP84, PCPF81]. Order [Hon83, Son83]. Ordered [Joh81]. Oriented [BK83, Fre80a]. OSPS [BLT89, BBD^{+}89, CDG^{+}89]. Outdoor [Coy82]. Output [BW83b, LS80c, Sal82, Rit84b, San85a, San85b]. Overall [BDL^{+}82a, BBP^{+}82]. Overflow [KSM81, Mor80a, Mor80b, Mor81b, MW82]. overload [Aki84]. Overtone [SCWB84]. Overview [ACSW82, ALS88, BBF^{+}81a, Bod81, CL82, DR81, Gig80, GK88, GY83, Hei89, HMSY82, KAM83, Pra83, RM82, SC83, TG83, CCG^{+}85, Ran84, SKP84]. Oxides [RS88].

Package [RM82]. Packages [KSM81, Rai84]. Packaging [AB89, GNS^{+}81, KBF83, AHC87]. Packet [HTB83, KSB88, Kar86, WA88, BN84, Lo87]. Packet-Switched [KSB88]. packet/circuit [BN84]. Page [Roo82]. Pair [LFG82]. PANACEA [RM82]. Papers [Ano80k, Ano80l, Ano80m, Ano80n, Ano80o, Ano80p, Ano80q, Ano80r, Ano80s, Ano80t, Ano81o, Ano81p, Ano81q, Ano81r, Ano81s, Ano81t, Ano81u, Ano81v, Ano81w, Ano81x, Ano82v, Ano82w, Ano82x, Ano82y, Ano82z, Ano82-27, Ano82-28, Ano82-29, Ano82-30, Ano82-31, Ano83s, Ano83t, Ano83u, Ano83v, Ano83w, Ano83x, Ano83y, Ano83z, Ano83-27, Ano83-28]. Parabolic [Sem83a, WM83]. Parallel [CDH81, Cho86, DH82, FD82, HHL89, LMP83]. Parameters [RWA81, RSL84]. Parasitic [LFG82]. Partial [Sun82]. Partial-Response [Sun82]. Pass [RW81]. password [BCW86]. past [Ros84b]. Path [Chu80a, Reu83]. Paths [LBP80, Rum80, Sch81]. Pattern [JSC83, RW81]. Pattern-Recognition [RW81]. Patterns [CGH81, Eic80, NWB^{+}80, Sha85]. pay [Bul87]. pay-per-view [Bul87]. PBX [KSM81]. PCM [SSW84, SWS84a, SWS84b]. Peak [Con80, Ste82, SB80]. Peakedness [Bur82]. Penetration [Wal83]. People [Ega83]. Perception [JB83]. Performance [Aca81a, AE83, Ano83n, AK88, BS88, Ben83, BDE^{+}81, BGGN^{+}86, Col88, DRM83, DG81, DFS82, Fla83, Fre86, FH88, FLGD83, GMMR83, GH82, Hey83, HTB83, KAM83, KSB88, Kav84, KM85, LZ82, Lo87, Loe83, MR83, McN80, PS81b, Rab82, She88, Whi83b, Yav82, Yue80a, Aki84, But84, CCD^{+}84, Fed84, Fos84, Jua84b, Kum88, LL88, RPS84, Ros84a, Tur85, Wer85]. Period [Yum80]. Periodic [Gra83, San81a, SA85]. Peripheral [AFS82]. Permeability [Ber82]. Permittivity [Ber82]. permutation [Jay87]. Permutations [JCMQ83]. Personnel [CK83]. Perspective [Nus82, Pfe83]. Phase [JDW89, Kar83, LS80a, LM81b, Log83a, PS81b, Sun82, Sun83b, WSX82, Yue80a, Yue80b]. Phase-Shift [Yue80b]. Phase-Shift-Keying [PS81b, Yue80a]. Phased [DG80]. Philosophy [Fra83, O’N80]. Photodetector [SF82]. Photodiodes [GAB^{+}80, SCS85]. Photonic [DSMBB89, Hin87, JDW89, Kam89b, Kam89a, MS89, AB89]. Physical [AFHK82, CS82, HP83b, JKT82, LFLB85, Smi82, BH84a]. Pico [Hol87]. Picosecond [LTT^{+}83]. picture [Hol87]. Pictures [GN83, JN81, Joh81, JSC83, NB82, Gha84]. Piecewise [Lee82]. Pilot [LS80b]. Pilot-Based [LS80b]. PIN [GAB^{+}80]. PKM [SS86]. placement [HJ88]. Plan [BNS82, HKRS83, Hoa81, KP88, LLS82, Pha82, HLL^{+}84, KV84]. Planar [KBZ80]. Plane [DG80]. Planning [ABS82, BW83a, BKL^{+}82, CP86, GH82, HPS86, KP88, War80, BO85, GE87]. Plant [Dat80, DFN81, DK83, GL80, Gol82, Bat84]. Point [Yue80b, Lav87]. Polarization [SS84, Gre85, Kav85a, KS85]. polarized [AS84b, Kav84, KS85]. Pole [LFG82]. policy [FW88, GGV84]. Polynomial [GC80, GC81]. Population [Hea81, WR83]. Populations [ND81, Wil85]. Port [Kur80, ABC89]. porting [BHK^{+}84]. Position [ADRW83, ARW80, BW83a, Ben80, BDV83, CHTW83, CS83, DVW83, DM83, GKS83, HHS83a, SC83]. Positive [Sch81]. Post [Jay81a]. Post-Filtering [Jay81a]. postfiltering [RJ84]. Potential [FLGD83]. Power [FYC81, Kar83, KLSW89, Lin81, LBHB83, Reu83, Sal80, SS83, SY82, Min84]. Powered [MDSL82]. Powerful [CDK88]. Practical [Tow83]. Practice [Eic80, Fuc86, Lin84]. Preattentive [JB83]. Precedence [Mor81b]. Precision [LS80a]. Prediction [Sch81, SSW80]. Predictions [Sch83b]. Predistortion [HHS83b]. Preemptive [HTB83]. Preface [Mar84, SB82]. Preparation [Ker89]. Presence [Sze80, Lav87]. Presentations [MF82]. Pressure [KKLS83]. Pressure-Volume-Temperature [KKLS83]. Primal [MS82]. Primary [MW82]. principle [WC85]. Printed [Rai81]. Priority [HTB83, Mor81a, WA88, Dos85, Lo87, RS85b]. priority-based [RS85b]. Privacy [CT83, CBB^{+}87, Jay87]. Private [KK82, MBJ81, She88, Col88]. Private-Line [She88, Col88]. Probabilistic [LRS83, MS84, Ros84a]. Probabilities [DH82]. Probability [Son83, Sun82, GGV84]. Probing [GC81]. Problem [MZ82, Oza80, Sal83, Sha81]. Problems [Log83b, Mor80a, Mor80b, LMS^{+}89, LF84]. Procedure [RBW82, RWJ86]. Process [And83, DHM^{+}86, LRS83, Nuc81, Fow86, HJ88, Pet86]. Processes [Hef80, SSW80, Son83, CV85, KS86, Leu86]. Processing [DRM83, LMP82, LMP83, LBHB83, RBH83, Row84a, AAJ^{+}86, CF86, HBH88, JLR^{+}86]. Processor [Aag81, Ang81, BBF^{+}81a, BS83a, BBF81b, Bod81, BDE^{+}81, BJM^{+}81, BST81, BGK^{+}86, BW83b, BKO81, Cox81, Cro81, DVW83, EC83, EG81, Fav81, FYC81, Gad81, GCD83, HP83a, HPW83, HK83, KAM83, KM83, KBF83, MBJ81, MBD^{+}83, QEG83, RBH83, RW83, Sca83, Sem81, TG83, YMNW83, BTOA84, SJ85, WM88]. processor-sharing [SJ85, WM88]. Product [CF88, LMOS83, Mar88, BCH^{+}86, GMT87]. Production [CR82, Tho80]. Productivity [FS88, FHH^{+}87]. products [CLMS87, Log84b]. Profile [CP83a, PPC81b, Pre81]. Profiles [PPC81a]. Program [AM82, BDL^{+}82a, BBP^{+}82, CRT82, DR81, DFS82, DS84, EY82, FM82, GLM82, Geh81, GFGB83, Gri83, HSS82, HMSY82, LLS82, PK84, Pra83, ST82, SW82]. Programmable [Kar80b]. Programmer [MT80]. Programming [Elk80, Jon88, Sha81, Ves88, LF84]. Programs [Rei81, Car84]. Project [KPW85]. Projection [IG82, Mor82]. Prologue [AM82, BHS82, Mar83, Mar81, Sca83]. proof [CC85]. proof-test [CC85]. Propagation [Ber82, CMLF80, NWB^{+}80, TA82, FPQ^{+}84, Pet84]. Properties [Chu83, EY81, PLR82, RSL83, RJLS85]. Proportions [Hea81]. Proposed [Kur85]. Protocol [FHK^{+}84, Hey82, Kur85, ASG85]. protocols [Hol85]. Proton [DJ80, HHG83]. Proton-Bombarded [HHG83]. Prototype [Gog82]. prototyping [BCMZ88]. Provide [Jay81b, RB82]. Providing [Tho80]. Pseudo [Rob82]. Pseudo-Random [Rob82]. Public [KRS83, CCD^{+}84]. Pulled [JCV80]. Pulse [AFPP82, BJM^{+}81, CB83, GS83b]. Pulse-Code-Modulation [BJM^{+}81]. Punctures [Ung80]. Purpose [LMP82, AAB^{+}86]. Pyramidal [But81, Sem83b, Sha85, Sil84].

QAM [FPQ^{+}84, Fos84, Kav84, Pra80]. QMP [Hoa81]. Quadrature [Aca81a, Mey82, Kav85a]. Quality [Eck86, EHS83, Fuc86, GMT87, Hoa81, Pha82, PKSG83, RS88, SSR81, Sur86, FT87, God86, GST86, Ing86, Pet86, SS84, WRM84, Wil85]. Quantitative [CHM83]. Quantization [McN80, RLS83, RPS84, RS85a, SRJR87]. Quantizer [RSL83, RSL84]. Quarter [DL84]. Quarter-wave [DL84]. Quartz [KKLS83]. Quest [KC88]. Queue [Dos83, HTB83, Mor80b, MW82, Fle84, Jag85b, Lo87, SJ85]. Queueing [AS83, BS83c, FM84, Fre82, Hef82, Kum83, MM82, RM82, Whi83b, Whi83c, Jag85a, RS85b, WC85]. Queues [Jag82a, KW84, Whi84b, Whi84c, Whi84d]. Queuing [Hef80, MMR81, Mor81a, Mor80a, Mor81b].

Radiated [Reu83]. Radiation [But81, Sha85]. Radio [ADG^{+}83, BCHR83, BMM^{+}83, CMN83, DRS83, FS83a, FS83b, GMMR83, GK83, GHP80, HP83b, HHS83b, HCH^{+}83, HG81, KP83, Mar83, PS82, Rum82, SS83, Sch81, Sun83a, Sun83b, Tim82, Vig81, Wal83, Wan83, AS84b, FPQ^{+}84, Fos85, Gla84, Gre85, GY85, Kar86, Kav85a, Kav85b, Row84b, RWRH85, YG85]. Radio-Line [HP83b]. Rain [Aca81b, LBP80]. Raman [SCWB84]. Random [Mey82, Rob82, Son83]. Range [BW83a]. Rapid [Mue81b]. Rate [Aha82a, AS83, CB83, DGH^{+}82, GS83a, Jay81b, Jay83, KSB88]. Rationale [Mac83]. Rayleigh [SSW84, SWS84a, Sun82, SWS84b, Yue80a]. RC [Ste82]. RC-Shaped [Ste82]. read [HJ88]. read-site [HJ88]. Reading [CK83]. Readings [AT&T86]. Real [AKK81, BBN82, BGGW83, CKKS88, GKS83].

Real-Time [AKK81, BBN82, CKKS88, GKS83]. Realization [CGH81, BCH^{+}86, GMT87]. Realized [LFG82]. Reasoning [Ega83, Bra88]. Receiver [BST81, HCH^{+}83, OOB83, Sun82]. Receivers [CC80, Oga81, Oza80, SF82]. Reciprocal [Mea80]. Recognition [BR82b, Hol83, KP88, LRS83, MR81, MRR81, RW81, RBW82, RRWK82, Rab82, RLS83, RRW80, TRW82, WR83, AO85, AAB^{+}86, BTOA84, Jua84a, Rab84, RWT84, RLS84, RSL84, RS85a, RWJ86, SRJR87]. recognize [Wil85]. Recognizer [DRM83, RWA81]. recognizers [RPS84, Ros84a]. Reconstruction [NB82]. Record [Lin82]. Recorded [RB82]. Recovery [GM87, HPW83, Wil82, Log84d, Log84b, Log84c]. Recursive [Hon83, Shi82]. Reduce [DG80, SWS84a]. Reduction [Sem83b, SB83, SB80]. Refinement [Geh81]. Reflection [Cla83, KD84]. Reflections [Sem80]. Reflector [But81, Dra81a, Sem80, Sem83b, Sha85, Sil84]. Reflectors [DG80]. Refractivity [Sch81]. Region [LTT^{+}83]. Regular [DH82]. Rehabilitation [SF80]. rehoming [Lav87]. Related [Geh81, Hef80]. Relational [Lee83]. Relationship [KKLS83]. Relationships [Inf80]. relay [Col88, DN88, Kum88]. Reliability [AM80, BKV82, CLMS87, CS83, SCS85, Tor81, Toy87, AH86, HFE85]. Reliable [DKZ80, KBZ80]. Relief [Koo80]. Remote [ACSW82, AFS82, AFHS82, BDL82b, BDL^{+}83, BNS82, Gig80, JKT82, KMM82, NS82, Pfe80]. remoted [Leu86]. Repair [BR82a, BBOV82, BM82b, CHW82, DRT82, FV82, GH82, GB82, Hol82, LZ82, Mar82, Ove82, Pau83, RD82, Rub82, SB82, Uhr82]. Repertory [RWR80]. Repertory-Dialer [RWR80]. Report [RD82]. Reports [Ega83, GFGB83]. Representation [Rum80, Bra88, Sha85]. Representations [MMR81, MM82, RJLS85, San85b]. required [Whi85]. requirements [MS84, PK87]. Research [KC87, AR86, BBW88, IABB^{+}89, Li87, McC86, Rei85]. Reservation [LS80c]. Resource [Aca81b, SW81, CV85]. resources [GE87]. Response [Sal80, San81a, Sun82, Ben84, But84, SA85, SJ85]. Responses [GC80, GC81]. Results [Gin83, Has81, Hef80, KLRS81, NR80, Row82, Wan83, MM85]. Retrial [Liu80]. retrieval [RWT84]. Retrofitting [DVW83]. Retrospective [Ega83, Ker89]. Reuse [Aca80, ABeB88]. Reveal [Ega83]. review [FT87]. Ribbons [BS83b]. ring [WC85]. robin [Fle84]. robot [And88, Bro85]. Robotic [LBHB83, Dec88, Str88]. Robotics [CC88, CKKS88]. Robust [KS86, Mor82]. Role [NS86]. round [Fle84]. round-robin [Fle84]. routine [Lin84]. Routing [ACM81, AKK81, FM82, Aki84, GGV84, Kam89a]. Rule [Ves88]. Rule-Based [Ves88]. Rules [Mey82]. running [FHR84].

s [BBDS86, CCJS82, DGH^{+}82, BC85, FPQ^{+}84, Wan83]. Sable [CF88]. Saddle [Yue80b]. Saddle-Point [Yue80b]. Safe [Alb82, Wei82]. Saito [Jua84b]. Sample [SB83]. Sampling [Lie82, ND81, HLL^{+}84]. Sandia [TNG88]. SARTS [Gig80]. Satellite [Aca81b, DG80, EH82, EH83b, EHS83, LBP80, Reu83, TA82, Yue80b]. Satellites [Aca80]. saturated [Tay84]. Scalar [PCP80]. Scale [BBF^{+}81a, BBW88, Gav85, Kli85]. Scaling [MF81]. scan [BH86, GP84]. scan-width [GP84]. Scanned [SBD82]. Scanning [Con80, PS81a, dlRR86]. SCARAB [FLRT81]. Scatterers [Blu80a]. Scattering [Blu80a, Eic80]. Scheduler [Hen84]. schedules [Ack85, Dos85]. schedules-high-priority [Dos85]. Scheme [SA81a, Tim80, RS85b]. Science [Ano83n, Ano83q, Ano83o, Ano83p, CHM83, CK83, DK83, Ega83, Eig83, Fla83, Fra83, FLGD83, Gin83, Han83, Hol83, JB83, KRS83, Koh83, Loe83, Mac83, Pau83, Ros83, SAT83, Sum83, Rei85]. Scramblers [JCMQ83]. Scrambling [CT83, Cho86, Jay87]. screen [EEF^{+}86]. SDLC [Kum88, WA88]. Search [ALS80, FT83]. Secant [PCP80]. Second [Rub82]. Second-Generation [Rub82]. Secondary [Mor80b]. Section [GL80]. Secure [KBN88]. Security [ALS88, BN88, PW88, RW84, BCW86, GM84, KC88]. segmental [RWJ86]. Selected [DRM83]. selection [GP84]. Selective [Sem83b]. Self [Jay81b]. Self-Contained [Jay81b]. semantic [WRM84]. Semantics [FLGD83]. Semiconductor [DSMBB89, SW83, CEGM89, JLN^{+}85]. Sense [Hey82]. Sensitive [Hol82, EEF^{+}86]. Sensitivity [CC80, Lon80, SF82]. Sequence [Aca81a]. sequences [Rab84]. Sequential [CB81, IG82, JCMQ83, Mor82]. Serial [Aha80]. Series [CDH81, DH82, Kne80, Log83b, KLSW89]. Series- [Kne80]. Series-Parallel [CDH81]. Serve [SA81a]. Service [ADRW83, ARW80, BBP^{+}82, BW83a, BR82a, BBF81b, BST81, BDV83, BBOV82, BM82b, CHW82, CHTW83, CRT82, CS83, DVW83, DRT82, DG81, DM83, EY82, FV82, GH82, Gig80, GKS83, GB82, HSS82, HHS83a, Hol82, LZ82, Mar82, Ove82, RD82, Rub82, SB82, SW82, Smi83, SC83, Uhr82, Fle84, Ran84, Whi84a, Whi85]. Services [Eig83, IG82, Nuc81, Wor82, BLT89, CDG^{+}89, DHK89, Hei89, MZCW84, ML84]. Servicing [AKK81, Sze80, Sze82]. serving [Lav87]. Sets [Mea80]. Several [Hea81, MR81, RS85a, Whi85]. Shallow [DJ80]. Shape [JCV83, Koh83, KW84]. Shaped [Ste82]. Shaping [SW80]. Share [Hen84]. Sharing [Aca81b, SW81, Rit84a, SJ85, WM88]. Shield [Ung80]. Shielded [Sem83a]. Shift [Ben80, PS81b, Yue80a, Yue80b]. Shifted [RCS86]. Short [DKZ80, OOB83, Sze82, Tay84]. Short-Channel [OOB83, Tay84]. Short-Haul [DKZ80]. Short-Term [Sze82]. Shunt [Kne80]. Shunt-Tuned [Kne80]. Si [OOB83]. Si-MOSFET [OOB83]. Sideband [ADG^{+}83, BCHR83, BMM^{+}83, DRS83, FS83a, GMMR83, GK83, HP83b, HHS83b, HCH^{+}83, KP83, Mar83, PS82]. Sidelobe [Sem83b, Sil84]. Sidetone [Mea80]. Sight [Rum80, Sch81]. Signal [Aag81, Ang81, BBF^{+}81a, BBF81b, Bod81, BDE^{+}81, BJM^{+}81, BST81, BGK^{+}86, BKO81, Cox81, Cro81, DRM83, EG81, Fav81, FYC81, Gad81, JCV83, KBG^{+}88, Lon80, MBJ81, Sem81, SSR81, Ste82]. Signal-to-Noise [Ste82]. Signaling [CRT82, Fan81, FM82, GLM82, Fos84]. Signals [CB83, EY81, EH83b, EHS83, Fla80, GC81, Has81, JS82, Kur80, Log84d, Log84b, Log84c, MF81, Sch83b, Ste82, SB83, TA82, Wal83, WSX82, Kav84, Kav85a]. Silica [PCPF81, SCWB84]. Silicon [BBF^{+}81a]. Simple [Fre80b, Sha85, Log85]. Simulation [Aha82b, FH80, Has81, LMM81, LM81a, LMP82, LMP83, DeT84, Gre85, GY85, RWRH85]. simulation-based [DeT84]. Simulations [EG81]. Simulator [Aag81, Man89]. Simultaneous [LM81b, SW80, SV81]. Simultaneously [PLR82, Whi85]. Single [AO85, AAB^{+}86, ADG^{+}83, BCHR83, BMM^{+}83, CMSP81, DRS83, DC80, FS83a, GMMR83, GPP87, GK83, HP83b, HHS83b, HCH^{+}83, JCMS82, KC87, KBZ80, KP83, LG84, LTT^{+}83, Mar83, Oga82, PPC81a, PPC81b, PLR82, PS82, RS85a, Tow83, AG84, Eis84, SS84]. Single-board [AAB^{+}86]. Single-Chip [DC80, AO85]. Single-frame [RS85a]. Single-Heterojunction [KBZ80]. single-layer [Eis84]. Single-Mode [CMSP81, GPP87, JCMS82, KC87, LTT^{+}83, Oga82, PPC81a, PPC81b, PLR82, AG84, SS84]. Single-Sideband [ADG^{+}83, BCHR83, BMM^{+}83, DRS83, FS83a, GMMR83, GK83, HP83b, HHS83b, HCH^{+}83, KP83, Mar83, PS82]. SiO [KKLS83]. Site [CM80, HJ88]. Size [Koh83, RLS84]. Sizing [Noe81]. Skills [CK83]. SLC [AS84a, BH84a, BBB^{+}84, Can81, COW82, LOS84, MMTV84, MSO84]. SLC-96 [Can81, COW82]. Slots [Max82]. Slow [Sun82]. Small [BO83]. Smoothing [Feu83, Gra83]. SNA [Kum88]. SNA* [WA88]. SNA*/SDLC [WA88]. SNA/SDLC [Kum88]. Soft [SWS84a, EEF^{+}86]. Software [Aag81, BT83, BR82a, BDL82b, BCMZ88, CH81, CRT82, CKKS88, FS88, FT87, Fra83, GKW82, Gin83, HHS83a, HPU81, LSY88, Mac83, QEG83, RM82, RW83, ABeB88, BBW88, DFOS85, EP84, Ing86, Leu86, RS86, SKP84]. Sojourn [RS85c, Fle84]. solder [Ray88]. solid [Min84]. solid-state [Min84]. soliton [Pet84]. solitons [Pet84]. Solution [MZ82, PCPB80]. Solutions [Log83b]. solver [DS85]. Solving [LMS^{+}89, LF84]. Some [BBW88, Maz82, Mor80a, Mor80b, ND81, NR80, Rab82, RJLS85]. sort [Lin84]. Source [GS83a, Oza80, Wit80, WW81, WWZ80]. Source-Coding [Oza80]. Sources [AMS82]. South [Car80]. Space [War80, GY85, YG85, Chu80a, Kar83, Rum82]. Space-Diversity [Rum82]. Spaced [GW81, GMW82]. Spanning [Str83]. Spanning-Tree [Str83]. sparing [Tol84]. Spatial [DG80, RN82, Row84b, Sha85]. SPC [DFS82, GLM82, HSS82, SW82]. Speaker [RLS83, RRW80, RLS84, SRJR87]. Speaker-Independent [RLS83, RRW80, RLS84]. Speaking [BGS^{+}89]. Special [BBF81b, FS83a, Feu81, Gig80, IG82, LMP82, Nuc81, Smi83]. Special-Information-Tone [Feu81]. Special-Purpose [LMP82]. Special-Service [BBF81b, Gig80, Smi83]. Special-Services [IG82]. specification [BCMZ88, Kur85]. Specified [Son83]. Spectra [JCMS82, SCWB84]. Spectral [EY81, LTT^{+}83, SW80, Son83]. spectrometer [Row84a]. Spectrum [Ein80, KM85]. Speech [AAJ^{+}86, AR86, BKO81, Cox81, CCJS82, CF86, GS83a, Hol83, JC80, Jay81a, Jay81b, JCMQ83, JLR^{+}86, LS80c, LRS83, MF81, Max80, SSR81, SV81, SB83, WSX82, AO85, AAB^{+}86, BGS^{+}89, Jua84a, Jua84b, RPS84, RJ84, WRM84, Wil85]. Speed [DKZ80, GAB^{+}80, SF81, Kav85b]. Spelled [RRW80]. Splice [Kum80]. Splices [WM83]. Splicing [Pau83, AFM87]. Splines [Gra83]. Splitting [Wit81]. Spoken [RRW80]. Spotting [MRR81]. Spread [Ein80, KM85]. Spread-Spectrum [Ein80]. Square [Ben81a]. Squares [Ben80, Hon83, MR83, Mue81a, Mue81b]. SSB [LS80b]. Stability [Kar80a]. Stabilized [SW83]. Stable [GMW82]. Stages [Ben81a]. Standard [Ing86, KRS83, Lie82, BBDS86]. Standards [Ald86, BN88, JLR^{+}86]. Star [Alb83]. StarKeeper [Mar88]. Start [Sal83]. Start-Up [Sal83]. State [Kar80a, CS85, Min84]. States [Ben81a]. station [Gla84]. Stationary [Ack84]. Statistical [AH86, FLGD83, Lin80, Lin81, Rum82, Var81, WA88, Eck85]. Steady [Kar80a]. Steady-State [Kar80a]. Stepwise [Geh81]. sticky [Wit84]. Stochastic [AMS82, GH81, MY83, Jua85]. Storage [Ano82f, BHS82, BKV82, CS82, GKW82, Max80, Nus82, Wor82, Wit84]. Store [Roo82]. Stored [AM82, BDL^{+}82a, BBP^{+}82, CRT82, DFS82, EY82, FM82, GLM82, GFGB83, HSS82, HMSY82, LLS82, ST82, SW82]. storeroom [Pul88]. Strategies [Uhr82, Wil82, SWS84b]. Strategy [BT83]. Stream [Rit84b]. stress [CC85]. Structural [BB87]. Structure [CH81, CP83a, CP83b, Roc82, Shi82]. Structured [ND81]. Studies [Aha82b, ABS82, CMSP81, EY82, LBP80, Pha86]. Study [AM80, BS88, CK83, EH81, Gin83, KFS83, MR81, Pre81, Var81, Gre85, GY85, Wil85, CCD^{+}84, HLL^{+}84]. Stylistic [CK83]. Sub [Cro81]. Sub-band [Cro81]. Subjective [CHN80]. submarine [HFE85]. Submersible [FLRT81]. Subnetworks [Coh82]. Subrates [AGS82]. Subsampling [Jay81b, RN82]. Subscriber [Aha82b, AS84a, BH84a, BBB^{+}84, LOS84, MMTV84, MSO84]. Subsequent [SB83]. Substitution [JSC83]. Subsystem [ABC^{+}81]. Subsystems [YMNW83]. Suburban [BDL^{+}83, CD80, GY83, IST^{+}83, LW83, CMN84]. Sum [Lin81]. Summary [LBP80]. Sums [SY82]. Supply [GK83]. Support [Bod81, Coh82]. suppression [Sil84]. Surface [Dra81b]. Surveillance [BGGW83]. survey [Bat84]. survivable [Aga89]. Switch [Hol82, BN84, CP86, DHK89, HPS86, WM88]. Switched [KSB88, LFG82, Pfe80, Cal88, CCD^{+}84]. Switches [Ben81a]. Switching [ACSW82, AFS82, AFHS82, BDV83, BDL82b, BDL^{+}83, BNS82, Chu80b, JKT82, KMM82, NS82, Smi82, Wil82, BHK^{+}85, CCG^{+}85, DFOS85, FGNS85, Gla84, GGV84, HLRW85, Hin87, MMTV84, TNG88]. Symbolic [Wes81]. Symmetric [Sal80]. Synchronization [EH83b, MRW81]. Synchronous [Kar80a, Yum80, Col88]. syntactic [WRM84]. Syntax [BK83]. Syntax-Directed [BK83]. Synthesis [BKO81, RS88, Kow88]. System [AT&T86, ACSW82, ADG^{+}83, AFS82, ADRW83, Aho84, AFHS82, AE83, And83, AMS82, Ano82f, AGS82, ARW80, AM80, BCHR83, BBN82, BS88, BO83, BDL^{+}82a, BPRS83, BS83a, BP83, BW83a, BHS82, BC83, BDV83, BM82b, BKV82, BDL82b, BDL^{+}83, BW83b, BNS82, BMM^{+}83, BGGW83, Can81, CHW82, CR82, COW82, CHTW83, Coh82, CS82, CBB^{+}87, CS83, DVW83, DM83, DRS83, DGH^{+}82, EH83a, Ein80, EC83, EY81, EH83b, EGKS81, FS83a, FV82, Fre82, FGP^{+}83, GMMR83, GKW82, GKS83, Gog82, GK83, GFGB83, GCD83, HHS83a, HP83a, HKRS83, HPW83, HP83b, HM83, HHS83b, HCH^{+}83, Hen81, HK83, Hol82, JKT82, KAM83, KMM82, Ker89, KP83, KM83, KKLS83, KBF83, LS80c, Lie82, Lin82, Mar83, MR83, Mar88, Mar82, Mar84, MBD^{+}83, Mor81b, MW82, NS82, Nus82, Pfe83, Pfe80, PK84]. System [QEG83, RWR80, RW84, Roc82, RBH83, RW83, Rub82, Rze83, ST82, Sca83, Sch83a, SC83, Tim80, Tim81, Tim82, Tol84, TG83, Ves88, WH82, Wes81, Wil82, Wor82, YMNW83, Yue80a, AS84a, AAB^{+}86, BH84a, BH84b, BHK^{+}85, BBD^{+}89, BC85, Ben84, BCW86, BHK^{+}84, BBB^{+}84, BH86, CCG^{+}85, CHL^{+}89, DTV85, DFOS85, Fed84, FMM84, FF88, FHR84, FGNS85, GM84, GE87, HJ88, HLRW85, HLL^{+}84, HHL89, Jay87, LOS84, MMTV84, MSO84, PR86, PK87, RS85b, RS85c, Rit84a, Rit84b, SMNV89, Yan85, AK88, Bat84, BBB^{+}85, DHM^{+}86, FMM84, FW88, Gig80, Hol83, KBN88]. System-1B [AM80]. System/370 [FMM84]. system/C [FF88]. Systems [AFPP82, ADG^{+}83, AFS82, Aho82, AHC87, BBN82, CR82, Coh82, CT83, Fre80a, Fre80b, FGP^{+}83, FLGD83, Gog82, Gol82, GFGB83, HP83a, HK83, HTB83, Kas82, LDGF82, Lin82, Oga82, PS81b, Ram88, Reu83, Roc82, Roo82, Sal82, SS83, San81a, San82a, San82b, San87, SW81, Sun83a, Sun83b, WH82, War80, Wei82, Yav82, And88, Bac84, Cal88, Eck86, Fos85, FHH^{+}87, Fre86, Gav85, Gla84, GMT87, HFE85, HB85, Li87, MS84, NS86, Rog87, Sur86, Tol84, Whi84a].

T [DR81, Yan85]. T-Carrier [DR81]. T1 [EGKS81, KLRS81]. Tailoring [RCS86]. Takes [Mor81b]. Tap [GMW82, GM87, OW84]. Tap-Leakage [GMW82]. tasks [Dos85]. TCM [EHS83].

TD [Wan83]. TDM [Max82]. TDMA [BH84b, Hon84]. Technical [CK83]. Technique [ALS80, EH81, Fre80b, Jag82b]. Techniques [CGH81, DRM83, EH82, JPW87, LZ82, MF81, Pau83, SV81, Uhr82, BGGN^{+}86, Cho86, GST86]. Technological [Ano83q]. Technologies [DSMBB89, Dut89, BGS^{+}89, BCMZ88, CC88, Hei89, MS89, SP89, SKP84]. Technology [Ano83p, HHC87, KR80, KBF83, LSY88, Str88, AB89, AFM87, BH88, CF86, DD87, GMT87, Hin87, Tho87]. TELBECC [Gri83]. Telecommuncations [MBD^{+}83]. Telecommunications [PC82, LMS^{+}89]. Telephone [Ben83, BBF81b, CHN80, Eig83, Gol82, Gri83, Inf80, Mea80, Tor81, WR83, WRM84, Wil85, You84]. telephone-quality [WRM84, Wil85]. Telephones [KRS83]. Teleterminal [BT83]. Teleterminals [HAK83]. Television [RN82, Rze83, Sch83b, SNR80, Bul87]. Temperature [KKLS83]. Templates [RRW80]. temporal [RPS84]. Terabit [AKH87]. Term [Sze82]. Terminal [AFHK82, AFHS82, DRS83, Pik84, TNG88, BM87]. Terminals [Fla83, GLMS82, GOP87]. Termination [WPG87]. Terrestrial [San87]. Test [FS83a, KM83, Pfe80, CC85]. Testing [ADRW83, DRT82, Eig83, Gig80, Pra83, Rob82, Rub82, Uhr82, DTV85]. Tests [QEG83]. Text [BK83]. Text-Oriented [BK83]. Textons [JB83]. Textures [JB83]. TFM [Has81]. TFSP [CT83]. Their [EH82, Kum83, Leu86]. Theoretical [Eis84, Gri82, HM83]. Theory [AMS82, Cla83, Eic80, Lee83, LRS83, Lin84, MZ82, MS81, Pra83, AS84b, Fuc86, Log84c]. Thermal [Coy82, GNS^{+}81, Len85]. Thermally [BS83b]. Thermoelastic [JCV80]. Thin [AFPP82, RS88]. Thin-Film [AFPP82]. Three [Ber84, But81, Cox81, EHS83, Oza80, Rei81, NS85].

Three-Address [Rei81]. Three-Dimensional [But81, NS85]. Throughput [AS83]. Ticket [SA81a]. tie [MS84]. Tight [PLR82]. Time [AKK81, BBN82, BGGW83, CKKS88, Ein80, EY81, EH81, EH82, EH83b, EHS83, Fre82, GKS83, Has81, JCMQ83, MR81, MRR81, Rab82, Rum80, San82b, Sch83b, BTOA84, But84, Jag85b, Jua84a, RS85c, Rit84a, SJ85]. Time- [Rum80]. Time-Companded [EH82]. Time-Compressed [EY81]. Time-Compression [EY81, EH81, EH83b, EHS83]. Time-Frequency [Has81, Sch83b]. Time-Frequency-Coded [Ein80]. time-sharing [Rit84a]. Time-Varying [San82b]. Time-Warping [MR81, BTOA84]. Times [Jag82a, Fle84]. Timing [GM87, Ros83]. token [DeT84, WC85]. token-ring [WC85]. Tolerant [Wil82]. tomography [SS86]. Tone [BST81, Fav81, Feu81, Gad81, HKMP85]. Tool [RD82, EP84, KC88]. Tools [BW83a, BR82a, BGK^{+}86, CK88, GST86, Ker89]. Topics [Geh81]. topologies [Kam89b]. Total [AE83, BO83, BPRS83, BP83, BC83, BGGW83, EH83a, FGP^{+}83, GFGB83, HKRS83, HM83, Pfe83, Sch83a]. Totempole [LMM81]. Touch [BST81, EEF^{+}86]. touch-sensitive [EEF^{+}86]. Touch-Tone [BST81]. Tracing [Hol85]. Tracking [ABS82, GM87, LM81b]. Traffic [ADRW83, ARW80, BW83a, BDV83, BS83c, CHTW83, CS83, DVW83, DM83, GKS83, HHS83a, Hef80, Mor82, Mor80b, MW82, SW81, SC83, FM84, Hon84, Jag84, Whi84a]. Training [RBW82, RWT84, RWJ86]. Transaction [Fre80a, MY83]. Transaction-Oriented [Fre80a]. Transceiver [Gad81]. Transfer [ASG85, Gri82]. Transform [Jag82b, MF81, Jag85a, Log84c]. transformations [DS84]. transformer [DL84]. Transient [Car80, Gri82, Jag82a, Nuc81, WM88]. Transients [CM80]. Transmission [BBF81b, BNS82, CTM81, CMSP81, CMN83, EH81, GL81, GS83a, GS83b, JS82, Joh81, KFS83, LM81b, McN80, SSW84, SV81, TZD^{+}83, Wan83, AS84b, CCD^{+}84, DeT84, KS85, Pet84, Rog87, Sal85b, SWS84a]. Transmissions [EH82]. transmitted [SWS84a, SWS84b]. Transmitter [HCH^{+}83, KR80]. Transmitter-Receiver [HCH^{+}83]. Transmitting [WSX82]. transmultiplexers [BMD84]. Transponder [EHS83]. Transport [Hey83, Ran84, SKP84]. Transversal [GW81, FPQ^{+}84]. Traveling [BCHR83, HHS83b]. Traveling-Wave-Tube [BCHR83, HHS83b]. Treatment [BBF81b]. Tree [Str83]. Trees [FT83]. Trellis [CMS83, CS85]. trial [Bul87]. Triplication [Toy87]. Tristate [LMM81]. Trouble [AE83, RD82]. Troubleshooter [Mar88]. Truck [DFN81]. Trunk [Els80, GPP87, KRS82, KV84, Nea80, Sze80, Sze82]. Trunking [BC83]. TSPS [CRT82]. TTL [LMM81]. Tube [BCHR83, HHS83b]. Tuned [Kne80]. Turning [Koh83]. Tutorial [Ang81]. TV [EY81, EH81, EH82, EH83b, EHS83, Has81]. Two [FT83, GH82, Has81, Jay81b, JSC83, Kur80, Oza80, Pau83, RW81, WW80, CV85, Lo87, Sem85, WC85]. Two-Level [JSC83]. Two-Pass [RW81]. Two-Port [Kur80]. two-priority [Lo87]. Type [Hef82, Sun82, KLSW89]. types [FM84].

UHF [LS80b, Reu83]. ULSI [RS88].

Ultimate [Aca80]. Ultra [JCMS82, RS88]. Ultra-Thin [RS88]. Ultra-Wide [JCMS82]. Uncertainty [Noe81, Sze80]. Undersea [Chu83, CC85]. Unidirectional [LF82]. Unified [CKP88, MS81, Jua84a]. Union [Tur85]. Unit [RBH83]. Units [AGS82, DGH^{+}82, HCH^{+}83]. Universal [ABC89, Hei89]. UNIX [AT&T86, Aho84, Bac84, BHK^{+}84, Fed84, FMM84, FF88, Fra83, FHR84, Gin83, GM84, Ker89, Mac83, Mar84, PK84, RW84, Rit84a, Roc82, Wei82, BS88, Ram88, Ves88]. Update [Bre81]. Upper [AS83, CMS83]. Urban [CD80]. Usage [MM82, You84]. Use [BR82b, DG80, GB82, Kum83, MRR81, RLS84]. User [BF80, Yav82, BGS^{+}89, But84, Far89, IABB^{+}89]. user-interface [BGS^{+}89, Far89, IABB^{+}89]. Using [Aca81b, BS88, CP83a, CT83, EH81, EH82, GAB^{+}80, GC81, HSS82, HPU81, Len83, LMP82, LMP83, LS80c, LTT^{+}83, MR83, OOB83, Pha82, PKSG83, RRW80, RB82, SA81a, Sha81, SW82, Shi82, SV81, SNR80, Ben84, Gla84, Jua84b, RPS84, RS85a]. Utilities [EC83]. Utilization [Yum80]. Utilizing [GKS83].

V [AB89, BS88, CEGM89, DSMBB89, Dut89, FW88, JDW89, SP89]. V/MLS [FW88]. vacation [WC85]. Vapor [JDW89]. Variable [GS83a, Jay83, KSB88, Mey82, SA81b, Gha84]. Variable-Bit-Rate [GS83a, KSB88]. variable-length [Gha84]. Variables [LMOS83]. variants [KLSW89]. Variation [KRS82, Max82]. Variations [CP83a]. Varying [RWA81, San82b]. Vector [PCP80, RLS83, RSL83, SRJR87, RPS84, RSL84, RS85a]. Velocity [Tay84]. Velocity-saturated [Tay84]. Verification [ARW80, JCMS82]. Versions [Rob82]. Versus [Hal83, Toy87]. Very [BBF^{+}81a, Yav82]. Very-Large-Scale-Integration [BBF^{+}81a]. via [Ack84, Mey82]. Video [JS82, PC82, HB85, Pet84]. Videotex [Gog82]. view [Bul87, Jua84a]. Viewed [Fre80a]. Virtual [She88, WA88, Col88, Kum88, FHK^{+}84]. Virtual-Circuit [WA88, Kum88]. Vision [BJW88, JB83, BH86]. visual [Ray88]. Viterbi [Tur85]. VLSI [DC80, GST86, Kow88]. Vocabularies [RRWK82]. vocabulary [RLS84]. Voice [Ano82f, BHS82, BCW86, BBF81b, BKV82, CS82, CT83, CBB^{+}87, GKW82, KSB88, Max82, Nus82, RWR80, Wor82, CCD^{+}84, DeT84, HB85, Hon84, Jay87, ML84, PR86]. Voice-Controlled [RWR80]. Voice-Frequency [BBF81b]. Voice/Data [Max82, ML84]. voice/data/video [HB85]. voiceband [CCD^{+}84]. Voltage [Ung80]. Volterra [San82b]. Volume [KKLS83]. vowel [RS85a].

Wafer [dlRR86]. Waiting

[Fre82, Jag82a, Jag85b]. Walls [CMN83]. Warping [MR81, MRR81, Rab82, BTOA84, Jua84a]. Water [SB80]. Wave [BCHR83, HHS83b, HM80, DL84]. Waveforms [JC80]. Waveguide [CP83a, CP83b, Nas81, KD84]. Waveguides [Dra81b, LBHB83]. Wavelength [OOB83, SF82]. Way [Sal80, Sha88]. Weather [Sem80]. Weather-Cover [Sem80]. Weight [JCV83, Str83]. Weight-Gain [JCV83]. Weighting [SWS84b]. Which [Mor81b]. White [Ben80]. Wide [Blu80a, JCMS82, Sil84]. wide-angle [Sil84]. Wideband [MDSL82, Min84]. width [GP84]. window [LL88]. window-based [LL88]. Wire [OW84, WPG87]. Wire-tap [OW84]. wireless [KM85]. without [Fos85]. Word [CB83, DRM83, FLGD83, MR81, MRR81, RWA81, RW81, RRWK82, Rab82, RLS83, TRW82, BTOA84, Rab84, RPS84, RSL84, RWJ86, Ros84a, WRM84]. word-detection [WRM84]. Words [BR82b, RLS84]. Work [Aho82]. workbench [FF88, BF80, Fra83, Gin83, Mac83, MT80, O’N80, Tho80]. working [Lin84]. workproducts [Fow86]. Workstation [LFHV88, Dec88]. Workstations [BJW88, AAJ^{+}86]. Worst [Wit81]. Worst-Case [Wit81]. write [WWKZ84]. write-once [WWKZ84].

Writer [Fra83, Gin83, Mac83].

Year [LBP80]. Yields [Max82].

Zero [Aha82a, Nas81, Ber84, LF84]. Zero- [Aha82a]. zero-one [LF84]. zones [Pul88].

References

Ackenhusen:1986:SBG

[AAB^{+}86] John G. Ackenhusen, Syed S. Ali, David Bishop, Louis F. Rosa, and Reed Thorkildsen. Singleboard general-purpose speech recognition system. AT&T Technical Journal, 65(5):48– 59, September 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Aagesen:1981:DSP

[Aag81] J. Aagesen. Digital signal processor: Software simulator. The Bell System Technical Journal, 60(7):1475–1481, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-7-1475.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-7-1475.pdf.

Ackenhusen:1986:SPA

[AAJ^{+}86] J. G. Ackenhusen, S. S. Ali, J. G. Josenhans, J. W. Moffett, R. R. Robertson, and J. R. Tormos. Speech processing for AT&T workstations. AT&T Technical Journal, 65(5):60–67, September/October 1986. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Alles:1989:PTI

[AB89] David S. Alles and Kevin J. Brady. Packaging technology for III–V photonic devices and integrated circuits. AT&T Technical Journal, 68(1):83–92, January/February 1989. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Anderson:1981:NEM

[ABC^{+}81] T. W. Anderson, J. H. Bobsin, R. F. Cook, L. Gingerich, M. A. Marouf, and R. J. Milczarek. No. 4 ESS: Mass announcement subsystem. The Bell System Technical Journal, 60(6):1083–1108, July/ August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-6-1083.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-6-1083.pdf.

Austin:1989:UPC

[ABC89] Gary P. Austin, Barry S. Bosik, and Christopher J. Capece. Universal port concept. AT&T Technical Journal, 68(2):14–22, 1989. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Anderson:1988:RSM

[ABeB88] Kathryn Anderson, Roger P. Beck, and Thomas e. Buonanno. Reuse of software modules. AT&T Technical Journal, 67(4):71–76, July 1988. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Amster:1982:PCF

[ABS82] S. J. Amster, G. G. Brush, and B. Saperstein. Planning and conducting field-tracking studies. The Bell System Technical Journal, 61(9):2333–2365, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2333.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2333.pdf.

Acampora:1980:UCF

[Aca80]  A. S. Acampora.  The ul-

timate capacity of frequencyreuse communication satellites. The Bell System Technical Journal, 59(7):1089–1122, September 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-7-1089.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-7-1089.pdf.

Acampora:1981:AML

[Aca81a] A. S. Acampora. Analysis of maximum-likelihood sequence estimation performance for quadrature amplitude modulation. The Bell System Technical Journal, 60(6):865–885, July/ August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-6-865.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-6-865.pdf.

Acampora:1981:RMI

[Aca81b] A. S. Acampora. Rain margin improvement using resource sharing in 12-GHz satellite downlinks. The Bell System Technical Journal, 60(2):167–192, February 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-2-167.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-2-167.pdf.

Ackroyd:1984:SCF

[Ack84] Martin H. Ackroyd. Stationary and cyclostationary finite buffer behaviour computation via Levinson’s method. AT&T Bell Laboratories Technical Journal, 63(10 part 1):2159–2170, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Ackroyd:1985:NCD

[Ack85] M. H. Ackroyd. Numerical computation of delays in clocked schedules. AT&T Technical Journal, 64(2 part 1):617–631, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Al-Chalabi:1988:DBA

[ACL88] Muayyad Al-Chalabi and William J.

Liss. Design of the Bank of America California data network. AT&T Technical Journal, 67(6):87–106, November 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Ash:1981:DON

[ACM81] G. R. Ash, R. H. Cardwell, and R. P. Murray. Design and optimization of networks with dynamic routing. The Bell System Technical Journal, 60(8):1787–1820, October 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-8-1787.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-8-1787.pdf.

Abbott:1982:NRS

[ACSW82] N. B. Abbott, K. J. S. Chadha, D. P. Smith, and T. F. Wickham. No. 10A remote switching system: System overview. The Bell System Technical Journal, 61(4):391–417, April 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-4-391.
pdf;   http://www.alcatel-

lucent.com/bstj/vol61-1982/
articles/bstj61-4-391.pdf.

Adams:1983:ASS

[ADG^{+}83] R. L. Adams, J. L. Donoghue, A. N. Georgiades, J. R. Sundquist, R. E. Sheehey, and C. F. Walker. The AR6A single-sideband microwave radio system: Systems networks. The Bell System Technical Journal, 62(10):3465– 3475, December 1983. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol62/bstj62-10-3465.pdf;
http://www.alcatel-lucent.
com/bstj/vol62-1983/articles/
bstj62-10-3465.pdf.

Ahmari:1983:TSP

[ADRW83] R. Ahmari, R. S. DiPietro, S. C. Reed, and J. R. Williams. Traffic service position system no. 1B: Integration and system testing. The Bell System Technical Journal, 62(3):885–905, March 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-3-885.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-3-885.pdf.

Anderson:1983:TND

[AE83] D. R. Anderson and M. J. Evans. Total network data system: Performance measurement/trouble location. The Bell System Technical Journal, 62(7):2433–2458, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-7-2433.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-7-2433.pdf.

Albert:1982:DDC

[AFHK82] W. G. Albert, A. G. Favale, J. R. Hall, and D. H. Klockow. D4 digital channel bank family: Digital terminal physical design. The Bell System Technical Journal, 61(9):2757–2790, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2757.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2757.pdf.

Anderson:1982:NRS

[AFHS82] D. A. Anderson, D. R. Fuller, D. R. Hanson, and R. B. Schmidt. No. 10A remote switching system: Remote terminal firmware. The Bell System Technical Journal, 61(4):565–596, April 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-4-565.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-4-565.pdf.

Anderson:1987:LSC

[AFM87] Jerry M. Anderson, Dean R. Frey, and Calvin M. Miller. Lightwave splicing and connector technology. AT&T Technical Journal, 66(1):45–64, January 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Adams:1982:DDC

[AFPP82] R. L. Adams, J. S. Fisher, O. G. Petersen, and I. G. Post. D4 digital channel bank family: Thin-film dual active filter for pulse code modulation systems. The Bell System Technical Journal, 61(9):2815–2834, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-9-2815.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-9-2815.pdf.

Adrian:1982:NRS

[AFS82] J. M. Adrian, L. Freimanis, and R. G. Sparber. No. 10A remote switching system: Peripheral systems architecture and circuit design. The Bell System Technical Journal, 61(4):451–489, April 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-4-451.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-4-451.pdf.

Anderson:1984:DDD

[AG84] W. T. Anderson and P. F.

Glodis. Design diagrams for depressed cladding single-mode fibers. AT&T Bell Laboratories Technical Journal, 63(3):425– 430, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Agarwal:1989:ADS

[Aga89] Yogesh K. Agarwal. An algorithm for designing survivable networks. AT&T Technical Journal, 68(3):64–76, 1989. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Aprille:1982:DDCb

[AGS82] T. J. Aprille, D. V. Gupta, and P. G. St. Amand. D4 digital channel bank family: Dataport — channel units for digital data system subrates. The Bell System Technical Journal, 61(9):2721–2740, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2721.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2721.pdf.

Amster:1986:SMR

[AH86] Sigmund J. Amster and Jeffrey H. Hopper. Statistical methods for reliability improvement. AT&T Technical Journal, 65(2): 69–76, March/April 1986. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Ahamed:1980:BJB

[Aha80] S. V. Ahamed. B.S.T.J. briefs: Serial coding for cyclic block codes. The Bell System Technical Journal, 59(2):269–276, February 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-2-269.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-2-269.pdf.

Ahamed:1982:GCZ

[Aha82a] S. V. Ahamed. A general class of zero- or minimumdelay fractional rate change circuits. The Bell System Technical Journal, 61(3):327–346, March 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-3-327.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-3-327.pdf.

Ahamed:1982:SDS

[Aha82b] S. V. Ahamed. Simulation and design studies of digital subscriber lines. The Bell System Technical Journal, 61(6):1003–1077, July/ August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

images/Vol61/bstj61-6-1003.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-6-1003.pdf.

Ambekar:1987:SP

[AHC87] S. M. Ambekar, W. E. Hamilton, and T. E. Cole. Systems packaging. AT&T Technical Journal, 66(4):81–95, 1987. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Aho:1982:DSD

[Aho82] A. V. Aho. Database systems: Database work at Bell Laboratories. The Bell System Technical Journal, 61(9): 2383–2385, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2383.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2383.pdf.

Aho:1984:USF

[Aho84] A. V. Aho. The UNIX system: Foreword. AT&T Bell Laboratories Technical Journal, 63(8 part 2):1573–1576, October 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Astmann:1988:PAS

[AK88] Robert H. Astmann and Joseph S.

Kaufman. Performance analysis of System 75. AT&T Technical Journal, 67(5):110–120, September/October 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Acampora:1987:TLN

[AKH87] Anthony S. Acampora, Mark J. Karol, and Michael G. Hluchyj. Terabit lightwave networks: The multihop approach. AT&T Technical Journal, 66(6):21– 34, November 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Akinpelu:1984:OPE

[Aki84] Jacqueline M. Akinpelu. The overload performance of engineered networks with nonhierarchical and hierarchical routing. AT&T Bell Laboratories Technical Journal, 63(7):1261– 1281, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Ash:1981:SRT

[AKK81] G. R. Ash, A. H. Kafker, and K. R. Krishnan. Servicing and real-time control of networks with dynamic routing. The Bell System Technical Journal, 60(8):1821–1845, October 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-8-1821.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-8-1821.pdf.

Albanese:1982:FSN

[Alb82] A. Albanese. Fail-safe nodes for lightguide digital networks. The Bell System Technical Journal, 61(2):247–256, February 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-2-247.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-2-247.pdf.

Albanese:1983:SNC

[Alb83]  A. Albanese.  Star network

with collision-avoidance circuits. The Bell System Technical Journal, 62(3):631–638, March 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-3-631.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-3-631.pdf.

Aldermeshian:1986:ISE

[Ald86] Hrair Aldermeshian. ISDN standards evolution. AT&T Technical Journal, 65(1):19–26, January/February 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Aldefeld:1980:MDS

[ALS80] B. Aldefeld, S. E. Levinson, and T. G. Szymanski. A minimum-distance search technique and its application to automatic directory assistance.

The Bell System Technical Journal, 59(8):1343–1356, October 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-8-1343.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-8-1343.pdf.

Andreassen:1988:ISO

[ALS88] Alf L. Andreassen, William J. Leighton III, and David F. Schreiber. Information security: An overview. AT&T Technical Journal, 67(3):2–8, May 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Aveyard:1980:SRC

[AM80] R. L. Aveyard and F. T. Man. A study on the reliability of the circuit maintenance system-1B. The Bell System Technical Journal, 59(8):1317–1332, October 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-8-1317.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-8-1317.pdf.

Andrews:1982:SPC

[AM82] F. T. Andrews, Jr. and K. E. Martersteck. Stored program controlled network: Prologue. The Bell System Technical Journal, 61(7):1575–1577, September 1982. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-7-1575.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-7-1575.pdf.

Anick:1982:STD

[AMS82] D. Anick, D. Mitra, and M. M. Sondhi. Stochastic theory of a data-handling system with multiple sources. The Bell System Technical Journal, 61(8):1871–1894, October 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-8-1871.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-8-1871.pdf.

Andrews:1983:LMS

[And83] J. M. Andrews. A lithographic mask system for MOS fineline process development. The Bell System Technical Journal, 62(4):1107–1160, April 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-4-1107.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-4-1107.pdf.

Andersson:1988:BFI

[And88] Russell L. Andersson. Building fast, intelligent robot systems. AT&T Technical Journal, 67(2):73–86, March/April 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Angelo:1981:DSP

[Ang81] E. J. Angelo, Jr. Digital signal processor: A tutorial introduction to digital filtering. The Bell System Technical Journal, 60(7):1499–1546, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-7-1499.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-7-1499.pdf.

Anonymous:1980:CIa

[Ano80a] Anonymous. Contributors to this issue. The Bell System Technical Journal, 59(1):139–141, January 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-1-139.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-1-139.pdf.

Anonymous:1980:CIb

[Ano80b] Anonymous. Contributors to this issue. The Bell System Technical Journal, 59(2):259–263, February 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-2-259.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol59-1980/
articles/bstj59-2-259.pdf.

Anonymous:1980:CIc

[Ano80c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 59(3):463–466, March 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-3-463.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-3-463.pdf.

Anonymous:1980:CId

[Ano80d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 59(4):657–659, April 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-4-657.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-4-657.pdf.

Anonymous:1980:CIe

[Ano80e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 59(5):797–798, May/ June 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-5-797.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-5-797.pdf.

Anonymous:1980:CIf

[Ano80f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 59(6):1073–1077, July/ August 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-6-1073.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-6-1073.pdf.

Anonymous:1980:CIg

[Ano80g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 59(7): 1277–1281, September 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-7-1277.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-7-1277.pdf.

Anonymous:1980:CIh

[Ano80h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 59(8):1525–1531, October 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-8-1525.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-8-1525.pdf.

Anonymous:1980:CIi

[Ano80i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 59(9): 1827–1832, November 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-9-1827.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-9-1827.pdf.

Anonymous:1980:CIj

[Ano80j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 59(10):2001– 2002, December 1980. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol59/bstj59-10-2001.pdf;
http://www.alcatel-lucent.
com/bstj/vol59-1980/articles/
bstj59-10-2001.pdf.

Anonymous:1980:PBLa

[Ano80k] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 59(1):143–147, January 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-1-143.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-1-143.pdf.

Anonymous:1980:PBLb

[Ano80l] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 59(2):265–267, February 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-2-265.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-2-265.pdf.

Anonymous:1980:PBLc

[Ano80m] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 59(3):467–470, March 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-3-467.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-3-467.pdf.

Anonymous:1980:PBLd

[Ano80n] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 59(4):661–667, April 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-4-661.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-4-661.pdf.

Anonymous:1980:PBLe

[Ano80o] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 59(5):799–803, May/June 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-5-799.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-5-799.pdf.

Anonymous:1980:PBLf

[Ano80p] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 59(6):1079–1082, July/ August 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-6-1079.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-6-1079.pdf.

Anonymous:1980:PBLg

[Ano80q] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 59(7):1283–1285, September 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-7-1283.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-7-1283.pdf.

Anonymous:1980:PBLh

[Ano80r] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 59(8):1533–1537, October 1980. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-8-1533.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-8-1533.pdf.

Anonymous:1980:PBLi

[Ano80s] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 59(9):1833–1835, November 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-9-1833.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-9-1833.pdf.

Anonymous:1980:PBLj

[Ano80t] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 59(10): 2003–2004, December 1980. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol59/bstj59-10-2003.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol59-1980/articles/
                                        bstj59-10-2003.pdf.

Anonymous:1981:AAa

[Ano81a] Anonymous. Acronyms and abbreviations. The Bell System Technical Journal, 60(6): 1225–1228, July/August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-6-1225.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-6-1225.pdf.

Anonymous:1981:AAb

[Ano81b] Anonymous. Acronyms and abbreviations. The Bell System Technical Journal, 60(7): 1699–1701, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-7-1699.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-7-1699.pdf.

Anonymous:1981:CIa

[Ano81c] Anonymous. Contributors to this issue. The Bell System Technical Journal, 60(1):117–118, January 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-1-117.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-1-117.pdf.

Anonymous:1981:CIb

[Ano81d] Anonymous. Contributors to this issue. The Bell System Technical Journal, 60(2):297–298, February 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-2-297.

                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-2-297.pdf.

Anonymous:1981:CIc

[Ano81e] Anonymous. Contributors to this issue. The Bell System Technical Journal, 60(3):445–447, March 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-3-445.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-3-445.pdf.

Anonymous:1981:CId

[Ano81f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 60(4):575–578, April 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-4-575.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-4-575.pdf.

Anonymous:1981:CIe

[Ano81g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 60(5):775–778, May/ June 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-5-775.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-5-775.pdf.

Anonymous:1981:CIf

[Ano81h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 60(6):1027–1033, July/ August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-6-1027.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-6-1027.pdf.

Anonymous:1981:CIg

[Ano81i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 60(6):1229–1232, July/ August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-6-1229.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-6-1229.pdf.

Anonymous:1981:CIh

[Ano81j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 60(7): 1423–1426, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-7-1423.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-7-1423.pdf.

Anonymous:1981:CIi

[Ano81k] Anonymous. Contributors to this issue. The Bell System Technical Journal, 60(7): 1703–1709, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-7-1703.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-7-1703.pdf.

Anonymous:1981:CIj

[Ano81l] Anonymous. Contributors to this issue. The Bell System Technical Journal, 60(8):1985–1991, October 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-8-1985.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-8-1985.pdf.

Anonymous:1981:CIk

[Ano81m] Anonymous. Contributors to this issue. The Bell System Technical Journal, 60(9): 2273–2278, November 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-9-2273.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-9-2273.pdf.

Anonymous:1981:CIl

[Ano81n] Anonymous. Contributors to this issue. The Bell System Technical Journal, 60(10):2421– 2423, December 1981. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol60/bstj60-10-2421.pdf;
http://www.alcatel-lucent.
com/bstj/vol60-1981/articles/
bstj60-10-2421.pdf.

Anonymous:1981:PBLa

[Ano81o] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 60(1):119–120, January 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-1-119.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-1-119.pdf.

Anonymous:1981:PBLb

[Ano81p] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 60(2):299–300, February 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-2-299.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-2-299.pdf.

Anonymous:1981:PBLc

[Ano81q] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 60(3):449–453, March 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-3-449.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-3-449.pdf.

Anonymous:1981:PBLd

[Ano81r] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 60(4):579–581, April 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-4-579.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-4-579.pdf.

Anonymous:1981:PBLe

[Ano81s] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 60(5):779–782, May/June 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-5-779.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-5-779.pdf.

Anonymous:1981:PBLf

[Ano81t] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 60(6):1035–1038, July/ August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-6-1035.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-6-1035.pdf.

Anonymous:1981:PBLg

[Ano81u] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 60(7):1427–1429, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-7-1427.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-7-1427.pdf.

Anonymous:1981:PBLh

[Ano81v] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 60(8):1993–??, October 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-8-1993.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-8-1993.pdf.

Anonymous:1981:PBLi

[Ano81w] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 60(9):2279–2280, November 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-9-2279.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-9-2279.pdf.

Anonymous:1981:PBLj

[Ano81x] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 60(10): 2425–2426, December 1981. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol60/bstj60-10-2425.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol60-1981/articles/
                                        bstj60-10-2425.pdf.

Anonymous:1982:AAa

[Ano82a] Anonymous. Acronyms and abbreviations. The Bell System Technical Journal, 61(4): 643–646, April 1982. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol61/bstj61-4-643.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol61-1982/articles/
                                        bstj61-4-643.pdf.

Anonymous:1982:AAb

[Ano82b] Anonymous. Acronyms and abbreviations. The Bell System Technical Journal, 61(6): 1345–1347, July/August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol61/bstj61-6-1345.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-6-1345.pdf.

Anonymous:1982:AAc

[Ano82c] Anonymous. Acronyms and abbreviations. The Bell System Technical Journal, 61(7): 1799–1803, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-7-1799.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-7-1799.pdf.

Anonymous:1982:AAd

[Ano82d] Anonymous. Acronyms and abbreviations. The Bell System Technical Journal, 61(9): 2597–2598, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2597.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2597.pdf.

Anonymous:1982:AAe

[Ano82e] Anonymous. Acronyms and abbreviations. The Bell System Technical Journal, 61(9): 2835–2838, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2835.

                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-9-2835.pdf.

Anonymous:1982:AAV

[Ano82f] Anonymous. Acronyms and abbreviations for voice storage system. The Bell System Technical Journal, 61(5):913–914, May/ June 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-5-913.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-5-913.pdf.

Anonymous:1982:CIa

[Ano82g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 61(1):117–118, January 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-1-117.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-1-117.pdf.

Anonymous:1982:CIb

[Ano82h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 61(2):257–258, February 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-2-257.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-2-257.pdf.

Anonymous:1982:CIc

[Ano82i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 61(3):377–378, March 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-3-377.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-3-377.pdf.

Anonymous:1982:CId

[Ano82j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 61(4):647–652, April 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-4-647.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-4-647.pdf.

Anonymous:1982:CIe

[Ano82k] Anonymous. Contributors to this issue. The Bell System Technical Journal, 61(5):915–921, May/ June 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-5-915.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-5-915.pdf.

Anonymous:1982:CIf

[Ano82l] Anonymous. Contributors to this issue. The Bell System Technical Journal, 61(6):1079–1081, July/ August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-6-1079.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-6-1079.pdf.

Anonymous:1982:CIg

[Ano82m] Anonymous. Contributors to this issue. The Bell System Technical Journal, 61(6):1349–1357, July/ August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-6-1349.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-6-1349.pdf.

Anonymous:1982:CIh

[Ano82n] Anonymous. Contributors to this issue. The Bell System Technical Journal, 61(7): 1551–1554, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-7-1551.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-7-1551.pdf.

Anonymous:1982:CIi

[Ano82o] Anonymous. Contributors to this issue. The Bell System Technical Journal, 61(7): 1805–1815, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-7-1805.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-7-1805.pdf.

Anonymous:1982:CIj

[Ano82p] Anonymous. Contributors to this issue. The Bell System Technical Journal, 61(8):2109–2115, October 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-8-2109.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-8-2109.pdf.

Anonymous:1982:CIk

[Ano82q] Anonymous. Contributors to this issue. The Bell System Technical Journal, 61(9): 2367–2372, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2367.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2367.pdf.

Anonymous:1982:CIl

[Ano82r] Anonymous. Contributors to this issue. The Bell System Technical Journal, 61(9): 2599–2603, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-9-2599.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-9-2599.pdf.

Anonymous:1982:CIm

[Ano82s] Anonymous. Contributors to this issue. The Bell System Technical Journal, 61(9): 2839–2847, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-9-2839.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-9-2839.pdf.

Anonymous:1982:CIn

[Ano82t] Anonymous. Contributors to this issue. The Bell System Technical Journal, 61(10):3041– 3045, December 1982. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol61/bstj61-10-3041.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol61-1982/articles/
                                        bstj61-10-3041.pdf.

Anonymous:1982:LE

[Ano82u] Anonymous. Letter to the editor. The Bell System Technical Journal, 61(3):375–??, March 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-3-375.

pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-3-375.pdf.

Anonymous:1982:PBLa

[Ano82v] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 61(1):119–120, January 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-1-119.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-1-119.pdf.

Anonymous:1982:PBLb

[Ano82w] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 61(2):259–260, February 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-2-259.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-2-259.pdf.

Anonymous:1982:PBLc

[Ano82x] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 61(3):379–381, March 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-3-379.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-3-379.pdf.

Anonymous:1982:PBLd

[Ano82y] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 61(4):653–657, April 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-4-653.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-4-653.pdf.

Anonymous:1982:PBLe

[Ano82z] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 61(5):923–928, May/June 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-5-923.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-5-923.pdf.

Anonymous:1982:PBLf

[Ano82-27] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 61(6):1083–1091, July/ August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-6-1083.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol61-1982/
articles/bstj61-6-1083.pdf.

Anonymous:1982:PBLg

[Ano82-28] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 61(7):1555–1564, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-7-1555.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-7-1555.pdf.

Anonymous:1982:PBLh

[Ano82-29] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 61(8):2117–2124, October 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-8-2117.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-8-2117.pdf.

Anonymous:1982:PBLi

[Ano82-30] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 61(9):2373–2379, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2373.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-9-2373.pdf.

Anonymous:1982:PBLj

[Ano82-31] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 61(10): 3047–3049, December 1982. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol61/bstj61-10-3047.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol61-1982/articles/
                                        bstj61-10-3047.pdf.

Anonymous:1983:AAa

[Ano83a] Anonymous. Acronyms and abbreviations. The Bell System Technical Journal, 62 (1):411–416, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-1-411.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-1-411.pdf.

Anonymous:1983:AAb

[Ano83b] Anonymous. Acronyms and abbreviations. The Bell System Technical Journal, 62(3): 979–983, March 1983. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol62/bstj62-3-979.pdf;
                                        http://www.alcatel-lucent.

com/bstj/vol62-1983/articles/
bstj62-3-979.pdf.

Anonymous:1983:AAc

[Ano83c] Anonymous. Acronyms and abbreviations. The Bell System Technical Journal, 62(6): 1551–1552, July/August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1551.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1551.pdf.

Anonymous:1983:AAd

[Ano83d] Anonymous. Acronyms and abbreviations. The Bell System Technical Journal, 62(7): 2473–2478, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-7-2473.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-7-2473.pdf.

Anonymous:1983:AAe

[Ano83e] Anonymous. Acronyms and abbreviations. The Bell System Technical Journal, 62(10):3491– 3492, December 1983. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol62/bstj62-10-3491.pdf;
http://www.alcatel-lucent.

                                        com/bstj/vol62-1983/articles/
                                        bstj62-10-3491.pdf.

Anonymous:1983:CIa

[Ano83f] Anonymous. Contributors to this issue. The Bell System Technical Journal, 62(1):153–157, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-1-153.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-1-153.pdf.

Anonymous:1983:CIb

[Ano83g] Anonymous. Contributors to this issue. The Bell System Technical Journal, 62(1):417–428, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-1-417.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-1-417.pdf.

Anonymous:1983:CIc

[Ano83h] Anonymous. Contributors to this issue. The Bell System Technical Journal, 62(2):583–586, February 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-2-583.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-2-583.pdf.

Anonymous:1983:CId

[Ano83i] Anonymous. Contributors to this issue. The Bell System Technical Journal, 62(3):743–746, March 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-3-743.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-3-743.pdf.

Anonymous:1983:CIe

[Ano83j] Anonymous. Contributors to this issue. The Bell System Technical Journal, 62(3):985–992, March 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-3-985.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-3-985.pdf.

Anonymous:1983:CIf

[Ano83k] Anonymous. Contributors to this issue. The Bell System Technical Journal, 62(4):1169–1171, April 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-4-1169.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-4-1169.pdf.

Anonymous:1983:CIg

[Ano83l] Anonymous. Contributors to this issue. The Bell System Technical Journal, 62(5):1337–1340, May/ June 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-5-1337.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-5-1337.pdf.

Anonymous:1983:HM

[Ano83m] Anonymous. History and methods. The Bell System Technical Journal, 62(6):1569, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1569.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1569.pdf.

Anonymous:1983:HFBa

[Ano83n] Anonymous. Human factors and behavioral science: Characteristics of human performance. The Bell System Technical Journal, 62(6):1617, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1617.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1617.pdf.

Anonymous:1983:HFBc

[Ano83o] Anonymous. Human factors and behavioral science: Interface design. The Bell System Technical Journal, 62(6):1751, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1751.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1751.pdf.

Anonymous:1983:HFBd

[Ano83p] Anonymous. Human factors and behavioral science: New functions for technology. The Bell System Technical Journal, 62(6):1881, July/August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1881.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1881.pdf.

Anonymous:1983:HFBb

[Ano83q] Anonymous. Human factors and behavioral science: New technological demands. The Bell System Technical Journal, 62(6):1699, July/August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1699.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1699.pdf.

Anonymous:1983:LE

[Ano83r] Anonymous. Letter to the editor.

The Bell System Technical Journal, 62(7):2113–2114, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-7-2113.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-7-2113.pdf.

Anonymous:1983:PBLa

[Ano83s] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 62(1):159–164, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-1-159.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-1-159.pdf.

Anonymous:1983:PBLb

[Ano83t] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 62(2):587–591, February 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-2-587.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-2-587.pdf.

Anonymous:1983:PBLc

[Ano83u] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 62(3):747–750, March 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-3-747.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-3-747.pdf.

Anonymous:1983:PBLd

[Ano83v] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 62(4):1173–1177, April 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-4-1173.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-4-1173.pdf.

Anonymous:1983:PBLe

[Ano83w] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 62(5):1341–1347, May/ June 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-5-1341.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-5-1341.pdf.

Anonymous:1983:PBLf

[Ano83x] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 62(6):1553–1556, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-6-1553.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-6-1553.pdf.

Anonymous:1983:PBLg

[Ano83y] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 62(7):2115–2118, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-7-2115.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-7-2115.pdf.

Anonymous:1983:PBLh

[Ano83z] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 62(8):2647–2659, October 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-8-2647.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-8-2647.pdf.

Anonymous:1983:PBLi

[Ano83-27] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 62(9):2845–2847, November 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-9-2845.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-9-2845.pdf.

Anonymous:1983:PBLj

[Ano83-28] Anonymous. Papers by Bell Laboratories authors. The Bell System Technical Journal, 62(10): 3063–3069, December 1983. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol62/bstj62-10-3063.pdf;
http://www.alcatel-lucent.
com/bstj/vol62-1983/articles/
bstj62-10-3063.pdf.

Aprille:1982:DDCa

[ANSW82] T. J. Aprille, S. Narayanan, P. G. St. Amand, and F. E. Weber. D4 digital channel bank family: Dataport — digital access through D4. The Bell System Technical Journal, 61(9):2703–2720, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2703.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2703.pdf.

Ackenhusen:1985:SCI

[AO85] J. G. Ackenhusen and Y. H.

Oh. Single-chip implementation of feature measurement for LPCbased speech recognition. AT&T Technical Journal, 64(8):1787– 1805, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Atal:1986:SRD

[AR86] Bishnu S. Atal and Lawrence R. Rabiner. Speech research directions. AT&T Technical Journal, 65(5):75–88, September 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Atkins:1980:TSP

[ARW80] J. Atkins, K. A. Raschke, and D. L. Woody. Traffic service position system no. 1: Busy line verification feature. The Bell System Technical Journal, 59(8):1397–1416, October 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-8-1397.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-8-1397.pdf.

Arthurs:1983:ULB

[AS83] E. Arthurs and B. W. Stuck. Upper and lower bounds on mean throughput rate and mean delay in memory-constrained queueing networks. The Bell System Technical Journal, 62 (2):541–581, February 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-2-541.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol62-1983/
articles/bstj62-2-541.pdf.

Abele:1984:SSL

[AS84a] Thomas A. Abele and Albert J.

Schepis. SLC 96 subscriber loop carrier system. AT&T Bell Laboratories Technical Journal, 63 (10 part 2):2273–2281, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Amitay:1984:LET

[AS84b] Noach Amitay and Jack Salz.

Linear equalization theory in digital data transmission over dually polarized fading radio channels. AT&T Bell Laboratories Technical Journal, 63(10 part 1):2215– 2259, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Aggarwal:1985:NFT

[ASG85] Sudhir Aggarwal, Krishan Sabnani, and B. Gopinath. New File Transfer Protocol. AT&T Technical Journal, 64(10):2387– 2411, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

ATT:1986:AUS

[AT&T86] AT&T. AT&T UNIX System Readings and Applications, volume II. Prentice-Hall, Inc., Upper Saddle River, NJ 07458, USA, 1986. ISBN 0-13-939845-7. xii + 324 pp. LCCN QA76.76.O63 U553 1986.

Bach:1984:MUO

[Bac84] M. J. Bach. Multiprocessor UNIX operating systems. AT&T Bell Laboratories Technical Journal, 63(8 part 2):1733–1749, October 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Batorsky:1984:BSN

[Bat84] D. V. Batorsky. 1980 Bell System noise survey of the loop plant. AT&T Bell Laboratories Technical Journal, 63(5):775–818, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Biffle:1987:NSA

[BB87] Johnny H. Biffle and Steven N. Burchett. Nonlinear structural analysis of complex electronic and electromechanical assemblies. AT&T Technical Journal, 66(6):51–62, November 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Bohn:1984:SSL

[BBB^{+}84] Peter P. Bohn, Charles A. Brackett, Michael J. Buckler, Tadikonda N. Rao, and Robert H. Saul. SLC 96 subscriber loop carrier system: the fiber SLC carrier system. AT&T Bell Laboratories Technical Journal, 63(10 part 2):2389–2416, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Baxter:1985:SCC

[BBB^{+}85] L. A. Baxter, P. R. Berkowitz, C. A. Buzzard, J. J. Horenkamp, and F. E. Wyatt. System 75: communications and control architecture. AT&T Technical Journal, 64(1 part 1):153– 173, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Basso:1989:OSA

[BBD^{+}89] Richard J. Basso, Hugh J.

Beuscher, Iris S. Dowden, Richard J. Piereth, and Samuel M. Salchenberger. OSPS system architecture. AT&T Technical Journal, 68(6):9–24, 1989. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Benvenuto:1986:KAC

[BBDS86] Nevio Benvenuto, Guido Bertocci, William R. Daumer, and D. K. Sparrell. The 32-kb/s AD-PCM coding standard. AT&T Technical Journal, 65(5):12–22, September/October 1986. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Barber:1981:DSP

[BBF^{+}81a] F. E. Barber, T. J. Bartoli, R. L.

Freyman, J. A. Grant, J. Kane, and R. N. Kershaw. Digital signal processor: An overview of the silicon very-large-scale-integration implementation. The Bell System Technical Journal, 60(7):1441–1447, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-7-1441.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-7-1441.pdf.

Blake:1981:DSP

[BBF81b] R. B. Blake, A. C. Bolling, and R. L. Farah. Digital signal processor: Voice-frequency transmission treatment for specialservice telephone circuits. The Bell System Technical Journal, 60(7):1585–1619, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-7-1585.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-7-1585.pdf.

Braunstein:1987:ACD

[BBM87] Marcy R. Braunstein, Christopher L. Burton, and Sandra D. McNabb. ASQIC 800 call data master. AT&T Technical Journal, 66(3):21–31, May/ June 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Barclay:1982:DSR

[BBN82] D. K. Barclay, E. R. Byrne, and F. K. Ng. Database systems: A real-time database management system for no. 5 ESS. The Bell System Technical Journal, 61(9):2423–2437, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2423.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2423.pdf.

Boggs:1982:ARSa

[BBOV82] P. S. Boggs, M. W. Bowker, E. A. Overstreet, and R. W. Vetter, Jr. Automated repair service bureau: Evolution. The Bell System Technical Journal, 61(6):1097–1114, July/ August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-6-1097.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-6-1097.pdf.

Basinger:1982:SPC

[BBP^{+}82] R. G. Basinger, M. Berger, E. M.

Prell, V. L. Ransom, and J. R. Williams. Stored program controlled network: Calling card service — overall description and operational characteristics. The Bell System Technical Journal, 61(7):1655–1673, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-7-1655.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-7-1655.pdf.

Belanger:1988:SRD

[BBW88] David G. Belanger, G. David Bergland, and Mike Wish. Some research directions for large-scale software development. AT&T Technical Journal, 67(4):77–92, July 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Bezdek:1983:TND

[BC83] P. V. Bezdek and J. P. Collins. Total network data system: Trunking system. The Bell System Technical Journal, 62(7):2345–2364, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-7-2345.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-7-2345.pdf.

Becker:1985:DSD

[BC85] R. A. Becker and J. M. Chambers. Design of the S system for data analysis. AT&T Technical Journal, 64(9):2131–2151, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Boza:1986:CIA

[BCH^{+}86] L. B. Boza, C. D. Chavous, J. C. Hsu, R. D. Lake, and M. K. Pratt. CIMA — an integrated architecture for product realization. AT&T Technical Journal, 65(4):17–24, July/August 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Balicki:1983:ASS

[BCHR83] J. F. Balicki, E. F. Cook, R. C. Heidt, and V. E. Rutter. The AR6A single-sideband microwave radio system: The traveling-wave-tube amplifier. The Bell System Technical Journal, 62(10):3429–3445, December 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol62/bstj62-10-3429.pdf;
http://www.alcatel-lucent.
com/bstj/vol62-1983/articles/
bstj62-10-3429.pdf.

Brown:1988:SSP

[BCMZ88] Donald W. Brown, Christopher D. Carson, Warren A. Montgomery, and Paul M. Zislis. Software specification and prototyping technologies. AT&T Technical Journal, 67(4):33–45, July 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Birnbaum:1986:VPS

[BCW86] Martha Birnbaum, Larry A. Cohen, and Frank X. Welsh. Voice password system for access security. AT&T Technical Journal, 65(5):68–74, September 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Boddie:1981:DSPb

[BDE^{+}81] J. R. Boddie, G. T. Daryanani, I. I. Eldumiati, R. N. Gadenz, J. S. Thompson, and S. M. Walters. Digital signal processor: Architecture and performance. The Bell System Technical Journal, 60(7):1449–1462, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-7-1449.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-7-1449.pdf.

Barrere:1982:SPC

[BDL^{+}82a] R. G. Barrere, J. P. Delatore, J. W. Lurtz, R. J. Piereth, and C. M. Rubald. Stored program controlled network: Data base administration system — overall description and operational characteristics. The Bell System Technical Journal, 61(7):1759–1777, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-7-1759.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-7-1759.pdf.

Brown:1982:NRS

[BDL82b] D. W. Brown, J. J. Driscoll, and F. M. Lax. No. 10A remote switching system: Host software. The Bell System Technical Journal, 61(4):491–524, April 1982. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-4-491.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-4-491.pdf.

Brown:1983:MSE

[BDL^{+}83] D. W. Brown, J. J. Driscoll, F. M. Lax, M. W. Saad, and J. G. Whitemyer. Modernization of the suburban ESS: Hosting the no. 10A remote switching system. The Bell System Technical Journal, 62(6):1497–1536, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1497.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1497.pdf.

Bodnar:1983:TSP

[BDV83] J. J. Bodnar, J. R. Daino, and K. A. VanderMeulen. Traffic service position system no. 1B: Switching control center system interface. The Bell System Technical Journal, 62(3):941–957, March 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-3-941.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-3-941.pdf.

Benes:1980:BJB

[Ben80] V. E. Beneš. B.S.T.J. briefs:

Least-squares estimator for frequency-shift position modulation in white noise. The Bell System Technical Journal, 59(7):1289–1296, September 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-7-1289.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-7-1289.pdf.

Benes:1981:BSC

[Ben81a] V. E. Beneš. Blocking states in connecting networks made of square switches arranged in stages. The Bell System Technical Journal, 60(4):511–521, April 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-4-511.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-4-511.pdf.

Benes:1981:DOE

[Ben81b] V. E. Beneš. Detecting the occurrence of an event by FM through noise. The Bell System Technical Journal, 60(9):2227–2233, November 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-9-2227.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-9-2227.pdf.

Benes:1983:GCP

[Ben83] V. E. Beneš. Growth, complexity, and performance of telephone connecting networks. The Bell System Technical Journal, 62(2):499–539, February 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-2-499.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-2-499.pdf.

Benvenuto:1984:DAM

[Ben84] N. Benvenuto. Distortion analysis on measuring the impulse response of a system using a crosscorrelation. AT&T Bell Laboratories Technical Journal, 63(10 part 1):2171–2192, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Bergmann:1982:EPH

[Ber82]  E. E. Bergmann.  Electro-

magnetic propagation in homogeneous media with Hermitian permeability and permittivity. The Bell System Technical Journal, 61(6):935–948, July/ August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-6-935.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-6-935.pdf.

Bernstein:1984:TDA

[Ber84] Jeremy Bernstein. Three degrees above zero: Bell Labs in the information age. C. Scribner’s, New York, NY, USA, 1984. ISBN 0-684-18170-3. xiii + 241 + 8 pp. LCCN T178.B45 B47 1984. US$17.95.

Breiland:1980:DWU

[BF80] J. R. Breiland and R. A. Friedenson. Designer’s Workbench: The user environment. The Bell System Technical Journal, 59(9):1767–1792, November 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-9-1767.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-9-1767.pdf.

Burman:1986:PAT

[BGGN^{+}86] D. Y. Burman, F. J. Gurrola-Gal, A. Nozari, S. Sathaye, and J. P. Sitarik. Performance analysis techniques for IC manufacturing lines. AT&T Technical Journal, 65(4):46–57, July/August 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Byrne:1983:TND

[BGGW83] C. J. Byrne, D. J. Gagne, J. A. Grandle, Jr., and G. H. Wedemeyer. Total network data system: Data acquisition and near-real-time surveillance by EADAS. The Bell System Technical Journal, 62 (7):2209–2237, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-7-2209.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-7-2209.pdf.

Boddie:1986:DDS

[BGK^{+}86] James R. Boddie, Renato N.

Gadenz, Robert N. Kershaw, W. Patrick Hays, and James Tow. The DSP32 digital signal processor and its application development tools. AT&T Technical Journal, 65(5):89– 104, September 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Bennett:1989:STC

[BGS^{+}89] Raymond W. Bennett, Steven L.

Greenspan, Ann K. Syrdal, Judith E. Tschirgi, and John J. Wisowaty. Speaking to, from, and through computers. speech technologies and user-interface design. AT&T Technical Journal, 68(5):17–30, September/October 1989. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Bainbridge:1984:SSL

[BH84a] Gary D. Bainbridge and Robert W.

Henn. SLC 96 subscriber loop carrier system: physical design. AT&T Bell Laboratories Technical Journal, 63(10 part 2):2333– 2362, 1984. CODEN ABLJER.

ISSN 0748-612X (print), 2376-7162 (electronic).

Barta:1984:ADA

[BH84b] S. M. Barta and M. L. Honig. Analysis of a demand assignment TDMA blocking system. AT&T Bell Laboratories Technical Journal, 63(1):89–114, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Budrikis:1984:MCD

[BH84c] Z. L. Budrikis and M. Hatamian. Moment calculations by digital filters. AT&T Bell Laboratories Technical Journal, 63(2):217– 229, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Brunelle:1986:LSV

[BH86] Norman R. Brunelle and Frank P. Higgins. Line scan vision system. AT&T Technical Journal, 65(4):58–65, July/August 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Brachman:1988:EAI

[BH88] R. J. Brachman and F. H. Henig. The emergence of artificial intelligence technology. AT&T Technical Journal, 67(1):3–6, January 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Bodenstab:1984:UOS

[BHK^{+}84] D. E. Bodenstab, Thomas F. Houghton, Keith A. Kelleman, George Ronkin, and Edward P.

Schan. UNIX operating system porting experiences. AT&T Bell Laboratories Technical Journal, 63(8 part 2):1769–1790, October 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Basinger:1985:SSS

[BHK^{+}85] R. G. Basinger, J. A. Herndon, B. Kaskey, J. A. Lindner, and J. M. Milner. The 5ESS switching system: system development environment. AT&T Technical Journal, 64(6 part 1):1485– 1502, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Bergland:1982:VSS

[BHS82] G. D. Bergland, W. T. Hartwell, and G. W. Smith, Jr. 1A voice storage system: Prologue. The Bell System Technical Journal, 61(5):815–819, May/ June 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-5-815.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-5-815.pdf.

Boddie:1981:DSPc

[BJM^{+}81] J. R. Boddie, J. D. Johnston, C. A. McGonegal, J. W. Upton, D. A. Berkley, R. E. Crochiere, and J. L. Flanagan. Digital signal processor: Adaptive differential pulsecode-modulation coding. The Bell System Technical Journal, 60(7):1547–1561, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-7-1547.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-7-1547.pdf.

Berk:1988:MVA

[BJW88] Donald A. Berk, Frank F. Judd, and Stanley C. Wisniewski II. Machine-vision-assisted workstations for lightwave device manufacture. AT&T Technical Journal, 67(2):35–46, March/April 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Bottos:1983:GSD

[BK83] B. A. Bottos and C. M. R. Kintala. Generation of syntaxdirected editors with textoriented features. The Bell System Technical Journal, 62(10): 3205–3224, December 1983. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol62/bstj62-10-3205.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol62-1983/articles/
                                        bstj62-10-3205.pdf.

Bulcha:1982:FPM

[BKL^{+}82] B. Bulcha, L. E. Kodrich, D. B. Luber, W. J. Mitchell, M. A. Schwartz, and F. N. Woomer. Feeder planning methods for digital loop carrier.

The Bell System Technical Journal, 61(9):2129–2141, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2129.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2129.pdf.

Buric:1981:DSP

[BKO81] M. R. Buric, J. Kohut, and J. P. Olive. Digital signal processor: Speech synthesis. The Bell System Technical Journal, 60(7):1621–1631, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-7-1621.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-7-1621.pdf.

Boakye:1988:ADN

[BKP88] Kwame A. Boakye, James C.

Kaufeld, Jr., and John W. Palmer. AT&T data networking architecture. AT&T Technical Journal, 67(6):23–34, November 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Bowman:1982:VSS

[BKV82] P. W. Bowman, J. C. Kennedy, and G. A. VanDine. 1A voice storage system: Office engineering, maintenance, and reliability. The Bell System Technical Journal, 61(5):885–911, May/ June 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-5-885.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-5-885.pdf.

Basso:1989:OOS

[BLT89] R. J. Basso, J. C. Lund, Jr., and J. H. Tendick. OSPS operator services applications. AT&T Technical Journal, 68(6): 25–37, 1989. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Blue:1980:IEE

[Blu80a] J. L. Blue. Integral equations for electromagnetic scattering by wide scatterers. The Bell System Technical Journal, 59(10): 1893–1908, December 1980. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol59/bstj59-10-1893.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol59-1980/articles/
                                        bstj59-10-1893.pdf.

Blue:1980:ECC

[Blu80b]  James L. Blue.  Efficient

calculation of current flow in electroplating cells. The Bell System Technical Journal, 59(2):169–182, February 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol59/bstj59-2-169.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-2-169.pdf.

Benjamin:1982:DDC

[BM82a] R. E. Benjamin and H. H. Mahn.

D4 digital channel bank family: The maintenance bank. The Bell System Technical Journal, 61(9):2665–2676, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2665.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2665.pdf.

Boggs:1982:ARSb

[BM82b] P. S. Boggs and J. R. Mashey.

Automated repair service bureau: Cable repair administrative system. The Bell System Technical Journal, 61(6): 1275–1291, July/August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-6-1275.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-6-1275.pdf.

Bernet:1987:MGT

[BM87] Yoram Bernet and James M.

McCarthy. The 630 Multitasking Graphics Terminal. AT&T Technical Journal, 66(6): 3–14, November 1987. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Bleisch:1984:LCF

[BMD84] George W. Bleisch, William J. Mitchell, and Stanley Dodds. The LT-1 connector family of transmultiplexers. AT&T Bell Laboratories Technical Journal, 63(6 part 1):879–900, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Burgess:1983:ASS

[BMM^{+}83] N. O. Burgess, R. C. MacLean, G. J. Mandeville, D. I. McLean, M. E. Sands, and R. P. Snicer. The AR6A single-sideband microwave radio system: Equalization for multipath fading. The Bell System Technical Journal, 62(10):3409–3428, December 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol62/bstj62-10-3409.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol62-1983/articles/
                                        bstj62-10-3409.pdf.

Budrikis:1984:PCS

[BN84] Zigmantas L. Budrikis and Arun N. Netravali. A packet/ circuit switch. AT&T Bell Laboratories Technical Journal, 63 (8 part 1):1499–1520, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Barker:1988:SSG

[BN88] L. Kirk Barker and Larry D. Nelson. Security standards: government and commercial. AT&T Technical Journal, 67(3):9–18, May 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Burgess:1982:NRS

[BNS82] P. N. Burgess, J. L. Neigh, and R. G. Sparber. No. 10A remote switching system: Transmission plan. The Bell System Technical Journal, 61(4):627–642, April 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-4-627.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-4-627.pdf.

Barnes:1983:TND

[BO83] D. H. Barnes and J. J. O’Connor.

Total network data system: Small office network data system. The Bell System Technical Journal, 62(7):2397–2431, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-7-2397.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-7-2397.pdf.

Byrne:1985:AP

[BO85] William R. Byrne and Gerard P.

O’Reilly. Applications planning. AT&T Technical Journal, 64(6 part 1):1315–1337, 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Boddie:1981:DSPa

[Bod81]  J. R. Boddie.  Digital sig-

nal processor: Overview: The device, support facilities, and applications. The Bell System Technical Journal, 60(7): 1431–1439, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-7-1431.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-7-1431.pdf.

Bartz:1983:TND

[BP83] W. S. Bartz and R. W. Patterson. Total network data system: National network management. The Bell System Technical Journal, 62(7):2261–2280, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-7-2261.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-7-2261.pdf.

Barrese:1983:TND

[BPRS83] A. L. Barrese, D. E. Parker, T. E. Robbins, and L. M. Steele. Total network data system: Environment and objectives. The Bell System Technical Journal, 62(7):2127–2146, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-7-2127.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-7-2127.pdf.

Bergeron:1982:ARS

[BR82a] R. F. Bergeron and M. J.

Rochkind. Automated repair service bureau: Software tools and components. The Bell System Technical Journal, 61(6):1177–1195, July/ August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-6-1177.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-6-1177.pdf.

Brown:1982:UEL

[BR82b] M. K. Brown and L. R. Rabiner. On the use of energy in LPC-based recognition of isolated words. The Bell System Technical Journal, 61(10): 2971–2987, December 1982. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol61/bstj61-10-2971.pdf;
http://www.alcatel-lucent.
com/bstj/vol61-1982/articles/
bstj61-10-2971.pdf.

Brachman:1988:BKR

[Bra88]  Ronald J. Brachman.  Ba-

sics of knowledge representation and reasoning. AT&T Technical Journal, 67(1):7–24, January 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Brecher:1981:MFU

[Bre81]  S. M. Brecher.  Micropro-

cessor firmware update inventory model. The Bell System Technical Journal, 60(10):2293– 2306, December 1981. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol60/bstj60-10-2293.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol60-1981/articles/
                                        bstj60-10-2293.pdf.

Brown:1985:CIR

[Bro85] M. K. Brown. A controlled impedance robot gripper. AT&T Technical Journal, 64(4):937– 969, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Barton:1983:PAD

[BS83a] M. E. Barton and D. A. Schmitt. The 3B20D processor & DMERT operating system: 3B20D craft interface. The Bell System Technical Journal, 62(1):383–397, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-1-383.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-1-383.pdf.

Brockway:1983:ATI

[BS83b] G. S. Brockway and M. R. Santana. Analysis of thermally induced loss in fiber-optic ribbons. The Bell System Technical Journal, 62(4):993–1018, April 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-4-993.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-4-993.pdf.

Burman:1983:AAQ

[BS83c] D. Y. Burman and D. R. Smith.

Asymptotic analysis of a queueing model with bursty traffic. The Bell System Technical Journal, 62(6):1433–1453, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1433.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1433.pdf.

Barkley:1988:PSU

[BS88] Ronald E. Barkley and Curt F.

Schimmel. A performance study of the Unix System V fork system call using Casper. AT&T Technical Journal, 67(5):100– 109, September/October 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Boddie:1981:DSPd

[BST81] J. R. Boddie, N. Sachs, and J. Tow. Digital signal processor: Receiver for touch-tone service. The Bell System Technical Journal, 60(7):1573–1583, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-7-1573.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-7-1573.pdf.

Bayer:1983:ETS

[BT83] D. L. Bayer and R. A. Thompson. An experimental teleterminal — the software strategy. The Bell System Technical Journal, 62(1):121–144, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-1-121.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-1-121.pdf.

Brown:1984:DLB

[BTOA84] M. K. Brown, R. Thorkildsen, Y. H. Oh, and S. S. Ali. The DTWP: an LPC-based dynamic time-warping processor for isolated word recognition. AT&T Bell Laboratories Technical Journal, 63(3):441–457, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Bulfer:1987:APP

[Bul87] A. F. Bulfer. AT&Tt’s payper-view television trial. AT&T Technical Journal, 66(3):54– 63, May/June 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Burke:1982:LBO

[Bur82] P. J. Burke. The limit of the blocking as offered load decreases with fixed peakedness. The Bell System Technical Journal, 61 (10):2911–2916, December 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol61/bstj61-10-2911.pdf;
http://www.alcatel-lucent.
com/bstj/vol61-1982/articles/
bstj61-10-2911.pdf.

Butzien:1981:TDR

[But81] P. E. Butzien. Three-dimensional radiation characteristics of a pyramidal horn-reflector antenna. The Bell System Technical Journal, 60(6):913–921, July/ August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-6-913.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-6-913.pdf.

Butler:1984:CRT

[But84]  Thomas W. Butler.  Com-

puter response time and user performance during data entry.

AT&T Bell Laboratories Technical Journal, 63(6 part 2):1007– 1018, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Bastien:1983:TSP

[BW83a] P. L. Bastien and B. R. Wycherley. Traffic service position system no. 1B: Long-range planning tools. The Bell System Technical Journal, 62(3):959–978, March 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-3-959.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-3-959.pdf.

Budlong:1983:PAD

[BW83b] A. H. Budlong and F. W. Wendland. The 3B20D processor & DMERT operating system: 3B20D input/output system. The Bell System Technical Journal, 62(1):255–273, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-1-255.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-1-255.pdf.

Callahan:1988:ESA

[Cal88] Paul H. Callahan. Expert systems for AT&T switched network maintenance. AT&T Technical Journal, 67(1):93–103, January 1988. CODEN ATJOEM.

ISSN 2376-676X (print), 8756-2324 (electronic).

Campbell:1987:AFC

[Cam87] R. L. Campbell. An architecture for factory control automation. AT&T Technical Journal, 66(5): 77–85, September/October 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Canniff:1981:DCS

[Can81] R. J. Canniff. A digital concentrator for the SLC-96 system. The Bell System Technical Journal, 60(2):121–158, February 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-2-121.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-2-121.pdf.

Carroll:1980:LTM

[Car80]  R. L. Carroll.  Loop tran-

sient measurements in cleveland, south carolina. The Bell System Technical Journal, 59(9):1645–1680, November 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-9-1645.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-9-1645.pdf.

Cargill:1984:DCP

[Car84] Thomas A. Cargill. Debugging C programs with the Blit. AT&T Bell Laboratories Technical Journal, 63(8 part 2):1633–1647, October 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Chang:1981:FCA

[CB81] S. J. Chang and M. A. Breuer. A fault-collapsing analysis in sequential logic networks. The Bell System Technical Journal, 60(9):2259–2271, November 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-9-2259.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-9-2259.pdf.

Candy:1983:CCW

[CB83] J. C. Candy and O. J. Benjamin. A circuit that changes the word rate of pulse code modulated signals. The Bell System Technical Journal, 62(4):1161–1168, April 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-4-1161.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-4-1161.pdf.

Cox:1987:AVP

[CBB^{+}87] Richard V. Cox, Donald E. Bock, Keith B. Bauer, James D. Johnston, and James H. Snyder.

The analog voice privacy system. AT&T Technical Journal, 66(1): 119–131, January 1987. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Chen:1980:SLD

[CC80] F. S. Chen and Y. S. Chen.

Sensitivity loss of digital optical receivers caused by intersymbol interference. The Bell System Technical Journal, 59(10): 1877–1891, December 1980. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol59/bstj59-10-1877.pdf;
http://www.alcatel-lucent.
com/bstj/vol59-1980/articles/
bstj59-10-1877.pdf.

Chu:1985:DFP

[CC85] T. C. Chu and H. C. Chandan.

Determination of fiber proof-test stress for undersea lightguide cable. AT&T Technical Journal, 64(4):971–982, 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Cohen:1988:RTK

[CC88] M. I. Cohen and R. L. Campbell.

Robotics technologies — a key element in achieving manufacturing excellence. AT&T Technical Journal, 67(2):2–4, March/ April 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Carey:1984:EOC

[CCD^{+}84] Michele B. Carey, Han-Tee Chen, Alfred Descloux, James F. Ingle, and Kun I. Park. 1982/83 End Office Connection Study: analog voice and voiceband data transmission performance characterization of the public switched network. AT&T Bell Laboratories Technical Journal, 63(9):2059– 2119, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Carney:1985:SSA

[CCG^{+}85] D. L. Carney, J. I. Cochrane, L. J. Gitten, E. M. Prell, and R. Staehler. The 5ESS switching system: architectural overview. AT&T Technical Journal, 64(6 part 1):1339–1356, 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Crochiere:1982:KDS

[CCJS82] R. E. Crochiere, R. V. Cox, J. D. Johnston, and L. A. Seltzer. A 9.6-kb/s DSP speech coder. The Bell System Technical Journal, 61(9):2263–2288, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-9-2263.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-9-2263.pdf.

Ciesielka:1980:ESL

[CD80] A. J. Ciesielka and D. C. Douglas. Electronics in the suburban and light urban loop networks. The Bell System Technical Journal, 59(3):417–439, March 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-3-417.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-3-417.pdf.

Crane:1989:IOD

[CDG^{+}89] Bently A. Crane, John C. Dalby, Kamla Garg, Michael H. Poepping, and Charles E. W. Ward. International OSPS database services. AT&T Technical Journal, 68(6):38–50, 1989. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Chang:1981:CSP

[CDH81] X. M. Chang, D. Z. Du, and F. K. Hwang. Characterization for series-parallel channel graphs. The Bell System Technical Journal, 60(6):887–891, July/ August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-6-887.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-6-887.pdf.

Coplien:1988:CET

[CDK88] James O. Coplien, Stephen C.

Dewhurst, and Andrew R. Koenig. C++: Evolving toward a more powerful language.

AT&T Technical Journal, 67 (4):19–32, July 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Chu:1983:EBI

[CE83] T. S. Chu and R. W. England. An experimental broadband imaging feed. The Bell System Technical Journal, 62(5):1233–1250, May/ June 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-5-1233.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-5-1233.pdf.

Clemans:1989:BIV

[CEGM89] Jim E. Clemans, Theophilus I. Ejim, William A. Gault, and Eric M. Monberg. Bulk III–V compound semiconductor crystal growth. AT&T Technical Journal, 68(1):29–42, January/February 1989. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (elec-

tronic).  URL http://www.
osti.gov/energycitations/
product.biblio.jsp?osti_id=
7152214.

Crochiere:1986:SPE

[CF86] Ronald E. Crochiere and James L. Flanagan. Speech processing: an evolving technology. AT&T Technical Journal, 65(5):2–11, September/October 1986. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Cichinski:1988:PAT

[CF88] Steve Cichinski and Glenn S.

Fowler. Product administration through Sable and Nmake. AT&T Technical Journal, 67(4): 59–70, July 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Crue:1982:DDC

[CGG^{+}82] C. R. Crue, W. B. Gaunt, Jr., J. H. Green, J. E. Landry, and D. A. Spires. D4 digital channel bank family: The channel bank. The Bell System Technical Journal, 61(9):2611–2664, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2611.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2611.pdf.

Chung:1981:ERT

[CGH81] F. R. K. Chung, R. L. Graham, and F. K. Hwang. Efficient realization techniques for network flow patterns. The Bell System Technical Journal, 60(8):1771–1786, October 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-8-1771.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-8-1771.pdf.

Carestia:1981:NEE

[CH81] P. D. Carestia and F. S. Hudson. No. 4 ESS: Evolution of software structure. The Bell System Technical Journal, 60(6):1167–1201, July/ August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-6-1167.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-6-1167.pdf.

Cheng:1989:AKS

[CHL^{+}89] Yun-Chian Cheng, David J. Houck, Jr., Jun-Min Liu, Marc S. Meketon, Lev Slutsman, Robert J. Vanderbei, and Pyng Wang. The AT&T KORBX system. AT&T Technical Journal, 68(3):7–19, 1989. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Cleveland:1983:HFB

[CHM83] W. S. Cleveland, C. S. Harris, and R. McGill. Human factors and behavioral science: Experiments on quantitative judgements of graphs and maps. The Bell System Technical Journal, 62(6):1659–1674, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1659.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1659.pdf.

Cavanaugh:1980:MSE

[CHN80] J. R. Cavanaugh, R. W. Hatch, and J. L. Neigh. A model for the subjective effects of listener echo on telephone connections. The Bell System Technical Journal, 59(6):1009–1060, July/ August 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-6-1009.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-6-1009.pdf.

Choi:1986:PST

[Cho86] DooWhan Choi. Parallel scrambling techniques for digital multiplexers. AT&T Technical Journal, 65(5):123–136, September/October 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Clark:1983:TSP

[CHTW83] G. T. Clark, H. A. Hilsinger, J. H. Tendick, and R. A. Weber. Traffic service position system no. 1B: Hardware configuration. The Bell System Technical Journal, 62(3):827–858, March 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-3-827.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-3-827.pdf.

Chu:1980:MDE

[Chu80a] T. S. Chu. Microwave depolarization of an Earth–Space path. The Bell System Technical Journal, 59(6):987–1007, July/ August 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-6-987.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-6-987.pdf.

Chung:1980:SNB

[Chu80b] F. R. K. Chung. On switching networks and block designs, II. The Bell System Technical Journal, 59(7):1165–1173, September 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-7-1165.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-7-1165.pdf.

Chu:1983:MCM

[Chu83]  T. C. Chu.  A method

to characterize the mechanical properties of undersea cables. The Bell System Technical Journal, 62(3):703–715, March 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-3-703.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-3-703.pdf.

Chappell:1982:ARS

[CHW82] S. G. Chappell, F. H. Henig, and D. S. Watson. Automated repair service bureau: The front-end system. The Bell System Technical Journal, 61(6):1165–1176, July/ August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-6-1165.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-6-1165.pdf.

Coke:1983:HFB

[CK83] E. U. Coke and M. E. Koether.

Human factors and behavioral science: A study of the match between the stylistic difficulty of technical documents and the reading skills of technical personnel. The Bell System Technical Journal, 62(6):1849–1864, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1849.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1849.pdf.

Cleaveland:1988:TBA

[CK88] J. C. Cleaveland and C. M. R.

Kintala. Tools for building application generators. AT&T Technical Journal, 67(4):46–58, July 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Cox:1988:RTS

[CKKS88] Ingemar J. Cox, David A. Kapilow, Walt J. Kropfl, and Jonathan E. Shopiro. Real-time software for robotics. AT&T Technical Journal, 67(2):61–72, March/April 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Cohen:1988:UNM

[CKP88] Roberta S. Cohen, Hsin K. Kan, and Raymond J. Pennotti. Unified network management from AT&T. AT&T Technical Journal, 67(6):121– 136, November 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Chernak:1982:DDC

[CL82] J. Chernak and J. J. Lang. D4 digital channel bank family: Overview. The Bell System Technical Journal, 61 (9):2607–2610, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-9-2607.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-9-2607.pdf.

Clarke:1983:TRA

[Cla83] R. H. Clarke. Theory of reflection from antireflection coatings. The Bell System Technical Journal, 62(10):2885–2891, December 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol62/bstj62-10-2885.pdf;
http://www.alcatel-lucent.
com/bstj/vol62-1983/articles/
bstj62-10-2885.pdf.

Cohen:1987:REI

[CLMS87] Howard M. Cohen, Michael J.

LuValle, J. Peter Mitchell, and Edward S. Sproles, Jr. Reliability evaluation of interconnection products. AT&T Technical Journal, 66(4):70–80, 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Carroll:1980:LTC

[CM80] R. L. Carroll and P. S. Miller.

Loop transients at the customer site. The Bell System Technical Journal, 59(9): 1609–1643, November 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-9-1609.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-9-1609.pdf.

Cohen:1980:PCD

[CMLF80] L. G. Cohen, W. L. Mammel, C. Lin, and W. G. French. Propagation characteristics of double-mode fibers. The Bell System Technical Journal, 59(6):1061–1072, July/ August 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-6-1061.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-6-1061.pdf.

Cox:1983:MMR

[CMN83] D. C. Cox, R. R. Murray, and A. W. Norris. Measurements of 800-MHz radio transmission into buildings with metallic walls. The Bell System Technical Journal, 62(9):2695–2717, November 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-9-2695.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-9-2695.pdf.

Cox:1984:MAM

[CMN84] Donald C. Cox, Roy R. Murray, and A. W. Norris. 800-MHz attenuation measured in and around suburban houses. AT&T Bell Laboratories Technical Journal, 63(6 part 1):921– 954, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Calderbank:1983:UBM

[CMS83] A. E. Calderbank, J. E. Mazo, and H. M. Shapiro. Upper bounds on the minimum distance of trellis codes. The Bell System Technical Journal, 62(8):2617–2646, October 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-8-2617.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-8-2617.pdf.

Cohen:1981:TSL

[CMSP81] L. G. Cohen, W. L. Mammel, J. Stone, and A. D. Pearson. Transmission studies of a long single-mode fiber — measurements and consideration for bandwidth optimization. The Bell System Technical Journal, 60(8):1713–1725, October 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-8-1713.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-8-1713.pdf.

Cohen:1982:DSI

[Coh82] D. Cohen. Database systems:

Implementation of a distributed database management system to support logical subnetworks. The Bell System Technical Journal, 61(9):2459–2474, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2459.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2459.pdf.

Cole:1988:PMB

[Col88] Robert G. Cole. Performance models of binary synchronous communications multipoint circuits over virtual private-line frame relay networks. AT&T Technical Journal, 67(5):41–56, September/October 1988. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Conte:1980:ESI

[Con80] R. Conte. Effects of scanning interval on peak load measurements. The Bell System Technical Journal, 59 (8):1513–1524, October 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-8-1513.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-8-1513.pdf.

Cho:1982:DDC

[COW82] Y. S. Cho, J. W. Olson, and D. H. Williamson. D4 digital channel bank family: The SLC-96 system. The Bell System Technical Journal, 61(9):2677–2702, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-9-2677.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-9-2677.pdf.

Cox:1981:CTS

[Cox81] R. V. Cox. A comparison of three speech coders to be implemented on the digital signal processor. The Bell System Technical Journal, 60(7):1411–1421, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-7-1411.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-7-1411.pdf.

Coyne:1982:ATM

[Coy82] J. C. Coyne. An approximate thermal model for outdoor electronics cabinets. The Bell System Technical Journal, 61(2):227–246, February 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-2-227.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-2-227.pdf.

Carnevale:1983:EEP

[CP83a] A. Carnevale and U. C. Paek.

Empirical evaluation of profile variations in an MCVD optical waveguide fiber using modal structure analysis. The Bell System Technical Journal, 62(7):1937–1954, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-7-1937.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-7-1937.pdf.

Carnevale:1983:MSM

[CP83b] A. Carnevale and U. C. Paek. Modal structure of an MCVD optical waveguide fiber. The Bell System Technical Journal, 62(6):1415–1431, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-6-1415.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-6-1415.pdf.

Carney:1986:PIS

[CP86] David L. Carney and Edward M. Prell. Planning for ISDN in the 5ESS switch. AT&T Technical Journal, 65(1):35–43, January/February 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Chiang:1982:DSD

[CR82] T. C. Chiang and G. R. Rose. Database systems: Design and implementation of a production database management system (DBM-2). The Bell System Technical Journal, 61(9):2511–2528, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-9-2511.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-9-2511.pdf.

Crochiere:1981:DSP

[Cro81] R. E. Crochiere. Digital signal processor: Sub-band coding. The Bell System Technical Journal, 60(7):1633–1653, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-7-1633.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-7-1633.pdf.

Confalone:1982:SPC

[CRT82] D. E. Confalone, B. W. Rogers, and R. J. Thornberry, Jr. Stored program controlled network: Calling card service — TSPS hardware, software, and signaling implementation. The Bell System Technical Journal, 61(7):1675–1714, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-7-1675.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-7-1675.pdf.

Cornell:1982:VSS

[CS82] R. G. Cornell and J. V. Smith.

1A voice storage system: Architecture and physical design. The Bell System Technical Journal, 61(5):841–861, May/ June 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-5-841.

                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-5-841.pdf.

Crane:1983:TSP

[CS83] B. A. Crane and D. S. Suk. Traffic service position system no. 1B: Capacity and reliability evaluation. The Bell System Technical Journal, 62(3):919–939, March 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-3-919.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-3-919.pdf.

Calderbank:1985:FDM

[CS85] A. R. Calderbank and N. J. A. Sloane. Four-dimensional modulation with an eight-state trellis code. AT&T Technical Journal, 64(5):1005–1018, 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Cox:1983:AVP

[CT83] R. V. Cox and J. M. Tribolet. Analog voice privacy systems using TFSP scrambling: Full duplex and half duplex. The Bell System Technical Journal, 62(1):47–61, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-1-47.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol62-1983/
articles/bstj62-1-47.pdf.

Chin:1981:TC

[CTM81] A. K. Chin, H. Temkin, and S. Mahajan. Transmission cathodoluminescence. The Bell System Technical Journal, 60(9):2187–2226, November 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-9-2187.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-9-2187.pdf.

Courcoubetis:1985:ORA

[CV85] Costas Courcoubetis and P. Varaiya.

Optimal resource allocation for two processes. AT&T Technical Journal, 64(1 part 1):1–14, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Cioffi:1985:EBD

[CW85] John M. Cioffi and Jean-Jacques Werner. Effects of biases on digitally implemented data-driven echo cancelers. AT&T Technical Journal, 64(1 part 1):115– 138, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Dattatreya:1980:MCF

[Dat80]  E. S. Dattatreya.  Multiple

commitment of feeder capacity in the loop plant. The Bell System Technical Journal, 59(1):67–79, January 1980.

CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-1-67.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-1-67.pdf.

Duttweiler:1980:SCV

[DC80] D. L. Duttweiler and Y. S. Chen. A single-chip VLSI echo canceler. The Bell System Technical Journal, 59(2):149–160, February 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-2-149.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-2-149.pdf.

Dixon:1987:LDT

[DD87] Richard W. Dixon and Niloy K. Dutta. Lightwave device technology. AT&T Technical Journal, 66(1):73–83, January 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Decelle:1988:DRW

[Dec88] L. S. Decelle. Design of a robotic workstation for component insertion. AT&T Technical Journal, 67(2):15–22, March/April 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

DeTreville:1984:SBC

[DeT84] J. D. DeTreville. A simulationbased comparison of voice transmission on CSMA/CD networks and on token buses. AT&T Bell Laboratories Technical Journal, 63(1):33–55, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Davis:1981:MDT

[DFN81] J. R. Davis, L. E. Forman, and L. T. Nguyen. The metropolitan digital truck plant. The Bell System Technical Journal, 60(6):933–964, July/August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-6-933.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-6-933.pdf.

Delatore:1985:SSO

[DFOS85] J. P. Delatore, R. J. Frank, H. Oehring, and L. C. Stecher. The 5ESS switching system: operational software. AT&T Technical Journal, 64(6 part 1):1357– 1384, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Davis:1982:SPC

[DFS82] E. A. Davis, C. J. Funk, and J. M. Sebeson. Stored program controlled network: CCIS and SPC network performance. The Bell System Technical Journal, 61(7):1637–1654, September 1982. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-7-1637.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-7-1637.pdf.

Dragone:1980:SPA

[DG80] C. Dragone and M. J. Gans. Satellite phased arrays: Use of imaging reflectors with spatial filtering in the focal plane to reduce grating lobes. The Bell System Technical Journal, 59(3):449–461, March 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-3-449.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-3-449.pdf.

Davis:1981:NEP

[DG81] E. A. Davis and P. K. Giloth. No. 4 ESS: Performance objectives and service experience. The Bell System Technical Journal, 60(6):1203–1224, July/ August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-6-1203.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-6-1203.pdf.

Dunbar:1982:DDC

[DGH^{+}82] B. J. Dunbar, D. V. Gupta, M. P. Horvath, S. P. Verma, and R. E. Sheehey. D4 digital channel bank family: Dataport — channel units for digital data system 56-kb/s rate. The Bell System Technical Journal, 61(9):2741–2756, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2741.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2741.pdf.

Du:1982:CBP

[DH82] D. Z. Du and F. K. Hwang.

Comparisons on blocking probabilities for regular series parallel channel graphs. The Bell System Technical Journal, 61(8):1965–1973, October 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-8-1965.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-8-1965.pdf.

Dowden:1989:OSF

[DHK89] D. C. Dowden, D. E. Herr, and B. Kaskey. Operator services feature of the 5ESS switch. AT&T Technical Journal, 68(6): 3–8, 1989. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Dunietz:1986:MMP

[DHM^{+}86] I. S. Dunietz, J. L. C. Hsu, M. T. McEachern, J. H. Stocking, M. A. Swartz, and R. M. Trombly. MPCS: The Manufacturing Process Control System. AT&T Technical Journal, 65(4):35–45, July/August 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Dixon:1980:CDG

[Dix80] R. W. Dixon. Current directions in GaAs laser device development. The Bell System Technical Journal, 59(5):669–722, May/ June 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-5-669.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-5-669.pdf.

Dixon:1980:AGD

[DJ80] R. W. Dixon and W. B. Joyce. (Al,Ga)As doubleheterostructure lasers: Comparison of devices fabricated with deep and shallow proton bombardment. The Bell System Technical Journal, 59 (6):975–985, July/August 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-6-975.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-6-975.pdf.

Donegan:1983:HFB

[DK83] J. Donegan and D. N. Koppes.

Human factors and behavioral science: Human factors engineering for the loop plant. The Bell System Technical Journal, 62(6):1701–1703, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1701.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1701.pdf.

Delatore:1988:IDN

[DKW88] John P. Delatore, Patrick Krause, and Randall J. Wilson. ISDN data networking applications in the corporate environment. AT&T Technical Journal, 67(6):107–120, November 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Dawson:1980:RHS

[DKZ80] L. R. Dawson, V. G. Keramidas, and C. L. Zipfel. Reliable, high-speed LEDs for shorthaul optical data links. The Bell System Technical Journal, 59(2):161–168, February 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-2-161.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-2-161.pdf.

Dragone:1984:QWC

[DL84] C. Dragone and W. E. Legg.

Quarter-wave corrugated transformer for broadband matching of a corrugated feed. AT&T Bell Laboratories Technical Journal, 63(2):207–215, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

delaRosa:1986:WIL

[dlRR86] John G. de la Rosa and Deni M. Rose. Wafer inspection with a laser scanning microscope. AT&T Technical Journal, 65(1): 68–77, January/February 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

DeLessio:1983:TSP

[DM83] N. X. DeLessio and N. A. Martellotto. Traffic service position system no. 1B: System description. The Bell System Technical Journal, 62(3):765–774, March 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-3-765.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-3-765.pdf.

Doshi:1988:CCI

[DN88] Bharat T. Doshi and Han Q. Nguyen. Congestion control in ISDN frame-relay networks. AT&T Technical Journal, 67(6): 35–46, November 1988. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Doshi:1983:MGQ

[Dos83]  B. T. Doshi.  An M/G/1

queue with a hybrid discipline. The Bell System Technical Journal, 62(5):1251–1271, May/ June 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-5-1251.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-5-1251.pdf.

Doshi:1985:ACS

[Dos85]  B. T. Doshi.  Analysis of

clocked schedules-high-priority tasks. AT&T Technical Journal, 64(2 part 1):633–660, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Davis:1981:CCP

[DR81] J. R. Davis and A. K. Reilly.

T-carrier characterization program — overview. The Bell System Technical Journal, 60 (6):929–932, July/August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-6-929.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-6-929.pdf.

Dragone:1981:CMC

[Dra81a] C. Dragone. Conformal mapping and complex coordinates on Cassegrainian and Gregorian reflector antennas. The Bell System Technical Journal, 60(10): 2397–2420, December 1981. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol60/bstj60-10-2397.pdf;
http://www.alcatel-lucent.
com/bstj/vol60-1981/articles/
bstj60-10-2397.pdf.

Dragone:1981:HFB

[Dra81b] C. Dragone. High-frequency behavior of waveguides with finite surface impedances. The Bell System Technical Journal, 60(1):89–116, January 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-1-89.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-1-89.pdf.

Dautrich:1983:ESS

[DRM83] B. A. Dautrich, L. R. Rabiner, and T. B. Martin. The effects of selected signal processing techniques on the performance of a filter-bank-based isolated word recognizer. The Bell System Technical Journal, 62(5):1311–1336, May/ June 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-5-1311.
pdf;   http://www.alcatel-

lucent.com/bstj/vol62-1983/
articles/bstj62-5-1311.pdf.

Dubois:1983:ASS

[DRS83] A. Dubois, D. N. Ritchie, and F. M. Smith. The AR6A single-sideband microwave radio system: Terminal multiplex equipment. The Bell System Technical Journal, 62(10): 3357–3375, December 1983. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol62/bstj62-10-3357.pdf;
http://www.alcatel-lucent.
com/bstj/vol62-1983/articles/
bstj62-10-3357.pdf.

Dale:1982:ARS

[DRT82] O. B. Dale, T. W. Robinson, and E. J. Theriot. Automated repair service bureau: Mechanized loop testing design. The Bell System Technical Journal, 61(6):1235–1256, July/ August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-6-1235.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-6-1235.pdf.

DeTreville:1984:PTD

[DS84] John D. DeTreville and W. David Sincoskie. Program transformations for data access in a local distributed environment. AT&T Bell Laboratories Technical Journal, 63(6 part 2):1019– 1028, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Derman:1985:HHE

[DS85] E. Derman and E. G. Sheppard. HEQS — a hierarchical equation solver. AT&T Technical Journal, 64(9):2061–2096, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Dautremont-Smith:1989:FTI

[DSMBB89] William C. Dautremont-Smith, R. J. McCoy, Randolph H. Burton, and Albert G. Baca. Fabrication technologies for III–V compound semiconductor photonic and electronic devices. AT&T Technical Journal, 68(1): 64–82, January/February 1989. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Delatore:1985:FST

[DTV85] John P. Delatore, Monte P. Tull, and D. Van Haften. Factory system testing. AT&T Technical Journal, 64(6 part 2):1537– 1558, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Dutta:1989:IVD

[Dut89] Niloy K. Dutta. III–V device technologies for lightwave applications. AT&T Technical Journal, 68(1):5–18, January/February 1989. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Dalby:1983:TSP

[DVW83] J. C. Dalby, Jr., D. Van Haften, and L. A. Weber. Traffic service position system no. 1B: Retrofitting the processor. The Bell System Technical Journal, 62(3):907–918, March 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-3-907.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-3-907.pdf.

Eldredge:1983:PAD

[EC83] G. P. Eldredge and J. C.

Chevalier. The 3B20D processor & DMERT operating system: 3B20 field utilities. The Bell System Technical Journal, 62(1):341–348, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-1-341.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-1-341.pdf.

Eckler:1985:SAL

[Eck85]  A. R. Eckler.  A statisti-

cal approach to laser certification. AT&T Technical Journal, 64(3):765–770, 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Eckel:1986:QAN

[Eck86] Eugene J. Eckel. Quality in AT&T network systems. AT&T Technical Journal, 65(2):30–38, March/April 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Elliot:1986:AST

[EEF^{+}86] Ray S. Elliot, David N. Ernst, William M. Flegal, Eugene A. Haney, and John T. Kamino. The AT&T soft touch-sensitive screen. AT&T Technical Journal, 65(1):56–61, January/ February 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Eldumiati:1981:DSP

[EG81] I. I. Eldumiati and R. N. Gadenz. Digital signal processor: Logic and fault simulations. The Bell System Technical Journal, 60(7):1463–1473, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-7-1463.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-7-1463.pdf.

Egan:1983:HFB

[Ega83] D. E. Egan. Human factors and behavioral science: Retrospective reports reveal differences in people’s reasoning. The Bell System Technical Journal, 62(6):1675–1697, July/ August 1983. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1675.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1675.pdf.

Erbrecht:1981:AMS

[EGKS81] G. F. Erbrecht, J. H. Gentry, Jr., T. C. Kaup, and R. J. Schweizer. Automated measurement system for T1 characterization. The Bell System Technical Journal, 60(6):1005–1026, July/ August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-6-1005.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-6-1005.pdf.

Eng:1981:STC

[EH81] K. Y. Eng and B. G. Haskell.

Study of a time-compression technique for TV transmission using a chirp filter and envelope detection. The Bell System Technical Journal, 60(10): 2373–2395, December 1981. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol60/bstj60-10-2373.pdf;
http://www.alcatel-lucent.
com/bstj/vol60-1981/articles/
bstj60-10-2373.pdf.

Eng:1982:TBC

[EH82] K. Y. Eng and B. G. Haskell.

TV bandwidth compression techniques using time-companded differentials and their applications to satellite transmissions. The Bell System Technical Journal, 61(10):2917–2927, December 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol61/bstj61-10-2917.pdf;
http://www.alcatel-lucent.
com/bstj/vol61-1982/articles/
bstj61-10-2917.pdf.

Ebner:1983:TND

[EH83a] G. C. Ebner and D. G. Haenschke. Total network data system: Network management. The Bell System Technical Journal, 62(7):2239–2260, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-7-2239.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-7-2239.pdf.

Eng:1983:SNT

[EH83b] K. Y. Eng and B. G. Haskell. Synchronization of noncolocated TV signals in a satellite time-compression multiplexing system. The Bell System Technical Journal, 62(10):2867– 2883, December 1983. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol62/bstj62-10-2867.pdf;

http://www.alcatel-lucent.
com/bstj/vol62-1983/articles/
bstj62-10-2867.pdf.

Eng:1983:TCM

[EHS83] K. Y. Eng, B. G. Haskell, and R. L. Schmidt. Timecompression multiplexing (TCM) of three broadcast-quality TV signals in a satellite transponder. The Bell System Technical Journal, 62(10):2853–2865, December 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol62/bstj62-10-2853.pdf;
http://www.alcatel-lucent.
com/bstj/vol62-1983/articles/
bstj62-10-2853.pdf.

Eichenbaum:1980:COF

[Eic80] B. R. Eichenbaum. The centering of optical fiber coatings by monitoring forward scattering patterns — theory and practice. The Bell System Technical Journal, 59(3):313–332, March 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-3-313.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-3-313.pdf.

Eigen:1983:HFB

[Eig83] D. J. Eigen. Human factors and behavioral science: Methods for field testing new telephone services. The Bell System Technical Journal, 62(6):1591–1616, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-6-1591.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-6-1591.pdf.

Einarsson:1980:AAT

[Ein80]  G. Einarsson.  Address as-

signment for a time-frequency-coded, spread-spectrum system. The Bell System Technical Journal, 59(7):1241–1255, September 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-7-1241.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-7-1241.pdf.

Eisenstein:1984:TDS

[Eis84] G. Eisenstein. Theoretical design of single-layer antireflection coatings on laser facets. AT&T Bell Laboratories Technical Journal, 63(2):357–364, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Elken:1980:AMP

[Elk80] T. R. Elken. The application of mathematical programming to loop feeder allocation. The Bell System Technical Journal, 59(4):479–500, April 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol59/bstj59-4-479.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-4-479.pdf.

Elsner:1980:DTG

[Els80]  W. B. Elsner.  Dimension-

ing trunk groups for digital networks. The Bell System Technical Journal, 59(7): 1123–1138, September 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-7-1123.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-7-1123.pdf.

Erickson:1984:BSC

[EP84] Verlyn B. Erickson and John F.

Pellegrin. Build — a software construction tool. AT&T Bell Laboratories Technical Journal, 63(6 part 2):1049–1059, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Eng:1981:SPB

[EY81] K. Y. Eng and O. C. Yue. Spectral properties and band-limiting effects of time-compressed TV signals in a time-compression multiplexing system. The Bell System Technical Journal, 60(9):2167–2185, November 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-9-2167.

                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-9-2167.pdf.

Eigen:1982:SPC

[EY82] D. J. Eigen and E. A. Youngs. Stored program controlled network: Calling card service — human factors studies. The Bell System Technical Journal, 61(7):1715–1735, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-7-1715.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-7-1715.pdf.

Fang:1981:EEC

[Fan81] G. S. Fang. Effect of echo canceler on common-channel interoffice signaling continuity check. The Bell System Technical Journal, 60(7):1313–1334, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-7-1313.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-7-1313.pdf.

Farber:1989:AUI

[Far89] J. M. Farber. The AT&T userinterface architecture. AT&T Technical Journal, 68(5):9–16, September/October 1989. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Favin:1981:DSPa

[Fav81] D. L. Favin. Digital signal processor: Tone generation. The Bell System Technical Journal, 60(7):1655–1671, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-7-1655.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-7-1655.pdf.

Fulton:1982:BJB

[FD82] T. A. Fulton and L. N. Dunkleberger. B.S.T.J. briefs: A Josephson parallel multiplier. The Bell System Technical Journal, 61(5):931–933, May/ June 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-5-931.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-5-931.pdf.

Feder:1984:EUS

[Fed84]  Jerome Feder.  Evolution

of UNIX system performance. AT&T Bell Laboratories Technical Journal, 63(8 part 2): 1791–1814, October 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Feuer:1981:SIT

[Feu81] A. Feuer. Special-information-tone frequency detection. The Bell System Technical Journal, 60(7):1289–1312, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-7-1289.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-7-1289.pdf.

Feuer:1983:FAG

[Feu83] A. Feuer. Forecasting with adaptive gradient exponential smoothing. The Bell System Technical Journal, 62 (8):2561–2580, October 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-8-2561.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-8-2561.pdf.

Frakes:1988:CES

[FF88] William B. Frakes and Christopher J. Fox. CEST: an expert system function library and workbench for UNIX system/ C language. AT&T Technical Journal, 67(2):95–106, March/ April 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Fuhrer:1985:SSO

[FGNS85] P. T. Fuhrer, L. J. Gitten, B. A. Newman, and B. E. Snyder. The 5ESS switching system: operations, administration, and maintenance capabilities. AT&T Technical Journal, 64(6 part 1):1523–1536, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Fulton:1983:TND

[FGP^{+}83] N. D. Fulton, J. J. Galiardi, E. J.

Pasternak, S. A. Schulman, and H. E. Voigt. Total network data system: Equipment systems. The Bell System Technical Journal, 62(7):2281–2343, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-7-2281.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-7-2281.pdf.

Freedman:1980:DES

[FH80] H. T. Freedman and T. R.

Harms. A discrete-event simulation analysis of loop network assignment operations. The Bell System Technical Journal, 59(1):81–98, January 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-1-81.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-1-81.pdf.

Fredericks:1988:IPM

[FH88] Albert A. Fredericks and Jack M.

Holtzman. An introduction to performance modeling and analysis. AT&T Technical Journal, 67(5):2–3, September/October 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Franks:1987:PIS

[FHH^{+}87] Richard L. Franks, James P. Holtman, John L. C. Hsu, L. Gary Raymer, and Bernard E. Snyder. Productivity improvement systems for manufacturing. AT&T Technical Journal, 66(5): 61–76, September/October 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Fitton:1984:VPM

[FHK^{+}84] Michael J. Fitton, Carol J. Harkness, Keith A. Kelleman, Paul F. Long, and Carl Mee III. Virtual Protocol Machine. AT&T Bell Laboratories Technical Journal, 63(8 part 2):1859–1876, October 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Fritz:1984:NCR

[FHR84] T. E. Fritz, J. E. Hefner, and T. M. Raleigh. Network of computers running the UNIX system. AT&T Bell Laboratories Technical Journal, 63(8 part 2): 1877–1896, October 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Frank:1981:NEM

[FKSW81] R. J. Frank, R. J. Keevers, F. B. Strebendt, and J. E. Waninski. No. 4 ESS: Mass announcement capability. The Bell System Technical Journal, 60(6):1049–1081, July/ August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-6-1049.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-6-1049.pdf.

Flanagan:1980:DDA

[Fla80] J. L. Flanagan. Direct digitalto-analog conversion of acoustic signals. The Bell System Technical Journal, 59(9): 1693–1719, November 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-9-1693.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-9-1693.pdf.

Flamm:1983:HFB

[Fla83] L. E. Flamm. Human factors and behavioral science: Performance in locating terminals on a high-density connector. The Bell System Technical Journal, 62(6):1723–1731, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1723.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1723.pdf.

Fleming:1984:AAS

[Fle84] Philip J. Fleming. An approximate analysis of sojourn times in the M/G/1 queue with round-robin service discipline. AT&T Bell Laboratories Technical Journal, 63(8 part 1):1521– 1535, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Furnas:1983:HFB

[FLGD83] G. W. Furnas, T. K. Landauer, L. M. Gomez, and S. T. Dumais. Human factors and behavioral science: Statistical semantics: Analysis of the potential performance of key-word information systems. The Bell System Technical Journal, 62(6):1753–1806, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-6-1753.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-6-1753.pdf.

Francis:1981:MLB

[FLRT81] S. H. Francis, H. R. Lunde, G. A. Reinold, and T. E. Talpey. Magnetic localization of buried cable by the SCARAB submersible. The Bell System Technical Journal, 60(4):535–574, April 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-4-535.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol60-1981/
articles/bstj60-4-535.pdf.

Frerking:1982:SPC

[FM82] R. F. Frerking and M. A.

McGrew. Stored program controlled network: Routing of direct-signaling messagesin the CCIS network. The Bell System Technical Journal, 61(7):1599–1609, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-7-1599.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-7-1599.pdf.

Fraser:1984:QFD

[FM84] A. G. Fraser and Samuel P. Morgan. Queueing and framing disciplines for a mixture of data traffic types. AT&T Bell Laboratories Technical Journal, 63(6 part 2):1061–1087, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Felton:1984:USI

[FMM84] William A. Felton, Gerald L.

Miller, and J. Michael Milner. A UNIX system implementation for System/370. AT&T Bell Laboratories Technical Journal, 63(8 part 2):1751–1767, October 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Foschini:1984:CPF

[Fos84] Gerard J. Foschini. Contrasting performance of faster binary signaling with QAM. AT&T Bell Laboratories Technical Journal, 63(8 part 1):1419–1445, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Foschini:1985:EAD

[Fos85] G. J. Foschini. Equalizing without altering or detecting data [digital radio systems]. AT&T Technical Journal, 64(8):1885– 1911, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Fowler:1986:PIW

[Fow86] Priscilla J. Fowler. In-process inspections of workproducts at AT&T. AT&T Technical Journal, 65(2):102–112, March/April 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Fenderson:1984:ATE

[FPQ^{+}84] Gerald L. Fenderson, James W. Parker, Patrick D. Quigley, Scott R. Shepard, and Curtis A. Siller Jr. Adaptive transversal equalizaton of multipath propagation for 16-QAM, 90-Mb/s digital radio. AT&T Bell Laboratories Technical Journal, 63 (8 part 1):1447–1463, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Frase:1983:HFB

[Fra83]  L. T. Frase.  Human fac-

tors and behavioral science: The UNIX Writer’s Workbench software: Philosophy. The Bell System Technical Journal, 62(6):1883–1890, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1883.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1883.pdf.

Fredericks:1980:ACV

[Fre80a]  A. A. Fredericks.  Approx-

imations for customer-viewed delays in multiprogrammed, transaction-oriented computer systems. The Bell System Technical Journal, 59(9): 1559–1574, November 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-9-1559.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-9-1559.pdf.

Fredericks:1980:CBS

[Fre80b]  A. A. Fredericks.  Conges-

tion in blocking systems — A simple approximation technique. The Bell System Technical Journal, 59(6):805–827, July/ August 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol59/bstj59-6-805.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-6-805.pdf.

Fredericks:1982:CAW

[Fre82] A. A. Fredericks. A class of approximations for the waiting time distribution in a GI/G/1 queueing system. The Bell System Technical Journal, 61(3):295–325, March 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-3-295.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-3-295.pdf.

Fredericks:1986:PAM

[Fre86]  Albert A. Fredericks.  Per-

formance analysis modeling for manufacturing systems. AT&T Technical Journal, 65(4):25–34, July/August 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Felsberg:1983:ASS

[FS83a] R. I. Felsberg and M. E. Sands. The AR6A single-sideband microwave radio system: Special test equipment. The Bell System Technical Journal, 62(10): 3477–3489, December 1983. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol62/bstj62-10-3477.pdf;
                                        http://www.alcatel-lucent.

com/bstj/vol62-1983/articles/
bstj62-10-3477.pdf.

Foschini:1983:DCF

[FS83b] G. J. Foschini and J. Salz.

Digital communications over fading radio channels. The Bell System Technical Journal, 62(2):429–456, February 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-2-429.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-2-429.pdf.

Factor:1988:DIS

[FS88] Robert M. Factor and William B.

Smith. A discipline for improving software productivity. AT&T Technical Journal, 67 (4):2–9, July 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Feigenbaum:1983:TNK

[FT83] J. Feigenbaum and R. E. Tarjan. Two new kinds of biased search trees. The Bell System Technical Journal, 62(10): 3139–3158, December 1983. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol62/bstj62-10-3139.pdf;
http://www.alcatel-lucent.
com/bstj/vol62-1983/articles/
bstj62-10-3139.pdf.

Fallah:1987:SDQ

[FT87] Hosein Fallah and George T. Tucker. Software design quality review. AT&T Technical Journal, 66(2):40–49, 1987. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Fuchs:1986:QTP

[Fuc86] E. Fuchs. Quality: theory and practice. AT&T Technical Journal, 65(2):4–8, March/ April 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Franklin:1982:ARS

[FV82] C. M. Franklin and J. F. Vogler. Automated repair service bureau: Data base system. The Bell System Technical Journal, 61(6):1131–1151, July/ August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-6-1131.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-6-1131.pdf.

Flink:1988:SVM

[FW88] Charles W. Flink and Jonathan D. Weiss. System V/MLS labeling and mandatory policy alternatives. AT&T Technical Journal, 67(3):53–64, May 1988. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Favin:1981:DSPb

[FYC81] D. L. Favin, D. P. Yorkgitis, and S. Cordray. Digital signal processor: Power measurements. The Bell System Technical Journal, 60(7):1673–1685, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-7-1673.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-7-1673.pdf.

Gloge:1980:HSD

[GAB^{+}80] D. Gloge, A. Albanese, C. A.

Burrus, E. L. Chinnock, J. A. Copeland, A. G. Dentai, T. P. Lee, Tingye Li, and K. Ogawa. High-speed digital lightwave communication using LEDs and PIN photodiodes at 1.3 μ m. The Bell System Technical Journal, 59(8):1365–1382, October 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-8-1365.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-8-1365.pdf.

Gadenz:1981:DSP

[Gad81] R. N. Gadenz. Digital signal processor: Tone detection for CCITT no. 5 transceiver. The Bell System Technical Journal, 60(7):1687–1698, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-7-1687.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-7-1687.pdf.

Gavish:1985:MCL

[Gav85] Bezalel Gavish. Models for configuring large-scale distributed computing systems. AT&T Technical Journal, 64(2 part 2):491– 532, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Glushko:1982:ARS

[GB82] R. J. Glushko and M. H. Bianchi. Automated repair service bureau: On-line documentation: Mechanizing develoment, delivery, and use. The Bell System Technical Journal, 61(6):1313–1323, July/ August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-6-1313.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-6-1313.pdf.

Greenstein:1980:PMM

[GC80] L. J. Greenstein and B. A. Czekaj. A polynomial model for multipath fading channel responses. The Bell System Technical Journal, 59(7): 1197–1225, September 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol59/bstj59-7-1197.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-7-1197.pdf.

Greenstein:1981:MMF

[GC81] L. J. Greenstein and B. A.

Czekaj. Modeling multipath fading responses using multitone probing signals and polynomial approximation. The Bell System Technical Journal, 60(2):193–214, February 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-2-193.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-2-193.pdf.

Grzelakowski:1983:PAD

[GCD83] M. E. Grzelakowski, J. H.

Campbell, and M. R. Dubman. The 3B20D processor & DMERT operating system: DMERT operating system. The Bell System Technical Journal, 62(1):303–322, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-1-303.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-1-303.pdf.

Gray:1987:BRP

[GE87] J. C. Gray and M. L. Ehlers.

The business resources planning system. AT&T Technical Journal, 66(5):49–60, September/October 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Gehani:1981:PDS

[Geh81]  N. Gehani.  Program de-

velopment by stepwise refinement and related topics. The Bell System Technical Journal, 60(3):347–378, March 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-3-347.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-3-347.pdf.

Grantges:1983:TND

[GFGB83] R. F. Grantges, V. L. Fahrmann, T. A. Gibson, and L. M. Brown. Total network data system: Central office equipment reports for stored program control systems. The Bell System Technical Journal, 62(7):2365–2395, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-7-2365.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-7-2365.pdf.

Gopinath:1984:BPS

[GGV84] B. Gopinath, J.-M. Garcia, and P. Varaiya. Blocking probability in a switching center with arbitrary routing policy. AT&T Bell Laboratories Technical Journal, 63(5):709–720, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Greenside:1981:NIS

[GH81] H. S. Greenside and E. Helfand.

Numerical integration of stochastic differential equations — II. The Bell System Technical Journal, 60(8):1927–1940, October 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-8-1927.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-8-1927.pdf.

Gauthier:1982:ARS

[GH82] R. F. Gauthier and W. A. Harris. Automated repair service bureau: Two examples of human performance analysis and design in planning the ARSB. The Bell System Technical Journal, 61(6):1301–1312, July/ August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-6-1301.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-6-1301.pdf.

Gharavi:1984:CVL

[Gha84]  H. Gharavi.  Conditional

variable-length coding for graylevel pictures. AT&T Bell Laboratories Technical Journal, 63 (2):249–260, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Goodman:1980:FHM

[GHP80] D. J. Goodman, P. S. Henry, and V. K. Prabhu. Frequencyhopped multilevel FSK for mobile radio. The Bell System Technical Journal, 59(7): 1257–1275, September 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-7-1257.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-7-1257.pdf.

Giguere:1980:SOR

[Gig80] W. J. Giguere. SARTS — an overview of remote specialservice testing in the Bell System. The Bell System Technical Journal, 59(4):501–527, April 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-4-501.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-4-501.pdf.

Gingrich:1983:HFB

[Gin83] P. S. Gingrich. Human factors and behavioral science: The UNIX Writer’s Workbench software: Results of a field study. The Bell System Technical Journal, 62(6):1909–1921, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1909.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1909.pdf.

Gonda:1983:ASS

[GK83] J. Gonda and J. M. Kiker, Jr. The AR6A single-sideband microwave radio system: Microwave carrier supply. The Bell System Technical Journal, 62 (10):3391–3408, December 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol62/bstj62-10-3391.pdf;
http://www.alcatel-lucent.
com/bstj/vol62-1983/articles/
bstj62-10-3391.pdf.

Gitlin:1988:DNI

[GK88] Richard D. Gitlin and James C.

Kaufeld, Jr. Data networking: An issue overview. AT&T Technical Journal, 67(6):3–6, November 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Gill:1983:TSP

[GKS83] R. J. Gill, G. J. Kujawinski, and E. H. Stredde. Traffic service position system no. 1B: Real-time architecture utilizing the DMERT operating system. The Bell System Technical Journal, 62(3):775–826, March 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-3-775.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-3-775.pdf.

Gates:1982:VSS

[GKW82] G. W. Gates, R. F. Kranzmann, and L. D. Whitehead. 1A voice storage system: Software. The Bell System Technical Journal, 61(5):863–883, May/ June 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-5-863.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-5-863.pdf.

Gibson:1980:CSM

[GL80] A. E. Gibson and D. B. Luber. Critical section methods for loop plant allocation. The Bell System Technical Journal, 59(1):99–117, January 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-1-99.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-1-99.pdf.

Gersho:1981:ACI

[GL81] A. Gersho and T. L. Lim. Adaptive cancellation of intersymbol interference for data transmission. The Bell System Technical Journal, 60(9): 1997–2021, November 1981.

CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-9-1997.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-9-1997.pdf.

Glance:1984:BDC

[Gla84] B. S. Glance. BER degradations caused by switching in digital mobile radio systems using base station diversity. AT&T Bell Laboratories Technical Journal, 63(4):545–564, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Gawron:1982:SPC

[GLM82] L. J. Gawron, S. J. Lueders, and K. L. Moeller. Stored program controlled network: No. 1/1A ESS — SPC network capabilities and signaling architecture. The Bell System Technical Journal, 61(7):1611–1636, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-7-1611.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-7-1611.pdf.

Goldstein:1982:DDC

[GLMS82] R. M. Goldstein, J. D. Leggett, G. L. Mowery, and K. F. Sodomsky. D4 digital channel bank family: Custom-integrated circuits for digital terminals. The Bell System Technical Journal, 61(9):2791–2813, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-9-2791.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-9-2791.pdf.

Grampp:1984:UOS

[GM84] Frederick T. Grampp and Robert H. Morris. UNIX operating system security. AT&T Bell Laboratories Technical Journal, 63(8 part 2):1649–1672, October 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic). URL https:/

                                        /ieeexplore.ieee.org/stamp/
                                        stamp.jsp?tp=&arnumber=6771912.

Gitlin:1987:CTT

[GM87] Richard D. Gitlin and Howard C. Meadors, Jr. Center-tap tracking algorithms for timing recovery. AT&T Technical Journal, 66 (6):63–78, November 1987. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Gammie:1983:ASS

[GMMR83] J. Gammie, J. P. Moffatt, R. H. Moseley, and W. A. Robinson. The AR6A single-sideband microwave radio system: System design and performance. The Bell System Technical Journal, 62(10):3255–3312, December 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol62/bstj62-10-3255.pdf;
http://www.alcatel-lucent.
com/bstj/vol62-1983/articles/
bstj62-10-3255.pdf.

Gundaker:1987:QTP

[GMT87] Bruce F. Gundaker, David E.

Martinich, and Michael J. Tortorella. Quality technology in product realization systems. AT&T Technical Journal, 66(5): 5–20, September/October 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Gitlin:1982:TLA

[GMW82] R. D. Gitlin, H. C. Meadors, Jr., and S. B. Weinstein. The tap-leakage algorithm: An algorithm for the stable operation of a digitally implemented, fractionally spaced adaptive equalizer. The Bell System Technical Journal, 61(8):1817–1839, October 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-8-1817.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-8-1817.pdf.

Gharavi:1983:CCC

[GN83] H. Gharavi and A. N. Netravali. CCITT compatible coding of multilevel pictures. The Bell System Technical Journal, 62(9):2765–2778, November 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-9-2765.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-9-2765.pdf.

Gardner:1981:EOF

[GNS^{+}81] W. B. Gardner, S. R. Nagel, M. I. Schwartz, F. V. DiMarcello, C. R. Lovelace, D. L. Brownlow, M. R. Santana, and E. V. Sigety. The effect of optical fiber core and cladding diameter on the loss added by packaging and thermal cycling. The Bell System Technical Journal, 60(6):859–864, July/ August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-6-859.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-6-859.pdf.

Godfrey:1986:HEQ

[God86] A. B. Godfrey. The history and evolution of quality in AT&T. AT&T Technical Journal, 65(2): 9–20, March/April 1986. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Goguen:1982:DSD

[Gog82]  N. H. Goguen.  Database

systems: Design of a prototype Videotex system. The Bell System Technical Journal, 61(9):2475–2486, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2475.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2475.pdf.

Goldstein:1982:DSD

[Gol82] A. J. Goldstein. Database systems: A directed hypergraph database: A model for the local loop telephone plant. The Bell System Technical Journal, 61(9):2529–2554, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2529.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2529.pdf.

Govind:1987:NIC

[GOP87] P. K. Govind, Glen E. Offord, and Lisa J. Piper. Networkinterface chip for ISDN terminals. AT&T Technical Journal, 66(2):27–39, 1987. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Garey:1984:OSW

[GP84] Michael R. Garey and Ron Y.

Pinter. Optimum scan-width selection under containment constraints. AT&T Bell Laboratories Technical Journal, 63(6 part 2):1191–1212, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Gartside:1987:SMC

[GPP87] Charles H. Gartside III, Andrew J. Panuska, and Parbhubhai D. Patel. Single-mode cable for long-haul, trunk and loop networks. AT&T Technical Journal, 66(1):84–94, January 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Graham:1983:SPC

[Gra83] N. Y. Graham. Smoothing with periodic cubic splines. The Bell System Technical Journal, 62(1):101–110, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-1-101.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-1-101.pdf.

Greenstein:1985:ASS

[Gre85]  L. J. Greenstein.  Analy-

sis/simulation study of crosspolarization cancellation in dualpolarization digital radio. AT&T Technical Journal, 64(10):2261– 2280, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Grimado:1982:TMT

[Gri82] P. B. Grimado. A theoretical model of transient heat transfer in a firestopped cable bundle. The Bell System Technical Journal, 61(9):2313–2331, November 1982. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2313.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2313.pdf.

Grimado:1983:TCM

[Gri83] P. B. Grimado. TELBECC — A computational method and computer program for analyzing telephone building energy consumption and control. The Bell System Technical Journal, 62(10): 2935–2960, December 1983. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol62/bstj62-10-2935.pdf;
http://www.alcatel-lucent.
com/bstj/vol62-1983/articles/
bstj62-10-2935.pdf.

Goodman:1983:CSC

[GS83a] D. J. Goodman and C. E. Sundberg. Combined source and channel coding for variable-bit-rate speech transmission. The Bell System Technical Journal, 62(7):2017–2036, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-7-2017.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-7-2017.pdf.

Goodman:1983:TEF

[GS83b] D. J. Goodman and C. E.

Sundberg. Transmission errors and forward error correction in embedded differential pulse code modulation. The Bell System Technical Journal, 62(9):2735–2764, November 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-9-2735.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-9-2735.pdf.

Gewirtz:1988:DND

[GS88] William L. Gewirtz and Peter H. Stuntebeck. Data networking directions. AT&T Technical Journal, 67(6):7–22, November 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Goksel:1986:TTV

[GST86] Ahmet Kemal Goksel, Warren T. Sekino, and William W. Troutman. Tools and techniques for VLSI quality. AT&T Technical Journal, 65(2):77–84, March/ April 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Gitlin:1981:FSE

[GW81] R. D. Gitlin and S. B. Weinstein. Fractionally-spaced equalization: An improved digital transversal equalizer. The Bell System Technical Journal, 60(2):275–296, February 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-2-275.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-2-275.pdf.

Grassman:1983:MSE

[GY83] T. E. Grassman and J. E.

Yates. Modernization of the suburban ESS: Overview: Evolution of the suburban ESS. The Bell System Technical Journal, 62(6):1455–1466, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1455.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1455.pdf.

Greenstein:1985:SSS

[GY85] L. J. Greenstein and Y. S. Yeh.

A simulation study of space diversity and adaptive equalization in microwave digital radio. AT&T Technical Journal, 64(4):907–935, 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Hagelbarger:1983:ETH

[HAK83] D. W. Hagelbarger, R. V. Anderson, and P. S. Kubik. Experimental teleterminals — hardware. The Bell System Technical Journal, 62(1):145–152, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-1-145.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-1-145.pdf.

Halfin:1983:BDV

[Hal83] S. Halfin. Batch delays versus customer delays. The Bell System Technical Journal, 62(7):2011–2015, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-7-2011.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-7-2011.pdf.

Hanson:1983:HFB

[Han83] B. L. Hanson. Human factors and behavioral science: A brief history of applied behavioral science at Bell Laboratories. The Bell System Technical Journal, 62(6):1571–1590, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-6-1571.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-6-1571.pdf.

Haskell:1981:TFM

[Has81] B. G. Haskell. Time-frequency multiplexing (TFM) of two NTSC color TV signals — simulation results. The Bell System Technical Journal, 60 (5):643–660, May/June 1981.

CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-5-643.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-5-643.pdf.

Hatamian:1985:HBV

[HB85] Mehdi Hatamian and Edward G.

Bowen. Homenet: a broadband voice/data/video network on CATV systems. AT&T Technical Journal, 64(2 part 1):347– 367, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Hirschberg:1988:NLP

[HBH88] Julia Hirschberg, Bruce W. Ballard, and Donald Hindle. Natural language processing. AT&T Technical Journal, 67(1):41– 57, January 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Heidt:1983:ASS

[HCH^{+}83] R. C. Heidt, E. F. Cook, R. P.

Hecken, R. W. Judkins, J. M. Kiker, Jr., F. J. Provenzano, Jr., and H. C. Wang. The AR6A single-sideband microwave radio system: Radio transmitterreceiver units. The Bell System Technical Journal, 62(10): 3337–3356, December 1983. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol62/bstj62-10-3337.pdf;

                                        http://www.alcatel-lucent.
                                        com/bstj/vol62-1983/articles/
                                        bstj62-10-3337.pdf.

Healy:1981:EME

[Hea81] J. D. Healy. The effects of misclassification error on the estimation of several population proportions. The Bell System Technical Journal, 60 (5):697–705, May/June 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-5-697.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-5-697.pdf.

Heffes:1980:CDT

[Hef80] H. Heffes. A class of data traffic processes — covariance function characterization and related queuing results. The Bell System Technical Journal, 59(6):897–929, July/August 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-6-897.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-6-897.pdf.

Heffes:1982:MFC

[Hef82] H. Heffes. Moment formulae for a class of mixed multi-job-type queueing networks. The Bell System Technical Journal, 61(5):709–745, May/June 1982. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-5-709.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-5-709.pdf.

Heiber:1989:OUI

[Hei89] Arnold Heiber. Overview of universal information services: concepts and technologies of future networks. AT&T Technical Journal, 68(2):5–13, 1989. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Henry:1981:BJB

[Hen81] P. S. Henry. B.S.T.J. briefs:

Fast decryption algorithm for the knapsack cryptographic system. The Bell System Technical Journal, 60(5):767–773, May/ June 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-5-767.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-5-767.pdf.

Henry:1984:FSS

[Hen84] Gary J. Henry. Fair Share Scheduler. AT&T Bell Laboratories Technical Journal, 63(8 part 2): 1845–1857, October 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Heyman:1982:ACS

[Hey82] D. P. Heyman. An analysis of the carrier-sense multiple-access protocol. The Bell System Technical Journal, 61(8):2023–2051, October 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-8-2023.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-8-2023.pdf.

Heyman:1983:DTP

[Hey83] D. P. Heyman. Data-transport performance analysis of fasnet. The Bell System Technical Journal, 62(8):2547–2560, October 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-8-2547.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-8-2547.pdf.

Hakki:1985:MML

[HFE85] B. W. Hakki, P. E. Fraley, and T. F. Eltringham. 1.3μm laser reliability determination for submarine cable systems. AT&T Technical Journal, 64(3):771–807, 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Henry:1981:NAH

[HG81] P. S. Henry and B. S. Glance. A new approach to high-capacity digital mobile radio. The Bell System Technical Journal, 60(8):1891–1904, October 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-8-1891.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-8-1891.pdf.

Hoover:1987:TI

[HHC87] Charles W. Hoover, Jr., William L.

Harrod, and Melvin I. Cohen. Technology of interconnection. AT&T Technical Journal, 66(4): 2–12, 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Hakki:1983:HFI

[HHG83] B. W. Hakki, W. R. Holbrook, and C. A. Gaw. Highfrequency impedance of protonbombarded injection lasers. The Bell System Technical Journal, 62(2):463–475, February 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-2-463.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-2-463.pdf.

Hatamian:1987:FII

[HHL^{+}87] Mehdi Hatamian, Lawrence A.

Hornak, Trevor E. Little, Stuart K. Tewksbury, and Paul Franzon. Fundamental interconnection issues [ICS]. AT&T Technical Journal, 66(4):13–30, 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Housos:1989:PAA

[HHL89] Efthymios C. Housos, Chih Chung Huang, and Jun-Min Liu. Parallel algorithms for the AT&T KORBX system. AT&T Technical Journal, 68(3):37–47, 1989. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (elec-

                                        tronic).  URL http://www.
                                        osti.gov/energycitations/
                                        product.biblio.jsp?query_
                                        id=6&page=0&osti_id=5280793.

Hack:1983:TSP

[HHS83a] T. G. Hack, T. Huang, and L. C. Stecher. Traffic service position system no. 1B: Software development system. The Bell System Technical Journal, 62(3):859–884, March 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-3-859.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-3-859.pdf.

Hecken:1983:ASS

[HHS83b] R. P. Hecken, R. C. Heidt, and D. E. Sanford. The AR6A single-sideband microwave radio system: Predistortion for the traveling-wave-tube amplifier. The Bell System Technical Journal, 62(10):3447–3464, December 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol62/bstj62-10-3447.pdf;
http://www.alcatel-lucent.
com/bstj/vol62-1983/articles/
bstj62-10-3447.pdf.

Hinton:1987:PST

[Hin87]  H. Scott Hinton.  Photonic

switching technology applications. AT&T Technical Journal, 66(3):41–53, May/June 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Hac:1988:DLB

[HJ88] A. Hac and T. J. Johnson. Dynamic load balancing through process and read-site placement in a distributed system. AT&T Technical Journal, 67(5):72–85, September/October 1988. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Howard:1988:ENN

[HJG88] R. E. Howard, L. D. Jackel, and H. P. Graf. Electronic neural networks. AT&T Technical Journal, 67(1):58–64, January 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Hetherington:1983:PAD

[HK83] I. K. Hetherington and P. Kusulas. The 3B20D processor & DMERT operating system: 3B20D processor memory systems. The Bell System Technical Journal, 62(1):207–220, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-1-207.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-1-207.pdf.

Hertz:1985:TLC

[HKMP85] David Hertz, Robert P. Kurshan, David Malah, and John T. Peoples. Tone location by cyclotomic filters. AT&T Technical Journal, 64(6 part 1):1161– 1180, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Hall:1983:TND

[HKRS83] M. S. Hall, Jr., J. A. Kohut, G. W. Riesz, and J. W. Steifle. Total network data system: System plan. The Bell System Technical Journal, 62(7):2147–2182, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-7-2147.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-7-2147.pdf.

Healy:1984:EOC

[HLL^{+}84] John D. Healy, Maurice Lampell, David G. Leeper, Thomas C. Redman, and Edward J. Vlacich. 1982/83 End Office Connection Study: ASPEN data acquisition system and sampling plan. AT&T Bell Laboratories Technical Journal, 63(9):2033– 2057, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Haugk:1985:SSM

[HLRW85] G. Haugk, F. M. Lax, R. D.

Royer, and J. R. Williams. The 5ESS switching system: maintenance capabilities. AT&T Technical Journal, 64(6 part 1):1385– 1416, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Hinderks:1980:CCM

[HM80] L. W. Hinderks and A. Maione.

Copper conductivity at millimeterwave frequencies. The Bell System Technical Journal, 59(1):43– 65, January 1980. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol59/bstj59-1-
43.pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-1-43.pdf.

Hayward:1983:TND

[HM83] W. S. Hayward and J. P. Moreland. Total network data system: Theoretical and engineering foundations. The Bell System Technical Journal, 62(7):2183–2207, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-7-2183.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-7-2183.pdf.

Hoppner:1981:NED

[HMPV81] K. M. Hoppner, H. Mann, S. F. Panyko, and J. VanZweden. No. 4 ESS: Digital interface. The Bell System Technical Journal, 60(6):1131–1166, July/ August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-6-1131.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-6-1131.pdf.

Horing:1982:SPC

[HMSY82] S. Horing, J. Z. Menard, R. E. Staehler, and B. J. Yokelson. Stored program controlled network: Overview. The Bell System Technical Journal, 61(7):1579–1589, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-7-1579.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-7-1579.pdf.

Hoadley:1981:QMP

[Hoa81]  Bruce Hoadley.  The qual-

ity measurement plan (QMP). The Bell System Technical Journal, 60(2):215–273, February 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol60/bstj60-2-215.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-2-215.pdf.

Holtman:1982:ARS

[Hol82] J. P. Holtman. Automated repair service bureau: The context-sensitive switch of the loop maintenance operations system. The Bell System Technical Journal, 61(6):1197–1208, July/ August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-6-1197.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-6-1197.pdf.

Holmgren:1983:HFB

[Hol83] J. E. Holmgren. Human factors and behavioral science: Toward Bell System applications of automatic speech recognition. The Bell System Technical Journal, 62(6):1865–1880, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1865.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1865.pdf.

Holzmann:1985:TP

[Hol85] G. J. Holzmann. Tracing protocols. AT&T Technical Journal, 64(10):2413–2433, 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Holzmann:1987:PPE

[Hol87] G. J. Holzmann. Pico — a picture editor. AT&T Technical Journal, 66(2):2–13, 1987. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Honig:1983:RFO

[Hon83] M. L. Honig. Recursive fixedorder covariance least-squares algorithms. The Bell System Technical Journal, 62(10):2961– 2992, December 1983. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol62/bstj62-10-2961.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol62-1983/articles/
                                        bstj62-10-2961.pdf.

Honig:1984:ATN

[Hon84] Michael L. Honig. Analysis of a TDMA network with voice and data traffic. AT&T Bell Laboratories Technical Journal, 63(8 part 1):1537–1563, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Haglund:1983:PAD

[HP83a] R. E. Haglund and L. D. Peterson. The 3B20D processor & DMERT operating system: 3B20D file memory systems. The Bell System Technical Journal, 62(1):235–254, January 1983. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-1-235.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-1-235.pdf.

Harvey:1983:ASS

[HP83b] S. A. Harvey and P. D. Patel.

The AR6A single-sideband microwave radio system: Radioline physical design. The Bell System Technical Journal, 62 (10):3313–3335, December 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol62/bstj62-10-3313.pdf;
http://www.alcatel-lucent.
com/bstj/vol62-1983/articles/
bstj62-10-3313.pdf.

Higdon:1986:PIS

[HPS86] Marda Higdon, Judith T. Page, and Peter Stuntebeck. Planning for ISDN in the 5ESS switch. AT&T Technical Journal, 65(1): 27–34, January/February 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Hester:1981:UDS

[HPU81] S. D. Hester, D. L. Parnas, and D. F. Utter. Using documentation as a software design medium. The Bell System Technical Journal, 60(8):1941–1977, October 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-8-1941.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-8-1941.pdf.

Hansen:1983:PAD

[HPW83] R. C. Hansen, R. W. Peterson, and N. O. Whittington. The 3B20D processor & DMERT operating system: Fault detection and recovery. The Bell System Technical Journal, 62(1):349–365, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-1-349.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-1-349.pdf.

Haas:1982:SPC

[HSS82] C. W. Haas, D. C. Salerno, and D. Sheinbein. Stored program controlled network: 800 service using SPC network capability — network implementation and administrative functions. The Bell System Technical Journal, 61(7):1745–1757, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-7-1745.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-7-1745.pdf.

Hluchyj:1983:PAP

[HTB83] M. G. Hluchyj, C. D. Tsao, and R. R. Boorstyn. Performance analysis of a preemptive priority queue with applications to packet communication systems. The Bell System Technical Journal, 62(10): 3225–3245, December 1983. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol62/bstj62-10-3225.pdf;
http://www.alcatel-lucent.
com/bstj/vol62-1983/articles/
bstj62-10-3225.pdf.

Israelski:1989:GUI

[IABB^{+}89] E. W. Israelski, J. S. Angiolillo-Bent, D. J. Brems, L. L. Hoag, L. A. Roberts, and R. S. Wells. Generalizable user-interface research. AT&T Technical Journal, 68(5):31–43, September/October 1989. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Ionescu-Graff:1982:SPA

[IG82] A. Ionescu-Graff. A sequential projection algorithm for special-services demand. The Bell System Technical Journal, 61(1):39–66, January 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-1-39.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-1-39.pdf.

Infosino:1980:RBD

[Inf80] W. J. Infosino. Relationships between the demand for local telephone calls and household characteristics. The Bell System Technical Journal, 59(6):931–953, July/August 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-6-931.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-6-931.pdf.

Inglis:1986:SSQ

[Ing86] James Inglis. Standard software quality metrics. AT&T Technical Journal, 65(2):113–118, March/ April 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Ishman:1983:MSE

[IST^{+}83] C. E. Ishman, R. B. Sanderson, L. M. Taff, D. P. Truax, and C. T. Tulloss. Modernization of the suburban ESS: Adding data links to an existing ESS. The Bell System Technical Journal, 62(6):1467–1495, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-6-1467.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-6-1467.pdf.

Jagerman:1982:AMW

[Jag82a] D. L. Jagerman. Approximate mean waiting times in transient GI/G/1 queues. The Bell System Technical Journal, 61(8):2003–2022, October 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-8-2003.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-8-2003.pdf.

Jagerman:1982:ITL

[Jag82b] D. L. Jagerman. An inversion technique for the Laplace transform. The Bell System Technical Journal, 61(8):1995–2002, October 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-8-1995.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-8-1995.pdf.

Jagerman:1984:MTC

[Jag84] David L. Jagerman. Methods in traffic calculations. AT&T Bell Laboratories Technical Journal, 63(7):1283–1310, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Jagerman:1985:LTI

[Jag85a] D. L. Jagerman. Laplace transform inequalities with application to queueing. AT&T Technical Journal, 64(7):1755– 1764, September 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Jagerman:1985:WTC

[Jag85b] David L. Jagerman. Waiting time convexity in the M/G/1 queue. AT&T Technical Journal, 64(1 part 1):33–41, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Jayant:1981:APF

[Jay81a] N. S. Jayant. Adaptive postfiltering of ADPCM speech. The Bell System Technical Journal, 60(5):707–717, May/June 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-5-707.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-5-707.pdf.

Jayant:1981:SDS

[Jay81b] N. S. Jayant. Subsampling of a DPCM speech channel to provide two ‘self-contained’ half-rate channels. The Bell System Technical Journal, 60(4):501–509, April 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-4-501.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-4-501.pdf.

Jayant:1983:VRA

[Jay83] N. S. Jayant. Variable rate AD-PCM based on explicit noise coding. The Bell System Technical Journal, 62(3):657–677, March 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-3-657.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-3-657.pdf.

Jayant:1987:ENK

[Jay87] N. S. Jayant. Effective number of keys in a voice privacy system based on permutation scrambling. AT&T Technical Journal, 66(2):132–136, January 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Julesz:1983:HFB

[JB83] B. Julesz and J. R. Bergen.

Human factors and behavioral science: Textons, the fundamental elements in preattentive vision and perception of textures. The Bell System Technical Journal, 62(6):1619–1645, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1619.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1619.pdf.

Jayant:1980:BJB

[JC80] N. S. Jayant and S. W. Christensen. B.S.T.J. briefs: Adaptive aperture coding for speech waveforms — II. The Bell System Technical Journal, 59(3):471–477, March 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-3-471.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-3-471.pdf.

Jayant:1983:ASS

[JCMQ83] N. S. Jayant, R. V. Cox, B. J. McDermott, and A. M. Quinn. Analog scramblers for speech based on sequential permutations in time and frequency. The Bell System Technical Journal, 62(1):25–46, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-1-25.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-1-25.pdf.

Jang:1982:BJB

[JCMS82] S. J. Jang, L. G. Cohen, W. L. Mammel, and M. A. Saifi. B.S.T.J. briefs: Experimental verification of ultrawide bandwidth spectra in double-clad single-mode fiber. The Bell System Technical Journal, 61(3):385–390, March 1982. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-3-385.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-3-385.pdf.

Jordan:1980:TAD

[JCV80] A. S. Jordan, R. Caruso, and A. R. VonNeida. A thermoelastic analysis of dislocation generation in pulled GaAs crystals. The Bell System Technical Journal, 59(4):593–637, April 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-4-593.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-4-593.pdf.

Jordan:1983:ADW

[JCV83] A. S. Jordan, R. Caruso, and A. R. VonNeida. An analysis of the derivative weightgain signal from measured crystal shape: Implications of diameter control of GaAs. The Bell System Technical Journal, 62(2):477–498, February 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-2-477.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-2-477.pdf.

Johnston:1989:LVP

[JDW89] W. Dexter Johnston, Jr., Michael A. DiGiuseppe, and Daniel P. Wilt. Liquid and vapor phase growth of III–V materials for photonic devices. AT&T Technical Journal, 68(1):53–63, January/February 1989. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Jeschke:1982:NRS

[JKT82] C. E. Jeschke, A. E. Kranenborg, and B. S. Thakkar. No. 10A remote switching system: Physical design. The Bell System Technical Journal, 61(4):525–564, April 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-4-525.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-4-525.pdf.

Joyce:1985:MAA

[JLN^{+}85] W. B. Joyce, K.-Y. Liou, F. R. Nash, P. R. Bossard, and R. L. Hartman. Methodology of accelerated aging [semiconductor devices]. AT&T Technical Journal, 64(3):717–764, 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Josenhans:1986:SPA

[JLR^{+}86] James G. Josenhans, John F. Lynch, Jr., Marian R. Rogers, Richard R. Rosinski, and Wendy P. VanDame. Speech processing application standards. AT&T Technical Journal, 65(5): 23–33, September/October 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Johnsen:1981:ECF

[JN81] O. Johnsen and A. N. Netravali.

An extension of the CCITT facsimile codes for dithered pictures. The Bell System Technical Journal, 60(3):391–404, March 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-3-391.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-3-391.pdf.

Johnsen:1981:NCT

[Joh81] O. Johnsen. A new code for transmission of ordered dithered pictures. The Bell System Technical Journal, 60(3):379–389, March 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-3-379.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-3-379.pdf.

Jones:1988:PCA

[Jon88]  Mark A. Jones.  Program-

ming connectionist architectures. AT&T Technical Journal, 67(1): 65–68, January 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Jablonowski:1987:OFM

[JPW87] Donald P. Jablonowski, Un Chul Pack, and Laurence S. Watkins. Optical fiber manufacturing techniques. AT&T Technical Journal, 66(1):33–44, January 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Janac:1982:DLE

[JS82] K. Janac and N. J. A. Sloane. The detection of long error bursts during transmission of video signals. The Bell System Technical Journal, 61(8):1911–1917, October 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-8-1911.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-8-1911.pdf.

Johnsen:1983:CTL

[JSC83] O. Johnsen, J. Segen, and G. L. Cash. Coding of twolevel pictures by pattern matching and substitution. The Bell System Technical Journal, 62(8):2513–2545, October 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-8-2513.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol62-1983/
articles/bstj62-8-2513.pdf.

Juang:1984:HMM

[Jua84a] Biing-Hwang Juang. On hidden Markov model and dynamic time warping for speech recognition — a unified view. AT&T Bell Laboratories Technical Journal, 63(7):1213–1243, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Juang:1984:UIS

[Jua84b] Biing-Hwang Juang. On using the Itakura–Saito measures for speech coder performance evaluation. AT&T Bell Laboratories Technical Journal, 63(8 part 1):1477–1498, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Juang:1985:MLE

[Jua85]  B.-H. Juang.  Maximum-

likelihood estimation for mixture multivariate stochastic observations of Markov chains. AT&T Technical Journal, 64(6 part 1):1235–1249, 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kane:1983:PAD

[KAM83] J. R. Kane, R. E. Anderson, and P. S. McCabe. The 3B20D processor & DMERT operating system: Overview, architecture, and performance of DMERT. The Bell System Technical Journal, 62(1):291–301, January 1983. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-1-291.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-1-291.pdf.

Kaminow:1989:PMAb

[Kam89a] I. P. Kaminow. Photonic multiple-access networks: routing and multiplexing. AT&T Technical Journal, 68(2):72–86, 1989. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kaminow:1989:PMAa

[Kam89b] I. P. Kaminow. Photonic multiple-access networks: topologies. AT&T Technical Journal, 68(2):61–71, 1989. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Karmakar:1980:SSS

[Kar80a] S. B. Karmakar. Steady-state stability of a synchronous machine. The Bell System Technical Journal, 59(8):1357–1364, October 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-8-1357.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-8-1357.pdf.

Kartalopoulos:1980:MIF

[Kar80b] S. V. Kartalopoulos. A memory implementation of a fully programmable digital controller. The Bell System Technical Journal, 59(9):1599–1607, November 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-9-1599.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-9-1599.pdf.

Karabinis:1983:MPA

[Kar83] P. D. Karabinis. Maximumpower and amplitude-equalizing algorithms for phase control in Space diversity combining. The Bell System Technical Journal, 62(1):63–89, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-1-63.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-1-63.pdf.

Karim:1986:PCM

[Kar86] M. R. Karim. Packet communications on a mobile radio channel. AT&T Technical Journal, 65(3):12–20, 1986. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kasper:1982:EMO

[Kas82] B. L. Kasper. Equalization of multimode optical fiber systems. The Bell System Technical Journal, 61(7):1367–1388, September 1982. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-7-1367.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-7-1367.pdf.

Kavehrad:1984:PCP

[Kav84] M. Kavehrad. Performance of cross-polarized M-ary QAM signals over nondispersive fading channels. AT&T Bell Laboratories Technical Journal, 63 (3):499–521, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Kavehrad:1985:BCP

[Kav85a] M. Kavehrad. Baseband cross-polarization interference cancellation for m-quadrature amplitude-modulated signals over multipath fading radio channels. AT&T Technical Journal, 64(8):1913–1926, 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kavehrad:1985:CCH

[Kav85b] Mohsen Kavehrad. Convolutional coding for high-speed microwave radio communications. AT&T Technical Journal, 64(7): 1625–1637, September 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kulpa:1983:PAD

[KBF83] S. H. Kulpa, J. M. Brown, and A. W. Fulton. The 3B20D processor & DMERT operating system: 3B20D packaging and technology. The Bell System Technical Journal, 62(1):221–234, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-1-221.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-1-221.pdf.

Kuzmak:1988:KBS

[KBG^{+}88] Sylvia D. Kuzmak, Mark D. Brittingham, Allen L. Gorin, Gregory A. Milich, and Joseph E. Shoenfelt. Knowledge-based signal interpretation. AT&T Technical Journal, 67(1):104– 120, January 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Karp:1988:SDN

[KBN88] Bennett C. Karp, L. Kirk Barker, and Larry D. Nelson. The Secure Data Network System. AT&T Technical Journal, 67(3):19–27, May 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Keramidas:1980:PFR

[KBZ80] V. G. Keramidas, G. W. Berkstresser, and C. L. Zipfel. Planar, fast, reliable, singleheterojunction light-emitting diodes for optical links. The Bell System Technical Journal, 59(9):1549–1557, November 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-9-1549.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-9-1549.pdf.

Kalish:1987:SMF

[KC87] David Kalish and Leonard G. Cohen. Single-mode fiber: From research and development to manufacturing. AT&T Technical Journal, 66(1):19–32, January 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kapilow:1988:QSA

[KC88] Sharon A. Kapilow and Mikhail Cherepov. Quest — a security auditing tool. AT&T Technical Journal, 67(3):65–71, May 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kaplan:1984:ECR

[KD84] Daniel R. Kaplan and Peter P. Deimel. Exact calculation of the reflection coefficient for coated optical waveguide devices. AT&T Bell Laboratories Technical Journal, 63 (6 part 1):857–877, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Kernighan:1989:USD

[Ker89] B. W. Kernighan. The UNIX system document preparation tools: A retrospective. AT&T Technical Journal, 68(4):5– 20, August 1989. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

King:1983:ESA

[KFS83] B. G. King, P. J. Fitzgerald, and H. A. Stein. An experimental study of atmospheric optical transmission. The Bell System Technical Journal, 62(3):607–629, March 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-3-607.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-3-607.pdf.

Kaufman:1982:EPN

[KK82] R. L. Kaufman and A. J. M.

Kester. Evaluation of private networks. The Bell System Technical Journal, 61(9): 2143–2166, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2143.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2143.pdf.

Kolb:1983:PVT

[KKLS83] E. D. Kolb, P. L. Key, R. A. Laudise, and E. E. Simpson. Pressure-volume-temperature behavior in the system H_{2}O–NaOH–SiO_{2} and its relationship to the hydrothermal growth of quartz. The Bell System Technical Journal, 62(3):639–656, March 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-3-639.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-3-639.pdf.

Klincewicz:1985:LSD

[Kli85] John G. Klincewicz. A largescale distribution and location model. AT&T Technical Journal, 64(7):1705–1730, September 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kaup:1981:TCC

[KLRS81] T. C. Kaup, D. G. Leeper, A. K. Reilly, and P. E. Scheffler. T1 carrier characterization — field-measurement results. The Bell System Technical Journal, 60(6):965–1003, July/ August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-6-965.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-6-965.pdf.

Karmarkar:1989:PSV

[KLSW89] Narendra K. Karmarkar, Jeffrey C. Lagarias, Lev Slutsman, and Pyng Wang. Power series variants of Karmarkar-type algorithms. AT&T Technical Journal, 68(3):20–36, 1989. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Klinksiek:1983:PAD

[KM83] W. F. Klinksiek and H. L.

Mitchell. The 3B20D processor & DMERT operating system: Integration and system test. The Bell System Technical Journal, 62(1):399–410, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-1-399.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-1-399.pdf.

Kavehrad:1985:PLC

[KM85] M. Kavehrad and P. J. McLane.

Performance of low-complexity channel coding and diversity for spread spectrum in indoor, wireless communication. AT&T Technical Journal, 64(8):1927– 1965, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Keeve:1982:NRS

[KMM82] F. H. Keeve, J. C. Martin, and T. L. McRoberts. No. 10A remote switching system: System maintenance. The Bell System Technical Journal, 61(4):597–625, April 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-4-597.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-4-597.pdf.

Knerr:1980:SST

[Kne80]  R. H. Knerr.  Series- and

shunt-tuned lumped-element microwave circulators. The Bell System Technical Journal, 59(5):723–730, May/June 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-5-723.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-5-723.pdf.

Kohl:1983:HFB

[Koh83]  G. A. Kohl.  Human fac-

tors and behavioral science: Effects of shape and size of knobs on maximal hand-turning forces applied by females. The Bell System Technical Journal, 62(6):1705–1712, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-6-1705.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-6-1705.pdf.

Koontz:1980:EEL

[Koo80] W. L. G. Koontz. Economic evaluation of loop feeder relief alternatives. The Bell System Technical Journal, 59(3):277–293, March 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-3-277.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-3-277.pdf.

Kowalski:1988:VDA

[Kow88] Thaddeus J. Kowalski. The VLSI design automation assistant: a synthesis expert. AT&T Technical Journal, 67(1):81–92, January 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kiker:1983:ASS

[KP83] J. M. Kiker, Jr. and S. B. Pirkau.

The AR6A single-sideband microwave radio system: Frequency control. The Bell System Technical Journal, 62(10): 3377–3390, December 1983. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol62/bstj62-10-3377.pdf;
http://www.alcatel-lucent.
com/bstj/vol62-1983/articles/
bstj62-10-3377.pdf.

Kautz:1988:PPR

[KP88] Henry A. Kautz and Edwin P. D. Pednault. Planning and plan recognition. AT&T Technical Journal, 67(1):25–40, January 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kennedy:1985:PDE

[KPW85] T. S. Kennedy, D. A. Pezzutti, and T. L. Wang. Project development environment. AT&T Technical Journal, 64(1 part 2):269–285, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Khanna:1980:EET

[KR80] S. P. Khanna and R. L. Remke. The EL2 electret transmitter: Technology development. The Bell System Technical Journal, 59(5):745–762, May/June 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-5-745.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-5-745.pdf.

Kashper:1982:EDD

[KRS82] A. Kashper, S. M. Rocklin, and C. R. Szelag. Effects of day-to-day load variation on trunk group blocking. The Bell System Technical Journal, 61(2):123–135, February 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-2-123.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-2-123.pdf.

Karhan:1983:HFB

[KRS83] C. J. Karhan, C. A. Riley, and M. S. Schoeffler. Human factors and behavioral science: Designing and evaluating standard instructions for public telephones.

The Bell System Technical Journal, 62(6):1827–1847, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1827.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1827.pdf.

Kavehrad:1985:CPC

[KS85] Mohsen Kavehrad and Jack Salz. Cross-polarization cancellation and equalization in digital transmission over dually polarized multipath fading channels. AT&T Technical Journal, 64(10):2211–2245, 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kackar:1986:RDC

[KS86] Raghu N. Kackar and Anne C.

Shoemaker. Robust design: A cost effective method for improving manufacturing processes. AT&T Technical Journal, 65(2):39–50, March/April 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Karanam:1988:PEV

[KSB88] V. R. Karanam, K. Sriram, and Duane O. Bowker. Performance evaluation of variablebit-rate voice in packet-switched networks. AT&T Technical Journal, 67(5):57–71, September/October 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kaufman:1981:OMD

[KSM81] L. Kaufman, J. B. Seery, and J. A. Morrison. Overflow models for dimension PBX feature packages. The Bell System Technical Journal, 60(5):661–676, May/ June 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-5-661.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-5-661.pdf.

Kummer:1980:LSL

[Kum80]  R. B. Kummer.  Light-

guide splice loss — effects of launch beam numerical aperture. The Bell System Technical Journal, 59(3):441–447, March 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-3-441.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-3-441.pdf.

Kumar:1983:EQN

[Kum83] A. Kumar. Equivalent queueing networks and their use in approximate equilibrium analysis. The Bell System Technical Journal, 62(10):2893–2910, December 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/

Vol62/bstj62-10-2893.pdf;
http://www.alcatel-lucent.
com/bstj/vol62-1983/articles/
bstj62-10-2893.pdf.

Kumar:1988:SSP

[Kum88] A. Kumar. SNA/SDLC performance over ISDN frame-relay, virtual-circuit data networks. AT&T Technical Journal, 67(5): 27–40, September/October 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kurth:1980:TPA

[Kur80] C. F. Kurth. Two-port analysis of SC networks with continuous input signals. The Bell System Technical Journal, 59(8):1297–1316, October 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-8-1297.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-8-1297.pdf.

Kurshan:1985:PSB

[Kur85] Robert P. Kurshan. Proposed specification of BX.25 link layer protocol. AT&T Technical Journal, 64(2 part 2):559–596, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kashper:1984:TIP

[KV84] A. N. Kashper and G. C. Varvaloucas. Trunk implementation plan for hierarchical networks.

AT&T Bell Laboratories Technical Journal, 63(1):57–88, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Klincewicz:1984:AQI

[KW84] J. G. Klincewicz and W. Whitt. On approximations for queues, II: Shape constraints. AT&T Bell Laboratories Technical Journal, 63(1):139–161, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Lavin:1987:ORL

[Lav87] Joanne S. Lavin. Optimally rehoming local serving offices to a new point of presence. AT&T Technical Journal, 66(2): 50–54, 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Lin:1983:HPL

[LBHB83] Chinlon Lin, G. Beni, S. Hackwood, and T. J. Bridges. High-power lasers and optical waveguides for robotic materialprocessing applications. The Bell System Technical Journal, 62(8):2479–2492, October 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-8-2479.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-8-2479.pdf.

Lin:1980:RAE

[LBP80] S. H. Lin, H. J. Bergmann, and M. V. Pursley. Rain attenuation on Earth-satellite paths — summary of 10-year experiments and studies. The Bell System Technical Journal, 59(2):183–228, February 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-2-183.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-2-183.pdf.

Landauer:1982:DSH

[LDGF82] T. K. Landauer, S. T. Dumais, L. M. Gomez, and G. W. Furnas. Database systems: Human factors in a data access. The Bell System Technical Journal, 61(9):2487–2509, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2487.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2487.pdf.

Leck:1980:BJB

[Lec80] R. P. Leck. B.S.T.J. briefs: A microprocessor-based automatic frequency controller. The Bell System Technical Journal, 59 (10):2007–2010, December 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/

                                        Vol59/bstj59-10-2007.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol59-1980/articles/
                                        bstj59-10-2007.pdf.

Lee:1982:PLA

[Lee82]  G. S. Lee.  Piecewise lin-

ear approximation of multivariate functions. The Bell System Technical Journal, 61 (7):1463–1486, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-7-1463.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-7-1463.pdf.

Lee:1983:ATR

[Lee83] T. T. Lee. An algebraic theory of relational databases. The Bell System Technical Journal, 62 (10):3159–3204, December 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol62/bstj62-10-3159.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol62-1983/articles/
                                        bstj62-10-3159.pdf.

Lenahan:1983:CMO

[Len83]  T. A. Lenahan.  Calcula-

tion of modes in an optical fiber using the finite element method and EISPACK. The Bell System Technical Journal, 62(9):2663–2694, November 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-9-2663.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-9-2663.pdf.

Lenahan:1985:TBD

[Len85] T. A. Lenahan. Thermal buckling of dual-coated fiber. AT&T Technical Journal, 64(7):1565– 1584, September 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Leung:1986:DSC

[Leu86] W. H. Francis Leung. The distributed software concept of remoted processes, their application and implementation. AT&T Technical Journal, 65(3): 2–11, 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Limb:1982:DFU

[LF82] J. O. Limb and C. Flores.

Description of fasnet — A unidirectional local-area communications network. The Bell System Technical Journal, 61(7):1413–1440, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-7-1413.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-7-1413.pdf.

Lipske:1984:NZO

[LF84] K. R. Lipske and J. H. Fletcher. Nonlinear zero-one combinatorial, goal-programming model and constructive algorithm for solving multiobjective assignment problems. AT&T Bell Laboratories Technical Journal, 63(4):665–677, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Laker:1982:PIB

[LFG82] K. R. Laker, P. E. Fleischer, and A. Ganesan. Parasitic insensitive, biphase switched capacitor filters realized with one operational amplifier per pole pair. The Bell System Technical Journal, 61(5):685–707, May/ June 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-5-685.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-5-685.pdf.

Lilienthal:1988:FMW

[LFHV88] Peter F. Lilienthal II, J. Peter Flemming, E. G. Harokopos, and Glenn C. Van Orden. A flexible manufacturing workstation. AT&T Technical Journal, 67(2): 5–14, March/April 1988. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Loverde:1985:PAD

[LFLB85] A. S. Loverde, Howard D. Frisch, C. R. Lindemulder, and Donn Baker. Physical architecture and design. AT&T Technical Journal, 64(1 part 2):175–195, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Levinson:1984:SGL

[LG84] F. H. Levinson and S. W.

Granlund. Single GRIN-lens directional couplers. AT&T Bell Laboratories Technical Journal, 63(3):431–439, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Li:1987:ALS

[Li87] Tingye Li. Advances in lightwave systems research. AT&T Technical Journal, 66(1):5–18, January 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Liebesman:1982:CEM

[Lie82] B. S. Liebesman. The continuing evolution of the military standard 105D sampling system. The Bell System Technical Journal, 61(2):137–157, February 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-2-137.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-2-137.pdf.

Lin:1980:SBM

[Lin80] S. H. Lin. Statistical behavior of multipair crosstalk. The Bell System Technical Journal, 59(6):955–974, July/August 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-6-955.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-6-955.pdf.

Lin:1981:SBC

[Lin81] S. H. Lin. Statistical behavior of crosstalk power sum with dominant components. The Bell System Technical Journal, 60(7):1363–1374, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-7-1363.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-7-1363.pdf.

Linderman:1982:DSI

[Lin82] J. P. Linderman. Database systems: Issues in the design of a distributed record management system. The Bell System Technical Journal, 61(9):2555–2566, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2555.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2555.pdf.

Linderman:1984:TPC

[Lin84] John P. Linderman. Theory and practice in the construction of a working sort routine. AT&T Bell Laboratories Technical Journal, 63(8 part 2):1827–1843, October 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Liu:1980:DDD

[Liu80]  K. S. Liu.  Direct dis-

tance dialing: Call completion and customer retrial behavior. The Bell System Technical Journal, 59(3):295–311, March 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-3-295.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-3-295.pdf.

Luan:1988:EBM

[LL88] D. T. Luan and D. M. Lucantoni. The effect of bandwidth management on the performance of a window-based flow control. AT&T Technical Journal, 67(5): 17–26, September/October 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Lawser:1982:SPC

[LLS82] J. J. Lawser, R. E. LeCronier, and R. L. Simms. Stored program controlled network: Generic network plan. The Bell System Technical Journal, 61(7):1589–1598, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-7-1589.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-7-1589.pdf.

Levendel:1981:FSM

[LM81a] Y. H. Levendel and P. R. Menon. Fault-simulation methods — extensions and comparison. The Bell System Technical Journal, 60(9):2235–2258, November 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-9-2235.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-9-2235.pdf.

Lim:1981:AEP

[LM81b] T. L. Lim and M. S. Mueller. Adaptive equalization and phase tracking for simultaneous analog/digital data transmission. The Bell System Technical Journal, 60(9):2039–2063, November 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-9-2039.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-9-2039.pdf.

Levendel:1981:ALS

[LMM81] Y. L. Levendel, P. R. Menon, and C. E. Miller. Accurate logic simulation models for TTL totempole and MOS gates and tristate devices. The Bell System Technical Journal, 60(7):1271–1287, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-7-1271.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-7-1271.pdf.

Logan:1983:APG

[LMOS83] B. F. Logan, Jr., J. E. Mazo, A. M. Odlyzko, and L. A. Shepp. On the average product of Gauss–Markov variables. The Bell System Technical Journal, 62(10):2993–3006, December 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol62/bstj62-10-2993.pdf;
http://www.alcatel-lucent.
com/bstj/vol62-1983/articles/
bstj62-10-2993.pdf.

Levendel:1982:SPC

[LMP82] Y. H. Levendel, P. R. Menon, and S. H. Patel. Special-purpose computer for logic simulation using distributed processing. The Bell System Technical Journal, 61(10):2873–2909, December 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol61/bstj61-10-2873.pdf;

                                        http://www.alcatel-lucent.
                                        com/bstj/vol61-1982/articles/
                                        bstj61-10-2873.pdf.

Levendel:1983:PFS

[LMP83] Y. H. Levendel, P. R. Menon, and S. H. Patel. Parallel fault simulation using distributed processing. The Bell System Technical Journal, 62(10):3107– 3137, December 1983. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol62/bstj62-10-3107.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol62-1983/articles/
                                        bstj62-10-3107.pdf.

Lee:1989:SLT

[LMS^{+}89] Dah Nain Lee, Karen T. Medhi, John L. Strand, Roger G. Cox, and Stephen Chen. Solving large telecommunications network loading problems. AT&T Technical Journal, 68(3):48–56, 1989. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Lo:1987:PAA

[Lo87] Chung-Yin Lo. Performance analysis and application of a twopriority packet queue. AT&T Technical Journal, 66(3):82– 99, May/June 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Loeb:1983:HFB

[Loe83]  K. M. Cohen Loeb.  Hu-

man factors and behavioral science: Membrane keyboards and human performance. The Bell System Technical Journal, 62(6):1733–1749, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1733.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1733.pdf.

Logan:1983:BCI

[Log83a] B. F. Logan, Jr. Bandwidthconserving independent amplitude and phase modulation. The Bell System Technical Journal, 62(10):3053–3062, December 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol62/bstj62-10-3053.pdf;
http://www.alcatel-lucent.
com/bstj/vol62-1983/articles/
bstj62-10-3053.pdf.

Logan:1983:SSC

[Log83b] B. F. Logan, Jr. Series solutions of companding problems. The Bell System Technical Journal, 62(10):3007–3052, December 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol62/bstj62-10-3007.pdf;
http://www.alcatel-lucent.
com/bstj/vol62-1983/articles/
bstj62-10-3007.pdf.

Logan:1984:CM

[Log84a] B. F. Logan, Jr. Click modulation. AT&T Bell Laboratories Technical Journal, 63(3):401– 423, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Logan:1984:SDRa

[Log84b] B. F. Logan, Jr. Signals designed for recovery after clipping — I. localization of infinite products. AT&T Bell Laboratories Technical Journal, 63(2):261–285, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Logan:1984:SDRb

[Log84c] B. F. Logan, Jr. Signals designed for recovery after clipping — II. Fourier transform theory of recovery. AT&T Bell Laboratories Technical Journal, 63(2):287–306, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Logan:1984:SDR

[Log84d] B. F. Logan, Jr. Signals designed for recovery after clipping — III: generalizations. AT&T Bell Laboratories Technical Journal, 63(3):379–399, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Logan:1985:BEE

[Log85] Benjamin F. Logan, Jr. Bandwidtherror exchange for a simple fading channel model. AT&T Technical Journal, 64(7):1687– 1704, September 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Longhitano:1980:SAS

[Lon80] R. Longhitano. Sensitivity analysis of the 47A signal. The Bell System Technical Journal, 59(9):1747–1756, November 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-9-1747.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-9-1747.pdf.

Luniewicz:1984:SSL

[LOS84] Michael M. Luniewicz, John W.

Olson, and Kenneth E. Stiefel. SLC 96 subscriber loop carrier system: channel bank. AT&T Bell Laboratories Technical Journal, 63(10 part 2):2283–2331, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Levinson:1983:IAT

[LRS83] S. E. Levinson, L. R. Rabiner, and M. M. Sondhi. An introduction to the application of the theory of probabilistic functions of a Markov process to automatic speech recognition. The Bell System Technical Journal, 62(4):1035–1074, April 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-4-1035.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-4-1035.pdf.

Lawrence:1980:FPD

[LS80a] V. B. Lawrence and A. C. Salazar. Finite precision design of linear-phase FIR filters. The Bell System Technical Journal, 59(9):1575–1598, November 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-9-1575.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-9-1575.pdf.

Leland:1980:IMS

[LS80b] K. W. Leland and N. R. Sollenberger. Impairment mechanisms for SSB mobile communications at UHF with pilotbased Doppler/fading correction. The Bell System Technical Journal, 59(10):1923–1942, December 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol59/bstj59-10-1923.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol59-1980/articles/
                                        bstj59-10-1923.pdf.

Levinson:1980:CMA

[LS80c] S. E. Levinson and K. L. Shipley. A conversational-mode airline information and reservation system using speech input and output. The Bell System Technical Journal, 59(1):119–137, January 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-1-119.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-1-119.pdf.

Lehder:1988:SET

[LSY88] Wilfred E. Lehder, Jr., D. Paul Smith, and Weider D. Yu. Software estimation technology. AT&T Technical Journal, 67 (4):10–18, July 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Lin:1983:CDM

[LTT^{+}83] C. Lin, A. R. Tynes, A. Tomita, P. L. Liu, and D. L. Philen. Chromatic dispersion measurements in single-mode fibers using picosecond InGaAsP injection lasers in the 1.2- to 1.5μ m spectral region. The Bell System Technical Journal, 62(2):457–462, February 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-2-457.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-2-457.pdf.

Lodwig:1983:MSE

[LW83] J. P. Lodwig and D. A. Ward.

Modernization of the suburban ESS: Billing and measurements modernization. The Bell System Technical Journal, 62(6):1537–1550, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1537.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1537.pdf.

Leonard:1982:ARS

[LZ82] G. H. Leonard and J. E. Zielinski. Automated repair service bureau: Human performance design techniques. The Bell System Technical Journal, 61(6):1293–1299, July/ August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-6-1293.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-6-1293.pdf.

Macdonald:1983:HFB

[Mac83]  N. H. Macdonald.  Human

factors and behavioral science: The UNIX Writer’s Workbench software: Rationale and design. The Bell System Technical Journal, 62(6):1891–1908, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1891.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1891.pdf.

Manighalam:1989:INS

[Man89] S. Manighalam. Interconnection Network Simulator (INS). AT&T Technical Journal, 68(6): 63–72, 1989. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Martersteck:1981:NEP

[Mar81]  K. E. Martersteck.  No. 4

ESS: Prologue. The Bell System Technical Journal, 60(6): 1041–1048, July/August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-6-1041.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-6-1041.pdf.

Martin:1982:ARS

[Mar82] R. L. Martin. Automated repair service bureau: System architecture. The Bell System Technical Journal, 61(6):1115–1129, July/ August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-6-1115.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-6-1115.pdf.

Markle:1983:ASS

[Mar83] R. E. Markle. The AR6A singlesideband microwave radio system: Prologue. The Bell System Technical Journal, 62(10): 3249–3253, December 1983. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol62/bstj62-10-3249.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol62-1983/articles/
                                        bstj62-10-3249.pdf.

Martin:1984:USP

[Mar84] R. L. Martin. The UNIX system: Preface. AT&T Bell Laboratories Technical Journal, 63(8 part 2): 1571–1572, October 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Marques:1988:SNT

[Mar88] Todd E. Marques. StarKeeper network troubleshooter: An expert system product. AT&T Technical Journal, 67(6):137– 154, November 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Maxemchuk:1980:ESS

[Max80] N. F. Maxemchuk. An experimental speech storage and editing facility. The Bell System Technical Journal, 59 (8):1383–1395, October 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-8-1383.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-8-1383.pdf.

Maxemchuk:1982:VCC

[Max82] N. F. Maxemchuk. A variation on CSMA/CD that yields movable TDM slots in integrated voice/data local networks. The Bell System Technical Journal, 61(7):1527–1550, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-7-1527.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-7-1527.pdf.

Mazo:1980:ADD

[Maz80] J. E. Mazo. Analysis of decisiondirected equalizer convergence. The Bell System Technical Journal, 59(10):1857–1876, December 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol59/bstj59-10-1857.pdf;
http://www.alcatel-lucent.
com/bstj/vol59-1980/articles/
bstj59-10-1857.pdf.

Mazo:1982:SEM

[Maz82]  J. E. Mazo.  Some ex-

tremal Markov chains. The Bell System Technical Journal, 61(8):2065–2080, October 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-8-2065.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-8-2065.pdf.

Mazo:1985:BDD

[Maz85] James E. Mazo. On binary differential detection for coherent lightwave communication. AT&T Technical Journal, 64 (10):2467–2483, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Mitze:1983:PAD

[MBD^{+}83] R. W. Mitze, H. L. Bosco, N. X. DeLessio, R. J. Frank, N. A. Martellotto, W. C. Schwartz, and R. M. Wolfe. The 3B20D processor & DMERT operating system: 3B20D processor & DMERT as a base for telecommuncations applications. The Bell System Technical Journal, 62(1):171–179, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-1-171.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-1-171.pdf.

McGonegal:1981:DSP

[MBJ81] C. A. McGonegal, D. A. Berkley, and N. S. Jayant. Digital signal processor: Private communications. The Bell System Technical Journal, 60(7):1563–1572, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-7-1563.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol60-1981/
articles/bstj60-7-1563.pdf.

McCallum:1986:ORM

[McC86] Charles J. McCallum, Jr. Operations research in manufacturing. AT&T Technical Journal, 65 (4):4–16, July/August 1986. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

McNees:1980:TPM

[McN80] G. K. McNees. Transmission performance of Mu-255 quantization in a local digital office. The Bell System Technical Journal, 59 (10):1943–1964, December 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol59/bstj59-10-1943.pdf;
http://www.alcatel-lucent.
com/bstj/vol59-1980/articles/
bstj59-10-1943.pdf.

Miller:1982:WBL

[MDSL82] R. C. Miller, B. C. DeLoach, T. S. Stakelon, and R. B. Lawry. Wideband, bidirectional lightguide communication with an optically powered audio channel. The Bell System Technical Journal, 61(7):1359–1365, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-7-1359.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-7-1359.pdf.

Means:1980:CLR

[Mea80] D. R. Means. A class of lossless, reciprocal anti-sidetone networks for telephone sets. The Bell System Technical Journal, 59(8):1483–1492, October 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-8-1483.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-8-1483.pdf.

Meyers:1982:CDR

[Mey82] M. H. Meyers. Computing the distribution of a random variable via Gaussian quadrature rules. The Bell System Technical Journal, 61(9):2245–2261, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-9-2245.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-9-2245.pdf.

Malah:1981:FSS

[MF81] D. Malah and J. L. Flanagan. Frequency scaling of speech signals by transform techniques. The Bell System Technical Journal, 60(9):2107–2156, November 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-9-2107.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol60-1981/
articles/bstj60-9-2107.pdf.

Miller:1982:GDC

[MF82] J. E. Miller and O. Fujimura.

Graphic displays of combined presentations of acoustic and articulatory information. The Bell System Technical Journal, 61(5):799–810, May/June 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-5-799.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-5-799.pdf.

Minkoff:1984:WON

[Min84] J. B. Minkoff. Wideband operation of nonlinear solid-state power amplifiers — comparisons of calculations and measurements. AT&T Bell Laboratories Technical Journal, 63(2):231– 248, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Mark:1984:IVD

[ML84] John W. Mark and John O.

Limb. Integrated voice/data services on Fasnet. AT&T Bell Laboratories Technical Journal, 63(2):307–336, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

McKenna:1982:IRA

[MM82] J. McKenna and D. Mitra. Integral representations and asymptotic expansions for closed Markovian queueing networks: Normal usage. The Bell System Technical Journal, 61(5):661–683, May/June 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-5-661.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-5-661.pdf.

McFarland:1985:IAR

[MM85] M. A. McFarland and J. A. Miller. Introduction activities and results. AT&T Technical Journal, 64(1 part 2):321– 332, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

McKenna:1981:CCM

[MMR81] J. McKenna, D. Mitra, and K. G. Ramakrishnan. A class of closed Markovian queuing networks: Integral representations, asymptotic expansions, and generalizations. The Bell System Technical Journal, 60(5):599–641, May/ June 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-5-599.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-5-599.pdf.

McRoy:1984:SSL

[MMTV84] Steven A. McRoy, James H. Miller, James B. Truesdale, and Robert W. Van Slooten. SLC 96 subscriber loop carrier system: integration with the 5ESS switching system. AT&T Bell Laboratories Technical Journal, 63(10 part 2):2417–2437, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Morrison:1980:ASO

[Mor80a] J. A. Morrison. Analysis of some overflow problems with queuing. The Bell System Technical Journal, 59(8):1427–1462, October 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-8-1427.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-8-1427.pdf.

Morrison:1980:STO

[Mor80b] J. A. Morrison. Some traffic overflow problems with a large secondary queue. The Bell System Technical Journal, 59(8):1463–1482, October 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-8-1463.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-8-1463.pdf.

Morris:1981:PQN

[Mor81a] R. J. T. Morris. Priority queuing networks. The Bell System Technical Journal, 60(8):1745–1769, October 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-8-1745.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-8-1745.pdf.

Morrison:1981:OSW

[Mor81b] J. A. Morrison. An overflow system in which queuing takes precedence. The Bell System Technical Journal, 60(1):1–12, January 1981. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol60/bstj60-1-
                                        1.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-1-1.pdf.

Moreland:1982:RSP

[Mor82]  J. P. Moreland.  A ro-

bust sequential projection algorithm for traffic load forecasting. The Bell System Technical Journal, 61(1):15–38, January 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-1-15.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-1-15.pdf.

Myers:1981:CSS

[MR81] C. S. Myers and L. R. Rabiner. A comparative study of several dynamic time-warping algorithms for connected-word recognition.

The Bell System Technical Journal, 60(7):1389–1409, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-7-1389.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-7-1389.pdf.

Marple:1983:PFA

[MR83] S. L. Marple, Jr. and L. R. Rabiner. Performance of a fast algorithm for FIR system identification using least-squares analysis. The Bell System Technical Journal, 62(3):717–742, March 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-3-717.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-3-717.pdf.

Myers:1981:UDT

[MRR81] C. S. Myers, L. R. Rabiner, and A. E. Rosenberg. On the use of dynamic time warping for word spotting and connected word recognition. The Bell System Technical Journal, 60(3):303–325, March 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-3-303.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-3-303.pdf.

Metz:1981:NEN

[MRW81] R. Metz, E. L. Reible, and D. F. Winchell. No. 4 ESS: Network clock synchronization. The Bell System Technical Journal, 60(6):1109–1129, July/ August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-6-1109.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-6-1109.pdf.

Mueller:1981:UTD

[MS81] M. S. Mueller and J. Salz. A unified theory of data-aided equalization. The Bell System Technical Journal, 60(9): 2023–2038, November 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-9-2023.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-9-2023.pdf.

Monma:1982:PAF

[MS82] C. L. Monma and M. Segal. A primal algorithm for finding minimum-cost flows in capacitated networks with applications. The Bell System Technical Journal, 61(6):949–968, July/ August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-6-949.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol61-1982/
articles/bstj61-6-949.pdf.

Monma:1984:PAI

[MS84] C. L. Monma and D. R. Smith.

Probabilistic analysis of interframe tie requirements for crossconnect systems. AT&T Bell Laboratories Technical Journal, 63(4):643–664, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Melliar-Smith:1989:PED

[MS89] C. Mark Melliar-Smith. Photonic and electronic device technologies. AT&T Technical Journal, 68(1):2–4, January/February 1989. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Morgen:1984:SSL

[MSO84] Dennis H. Morgen, Michael A.

Schwartz, and John W. Olson. SLC 96 subscriber loop carrier system: maintenance and operation. AT&T Bell Laboratories Technical Journal, 63(10 part 2):2363–2387, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

McDonald:1980:DWP

[MT80] P. H. McDonald and T. J.

Thompson. Designer’s Workbench: The programmer environment. The Bell System Technical Journal, 59(9): 1793–1809, November 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-9-1793.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-9-1793.pdf.

Mueller:1981:LSA

[Mue81a] M. S. Mueller. Least-squares algorithms for adaptive equalizers. The Bell System Technical Journal, 60(8):1905–1925, October 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-8-1905.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-8-1905.pdf.

Mueller:1981:RIC

[Mue81b] M. S. Mueller. On the rapid initial convergence of least-squares equalizer adjustment algorithms. The Bell System Technical Journal, 60(10):2345–2358, December 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol60/bstj60-10-2345.pdf;
http://www.alcatel-lucent.
com/bstj/vol60-1981/articles/
bstj60-10-2345.pdf.

Morrison:1982:TOS

[MW82] J. A. Morrison and P. E. Wright. A traffic overflow system with A large primary queue. The Bell System Technical Journal, 61(7):1487–1517, September 1982. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-7-1487.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-7-1487.pdf.

Morrison:1983:SAM

[MY83] J. A. Morrison and W. W.

Yale. Stochastic analysis of mechanizing transaction data bases. The Bell System Technical Journal, 62(1):91–99, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-1-91.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-1-91.pdf.

Malek-Zavarei:1982:AGT

[MZ82]  M. Malek-Zavarei.  Applica-

tion of graph theory to the solution of a nonlinear optimal assignment problem. The Bell System Technical Journal, 61(8):1863–1870, October 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-8-1863.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-8-1863.pdf.

Malek-Zavarei:1984:GAD

[MZCW84] M. Malek-Zavarei, M. C. Chow, and J. D. Williams. Generic approaches to the design of network services circuits. AT&T Bell Laboratories Technical Journal, 63(5):737–773, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Nassau:1981:MDZ

[Nas81] K. Nassau. The material dispersion zero in infrared optical waveguide materials. The Bell System Technical Journal, 60(3):327–337, March 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-3-327.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-3-327.pdf.

Netravali:1982:IRD

[NB82] A. N. Netravali and E. G. Bowen. Improved reconstruction of DPCM-coded pictures. The Bell System Technical Journal, 61(6):969–979, July/August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-6-969.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-6-969.pdf.

Nair:1981:SSP

[ND81] V. N. Nair and T. E. Dalenius. Sampling from structured populations: Some issues and answers. The Bell System Technical Journal, 60(7):1235–1256, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-7-1235.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-7-1235.pdf.

Neal:1980:BDT

[Nea80] S. R. Neal. Blocking distributions for trunk network administration. The Bell System Technical Journal, 59(6):829–844, July/ August 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-6-829.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-6-829.pdf.

Nelson:1984:EII

[NMR^{+}84] T. J. Nelson, D. J. Muehlner, V. V. S. Rana, B. J. Roman, and G. P. Vella-Coleiro. Experimental ion-implanted bubble memory device with 16-μm^{2} cell. AT&T Bell Laboratories Technical Journal, 63(2):337–355, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Noe:1981:IFU

[Noe81] N. H. Noe. Impact of forecast uncertainty on feeder-cable sizing. The Bell System Technical Journal, 60(5):677–696, May/ June 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-5-677.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-5-677.pdf.

Netravali:1980:MCC

[NR80] A. N. Netravali and J. D. Robbins. Motion-compensated coding: Some new results. The Bell System Technical Journal, 59(9):1735–1745, November 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-9-1735.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-9-1735.pdf.

Nichols:1982:NRS

[NS82] R. K. Nichols and T. J. J. Starr. No. 10A remote switching system: Control-complex architecture and circuit design. The Bell System Technical Journal, 61(4):419–450, April 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-4-419.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-4-419.pdf.

Netravali:1985:AET

[NS85] A. N. Netravali and J. Salz. Algorithms for estimation of threedimensional motion. AT&T Technical Journal, 64(2 part 1):335–346, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Neigh:1986:RIA

[NS86] James L. Neigh and Leslie A.

Spindel. Role of ISDN in AT&T information systems architecture. AT&T Technical Journal, 65(1):45–55, January/February 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Nantz:1987:LAL

[NS87] Timothy D. Nantz and Walter J.

Shenk. Lightguide applications in the loop. AT&T Technical Journal, 66(1):108–118, January 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Nucho:1981:TBK

[Nuc81] R. N. Nucho. Transient behavior of the Kendall birth-death process — applications to capacity expansion for special services. The Bell System Technical Journal, 60(1):57–87, January 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-1-57.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-1-57.pdf.

Nussbaum:1982:VSS

[Nus82] E. Nussbaum. 1A voice storage system: Voice storage in the network — perspective and history. The Bell System Technical Journal, 61(5):811–813, May/ June 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-5-811.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-5-811.pdf.

Nelson:1980:DBD

[NWB^{+}80] T. J. Nelson, R. Wolfe, S. L. Blank, P. I. Bonyhard, W. A. Johnson, B. J. Roman, and G. P. Vella Coleiro. Design of bubble device elements employing ion-implanted propagation patterns. The Bell System Technical Journal, 59(2):229–257, February 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-2-229.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-2-229.pdf.

Ogawa:1981:NCG

[Oga81] K. Ogawa. Noise caused by GaAs MESFETs in optical receivers. The Bell System Technical Journal, 60(6):923–928, July/ August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-6-923.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-6-923.pdf.

Ogawa:1982:CSM

[Oga82] K. Ogawa. Considerations for single-mode fiber systems. The Bell System Technical Journal, 61(8):1919–1931, October 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-8-1919.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-8-1919.pdf.

ONeill:1980:DWP

[O’N80]  L. A. O’Neill.  Designer’s

Workbench: Philosophy. The Bell System Technical Journal, 59(9):1757–1765, November 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-9-1757.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-9-1757.pdf.

Ogawa:1983:LWO

[OOB83] K. Ogawa, B. Owen, and H. J.

Boll. A long-wavelength optical receiver using a shortchannel Si-MOSFET. The Bell System Technical Journal, 62(5):1181–1188, May/ June 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-5-1181.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-5-1181.pdf.

Overstreet:1982:ARS

[Ove82]  E. A. Overstreet.  Auto-

mated repair service bureau: Economic evaluation. The Bell System Technical Journal, 61(6):1325–1344, July/ August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-6-1325.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-6-1325.pdf.

Ozarow:1984:WTC

[OW84] Lawrence H. Ozarow and Aaron D. Wyner. Wire-tap channel II. AT&T Bell Laboratories Technical Journal, 63(10 part 1):2135– 2157, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Ozarow:1980:SCP

[Oza80] L. Ozarow. On a source-coding problem with two channels and three receivers. The Bell System Technical Journal, 59(10): 1909–1921, December 1980. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol59/bstj59-10-1909.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol59-1980/articles/
                                        bstj59-10-1909.pdf.

Panish:1989:MBE

[Pan89] Morton B. Panish. Molecularbeam epitaxy. AT&T Technical Journal, 68(1):43–52, January/February 1989. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Paul:1983:HFB

[Pau83] L. M. Paul. Human factors and behavioral science: Human factors comparison of two fiberoptic continuous-groove fieldrepair splicing techniques. The Bell System Technical Journal, 62(6):1713–1721, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1713.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1713.pdf.

Presby:1982:VCD

[PC82] H. M. Presby and R. Chang.

Video colorization diagnostics in optical telecommunications. The Bell System Technical Journal, 61(3):267–282, March 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-3-267.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-3-267.pdf.

Peterson:1980:CVS

[PCP80] G. E. Peterson, A. Carnevale, and U. C. Paek. Comparison of vector and scalar modes in a lightguide with a hyperbolic secant index distribution.

The Bell System Technical Journal, 59(9):1681–1691, November 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-9-1681.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-9-1681.pdf.

Peterson:1980:ENS

[PCPB80] G. E. Peterson, A. Carnevale, U. C. Paek, and D. W. Berreman. An exact numerical solution to Maxwell’s equations for lightguides. The Bell System Technical Journal, 59(7):1175–1196, September 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-7-1175.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-7-1175.pdf.

Peterson:1981:NCO

[PCPF81] G. E. Peterson, A. Carnevale, U. C. Paek, and J. W. Fleming. Numerical calculation of optimum a for a germaniadoped silica lightguide. The Bell System Technical Journal, 60(4):455–470, April 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-4-455.
pdf;   http://www.alcatel-

lucent.com/bstj/vol60-1981/
articles/bstj60-4-455.pdf.

Peterson:1984:ETL

[Pet84] George E. Peterson. Electrical transmission lines as models for soliton propagation in materials: elementary aspects of video solitons. AT&T Bell Laboratories Technical Journal, 63 (6 part 1):901–919, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Pettijohn:1986:AQD

[Pet86] Caryl L. Pettijohn. Achieving quality in the development process. AT&T Technical Journal, 65(2):85–93, March/April 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Pfeufer:1980:HIS

[Pfe80] F. J. Pfeufer. The human interface to the switched access remote test system. The Bell System Technical Journal, 59(4):529–556, April 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-4-529.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-4-529.pdf.

Pfeiffer:1983:TND

[Pfe83]  J. Pfeiffer, Jr.  Total net-

work data system: Operating company perspective. The Bell System Technical Journal, 62(7):2459–2471, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-7-2459.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-7-2459.pdf.

Phadke:1982:QEP

[Pha82]  M. S. Phadke.  Quality

evaluation plan using adaptive Kalman filtering. The Bell System Technical Journal, 61(8):2081–2107, October 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-8-2081.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-8-2081.pdf.

Phadke:1986:DOC

[Pha86] M. S. Phadke. Design optimization case studies. AT&T Technical Journal, 65(2):51–68, March/ April 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Philen:1982:MOD

[Phi82] D. L. Philen. Measurements of OH diffusion in optical-fiber cores. The Bell System Technical Journal, 61(3):283–293, March 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

images/Vol61/bstj61-3-283.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-3-283.pdf.

Pike:1984:BMG

[Pik84] Rob Pike. The Blit: a multiplexed graphics terminal. AT&T Bell Laboratories Technical Journal, 63(8 part 2):1607–1631, October 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Pirsch:1982:AII

[Pir82]  P. Pirsch.  Adaptive intra-

interframe DPCM coder. The Bell System Technical Journal, 61(5):747–764, May/June 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-5-747.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-5-747.pdf.

Pike:1984:PDU

[PK84] Rob Pike and Brian W. Kernighan.

Program design in the UNIX system environment. AT&T Bell Laboratories Technical Journal, 63(8 part 2):1595–1605, October 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Pinnel:1987:ISR

[PK87] M. R. Pinnel and W. H. Knausenberger. Interconnection system requirements and modeling. AT&T Technical Journal, 66(4):

45–56, 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Phadke:1983:LQC

[PKSG83] M. S. Phadke, R. N. Kackar, D. V. Speeney, and M. J. Grieco. Off-line quality control in integrated circuit fabrication using experimental design. The Bell System Technical Journal, 62(5):1273–1309, May/ June 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-5-1273.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-5-1273.pdf.

Pearson:1982:BJB

[PLR82] A. D. Pearson, P. D. Lazay, and W. A. Reed. B.S.T.J. briefs: Fabrication and properties of single-mode optical fiber exhibiting low dispersion, low loss, and tight mode confinement simultaneously. The Bell System Technical Journal, 61(2):262–266, February 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-2-262.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-2-262.pdf.

Paek:1981:EFF

[PPC81a] U. C. Paek, G. E. Peterson, and A. Carnevale. Electromagnetic fields, field confinement, and energy flow in dispersionless single-mode lightguides with graded-index profiles. The Bell System Technical Journal, 60(8):1727–1743, October 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-8-1727.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-8-1727.pdf.

Paeke:1981:DSM

[PPC81b] U. C. Paeke, G. E. Peterson, and A. Carnevale. Dispersionless single-mode lightguides with a index profile. The Bell System Technical Journal, 60(5):583–598, May/June 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-5-583.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-5-583.pdf.

Perdue:1986:CVS

[PR86] R. J. Perdue and E. L. Rissanen. Conversant 1 voice system: architecture and applications. AT&T Technical Journal, 65(5): 34–47, September/October 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Prabhu:1980:DEA

[Pra80]  V. K. Prabhu.  The de-

tection efficiency of 16-ary QAM. The Bell System Technical Journal, 59(4):639–656, April 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-4-639.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-4-639.pdf.

Prather:1983:TPT

[Pra83] R. E. Prather. Theory of program testing — an overview. The Bell System Technical Journal, 62(10):3073–3105, December 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/
                                        Vol62/bstj62-10-3073.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol62-1983/articles/
                                        bstj62-10-3073.pdf.

Presby:1981:PCO

[Pre81] H. M. Presby. Profile characterization of optical fibers — A comparative study. The Bell System Technical Journal, 60(7):1335–1362, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-7-1335.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-7-1335.pdf.

Pferd:1981:OFS

[PS81a] W. Pferd and G. C. Stocker.

Optical fibers for scanning digitizers. The Bell System Technical Journal, 60(4):523–534, April 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-4-523.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-4-523.pdf.

Prabhu:1981:PPS

[PS81b] V. K. Prabhu and J. Salz. On the performance of phase-shift-keying systems. The Bell System Technical Journal, 60(10): 2307–2343, December 1981. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol60/bstj60-10-2307.pdf;
http://www.alcatel-lucent.
com/bstj/vol60-1981/articles/
bstj60-10-2307.pdf.

Prabhu:1982:FHS

[PS82] V. K. Prabhu and R. Steele.

Frequency-hopped single-sideband modulation for mobile radio. The Bell System Technical Journal, 61(7):1389–1411, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-7-1389.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-7-1389.pdf.

Pulat:1988:AMS

[Pul88] B. Mustafa Pulat. Automating manual storeroom zones in a factory. AT&T Technical Journal, 67(2):87–94, March/April 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Pack:1982:KFM

[PW82] C. D. Pack and B. A. Whitaker. Kalman filter models for network forecasting. The Bell System Technical Journal, 61(1):1– 14, January 1982. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

                                        http://bstj.bell-labs.com/
                                        BSTJ/images/Vol61/bstj61-1-
                                        1.pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-1-1.pdf.

Pierson:1988:SMC

[PW88] Lyndon G. Pierson and Edward L. Witzke. A security methodology for computer networks. AT&T Technical Journal, 67(3):28–36, May 1988. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Quinn:1983:PAD

[QEG83] J. L. Quinn, R. L. Engram, and F. M. Goetz. The 3B20D processor & DMERT operating system: Diagnostic tests and control software. The Bell System Technical Journal, 62(1):367–381, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-1-367.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-1-367.pdf.

Rabiner:1982:NSF

[Rab82] L. R. Rabiner. Note on some factors affecting performance of dynamic time warping algorithms for isolated word recognition. The Bell System Technical Journal, 61(3):363–373, March 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-3-363.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-3-363.pdf.

Rabiner:1984:AEC

[Rab84] Lawrence R. Rabiner. On the application of energy contours to the recognition of connected word sequences. AT&T Bell Laboratories Technical Journal, 63(9):1981–1995, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Rainal:1981:CCC

[Rai81] A. J. Rainal. Current-carrying capacity of fine-line printed conductors. The Bell System Technical Journal, 60(7): 1375–1388, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/

                                        images/Vol60/bstj60-7-1375.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-7-1375.pdf.

Rainal:1984:CIN

[Rai84] A. J. Rainal. Computing inductive noise of chip packages. AT&T Bell Laboratories Technical Journal, 63(1):177–195, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Ramamurthy:1988:AMU

[Ram88] Gopalakrishnan Ramamurthy. An analytical model for Unix systems. AT&T Technical Journal, 67(5):86–99, September/October 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Ransom:1984:LAD

[Ran84] M. Niel Ransom. Local area data transport service overview. AT&T Bell Laboratories Technical Journal, 63(6 part 2):1113– 1134, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Ray:1988:AVI

[Ray88] R. Ray. Automated visual inspection of solder bumps. AT&T Technical Journal, 67(2):47–60, March/April 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Rowley:1982:UMB

[RB82] J. E. Rowley and J. Bernardini. Using magnetic bubble memories to provide recorded announcements. The Bell System Technical Journal, 61 (8):1841–1862, October 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-8-1841.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-8-1841.pdf.

Rolund:1983:PAD

[RBH83] M. W. Rolund, J. T. Beckett, and D. A. Harms. The 3B20D processor & DMERT operating system: 3B20D central processing unit. The Bell System Technical Journal, 62(1):191–206, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-1-191.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-1-191.pdf.

Rabiner:1982:ITP

[RBW82] L. R. Rabiner, A. Bergh, and J. G. Wilpon. An improved training procedure for connected-digit recognition. The Bell System Technical Journal, 61(6):981–1001, July/August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-6-981.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-6-981.pdf.

Reed:1986:TOC

[RCS86] William A. Reed, Leonard G. Cohen, and Hen-Tai Shang. Tailoring optical characteristics of dispersion-shifted lightguides for applications near 1.55 μm. AT&T Technical Journal, 65(5): 105–122, September 1986. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Rhodes:1982:ARS

[RD82] S. P. Rhodes and L. S. Dickert. Automated repair service bureau: The trouble report evaluation and analysis tool. The Bell System Technical Journal, 61(6):1153–1163, July/ August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-6-1153.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-6-1153.pdf.

Reiser:1981:CTA

[Rei81] J. F. Reiser. Compiling threeaddress code for C programs. The Bell System Technical Journal, 60(2):159–166, February 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-2-159.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol60-1981/
articles/bstj60-2-159.pdf.

Reich:1985:MAI

[Rei85] Leonard S. Reich. The making of American industrial research: science and business at GE and Bell, 1876–1926. Studies in economic history and policy: the United States in the twentieth century / edited by Louis Galambos and Robert Gallman. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-30529-2. xvi + 309 pp. LCCN HD30.42.U5

R44 1985. URL http://www.
loc.gov/catdir/description/
cam023/85006645.html;   http:
//www.loc.gov/catdir/samples/
cam031/85006645.html;   http:
//www.loc.gov/catdir/toc/
cam031/85006645.html.

Reudink:1983:EPL

[Reu83] D. O. Reudink. Estimates of path loss and radiated power for UHF mobile-satellite systems. The Bell System Technical Journal, 62(8):2493–2512, October 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-8-2493.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-8-2493.pdf.

Ritchie:1984:EUT

[Rit84a] Dennis M. Ritchie. Evolution of the UNIX time-sharing system. AT&T Bell Laboratories Technical Journal, 63(8 part 2): 1577–1593, October 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Ritchie:1984:SIO

[Rit84b] Dennis M. Ritchie. Stream input-output system. AT&T Bell Laboratories Technical Journal, 63(8 part 2):1897–1910, October 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Ramamoorthy:1984:EAS

[RJ84] Venkatasubbarao Ramamoorthy and Nuggehally S. Jayant. Enhancement of ADPCM speech by adaptive postfiltering. AT&T Bell Laboratories Technical Journal, 63(8 part 1):1465–1475, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Rabiner:1985:SPC

[RJLS85] Lawrence R. Rabiner, Biing-Hwang Juang, Stephen E. Levinson, and Man Mohan Sondhi. Some properties of continuous hidden Markov model representations. AT&T Technical Journal, 64(6 part 1):1251–1270, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Rabiner:1983:AVQ

[RLS83] L. R. Rabiner, S. E. Levinson, and M. M. Sondhi. On the application of vector quantization and hidden Markov models to speaker-independent, isolated word recognition. The Bell System Technical Journal, 62(4):1075–1105, April 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-4-1075.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-4-1075.pdf.

Rabiner:1984:UHM

[RLS84] L. R. Rabiner, S. E. Levinson, and M. M. Sondhi. On the use of hidden Markov models for speaker-independent recognition of isolated words from a medium-size vocabulary. AT&T Bell Laboratories Technical Journal, 63(4):627–642, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Ramakrishnan:1982:OPS

[RM82] K. G. Ramakrishnan and D. Mitra. An overview of PANACEA, a software package for analyzing Markovian queueing networks. The Bell System Technical Journal, 61(10):2849–2872, December 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol61/bstj61-10-2849.pdf;
http://www.alcatel-lucent.
com/bstj/vol61-1982/articles/
bstj61-10-2849.pdf.

Robbins:1982:SSM

[RN82] J. D. Robbins and A. N.

Netravali. Spatial subsampling in motion-compensated television coders. The Bell System Technical Journal, 61 (8):1895–1910, October 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-8-1895.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-8-1895.pdf.

Roberts:1982:ITN

[Rob82] C. S. Roberts. Implementing and testing new versions of a good, 48-bit, pseudorandom number generator. The Bell System Technical Journal, 61(8):2053–2063, October 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-8-2053.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-8-2053.pdf.

Rochkind:1982:DSS

[Roc82] M. J. Rochkind. Database systems: Structure of a database file system for the UNIX operating system. The Bell System Technical Journal, 61 (9):2387–2405, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2387.
pdf;   http://www.alcatel-

lucent.com/bstj/vol61-1982/
articles/bstj61-9-2387.pdf.

Rocca:1986:IA

[Roc86] Richard T. Rocca. ISDN architecture. AT&T Technical Journal, 65(1):4–18, January/February 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Rogalski:1987:EGL

[Rog87] James E. Rogalski. Evolution of gigabit lightwave transmission systems. AT&T Technical Journal, 66(3):32–40, May/ June 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Roome:1982:DSI

[Roo82] W. D. Roome. Database systems: The intelligent store: A content-addressable page manager. The Bell System Technical Journal, 61(9): 2567–2596, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2567.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2567.pdf.

Rosenbaum:1983:HFB

[Ros83] D. A. Rosenbaum. Human factors and behavioral science: Central control of movement timing. The Bell System Technical Journal, 62(6):1647–1657, July/ August 1983. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-6-1647.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-6-1647.pdf.

Rosenberg:1984:PMP

[Ros84a] A. E. Rosenberg. A probabilistic model for the performance of word recognizers. AT&T Bell Laboratories Technical Journal, 63(1):1–32, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Rosler:1984:ECP

[Ros84b] Lawrence Rosler. Evolution of C — past and future. AT&T Bell Laboratories Technical Journal, 63(8 part 2):1685–1699, October 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Rowe:1982:MNG

[Row82] H. E. Rowe. Memoryless nonlinearities with Gaussian inputs: Elementary results. The Bell System Technical Journal, 61(7):1519–1525, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-7-1519.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-7-1519.pdf.

Rowe:1984:PCB

[Row84a] H. E. Rowe. Processing channelbank spectrometer data. AT&T Bell Laboratories Technical Journal, 63(4):565–585, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Rowe:1984:SFR

[Row84b] Harrison E. Rowe. Spatial filtering radio astronomical data: one-dimensional case. AT&T Bell Laboratories Technical Journal, 63(9):1997–2031, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Rabiner:1984:PIW

[RPS84] Lawrence R. Rabiner, Kok-Chin Pan, and Frank K. Soong. On the performance of isolated word speech recognizers using vector quantization and temporal energy contours. AT&T Bell Laboratories Technical Journal, 63(7):1245–1260, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Rosenberg:1980:RSS

[RRW80] A. E. Rosenberg, L. R. Rabiner, and J. G. Wilpon. Recognition of spoken spelled names for directory assistance using speaker-independent templates. The Bell System Technical Journal, 59(4):571–592, April 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-4-571.

                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-4-571.pdf.

Rabiner:1982:IWR

[RRWK82] L. R. Rabiner, A. E. Rosenberg, J. G. Wilpon, and W. J. Keilin. Isolated word recognition for large vocabularies. The Bell System Technical Journal, 61 (10):2989–3005, December 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol61/bstj61-10-2989.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol61-1982/articles/
                                        bstj61-10-2989.pdf.

Rabiner:1985:SFV

[RS85a] L. R. Rabiner and F. K. Soong. Single-frame vowel recognition using vector quantization with several distance measures. AT&T Technical Journal, 64(10):2319–2330, 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Rege:1985:PBA

[RS85b] K. M. Rege and B. Sengupta. A priority-based admission scheme for a multiclass queueing system. AT&T Technical Journal, 64(7): 1731–1753, September 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Rege:1985:STD

[RS85c] K. M. Rege and B. Sengupta.

Sojourn time distribution in a multiprogrammed computer system. AT&T Technical Journal, 64(5):1077–1090, 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Robertson:1986:EMS

[RS86] Lenard B. Robertson and Glenn A. Secor. Effective management of software development. AT&T Technical Journal, 65(2):94–101, March/April 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Roy:1988:SHQ

[RS88] Pradip K. Roy and Ashok K.

Sinha. Synthesis of high-quality ultra-thin gate oxides for ULSI applications. AT&T Technical Journal, 67(6):155–174, November 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Rabiner:1983:NPV

[RSL83] L. R. Rabiner, M. M. Sondhi, and S. E. Levinson. Note on the properties of a vector quantizer for LPC coefficients. The Bell System Technical Journal, 62(8):2603–2616, October 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-8-2603.

pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-8-2603.pdf.

Rabiner:1984:VQC

[RSL84] L. R. Rabiner, M. M. Sondhi, and S. E. Levinson. A vector quantizer combining energy and LPC parameters and its application to isolated word recognition. AT&T Bell Laboratories Technical Journal, 63(5):721–735, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Rubin:1982:ARS

[Rub82]  H. Rubin.  Automated re-

pair service bureau: Secondgeneration mechanized loop testing system — A distributed microprocessor application. The Bell System Technical Journal, 61(6):1257–1274, July/ August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-6-1257.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-6-1257.pdf.

Rummler:1980:TFD

[Rum80] W. D. Rummler. Time- and frequency-domain representation of multipath fading on line-of-sight microwave paths. The Bell System Technical Journal, 59(5):763–796, May/June 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-5-763.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-5-763.pdf.

Rummler:1982:SMM

[Rum82] W. D. Rummler. A statistical model of multipath fading on a Space-diversity radio channel. The Bell System Technical Journal, 61(9):2185–2219, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2185.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2185.pdf.

Rabiner:1981:TPP

[RW81] L. R. Rabiner and J. G.

Wilpon. A two-pass patternrecognition approach to isolated word recognition. The Bell System Technical Journal, 60(5):739–766, May/June 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-5-739.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-5-739.pdf.

Rowland:1983:PAD

[RW83] B. R. Rowland and R. J.

Welsch. The 3B20D processor & DMERT operating system: Software development system. The Bell System Technical Journal, 62(1):275–289, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-1-275.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-1-275.pdf.

Reeds:1984:FSU

[RW84] James A. Reeds and Peter J. Weinberger. File security and the UNIX system crypt command. AT&T Bell Laboratories Technical Journal, 63(8 part 2): 1673–1683, October 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic). Reprinted in [AT&T86, pp. 93– 103].

Rabiner:1981:EVA

[RWA81] L. R. Rabiner, J. G. Wilpon, and J. G. Ackenhusen. On the effects of varying analysis parameters on an LPC-based isolated word recognizer. The Bell System Technical Journal, 60(6):893–911, July/ August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-6-893.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-6-893.pdf.

Rabiner:1986:SMT

[RWJ86] L. R. Rabiner, J. G. Wilpon, and Biing-Hwang Juang. A segmental k-means training procedure for connected word recognition.

AT&T Technical Journal, 65(3): 21–31, 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Rabiner:1980:VCR

[RWR80] L. R. Rabiner, J. G. Wilpon, and A. E. Rosenberg. A voice-controlled, repertory-dialer system. The Bell System Technical Journal, 59(7):1153– 1163, September 1980. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol59/bstj59-7-1153.pdf;
http://www.alcatel-lucent.
com/bstj/vol59-1980/articles/
bstj59-7-1153.pdf.

Rustako:1985:LSF

[RWRH85] A. J. Rustako, Jr., C. B. Woodworth, R. S. Roman, and H. H. Hoffman. A laboratory simulation facility for multipath fading microwave radio channels. AT&T Technical Journal, 64 (10):2281–2317, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Rabiner:1984:AET

[RWT84] L. R. Rabiner, J. G. Wilpon, and S. G. Terrace. On the application of embedded training to connected letter recognition for directory listing retrieval. AT&T Bell Laboratories Technical Journal, 63(3):459–477, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Rzeszewski:1983:CHD

[Rze83] T. S. Rzeszewski. A compatible high-definition television system. The Bell System Technical Journal, 62(7):2091–2111, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-7-2091.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-7-2091.pdf.

Sharma:1981:FCF

[SA81a] D. K. Sharma and S. R. Ahuja. A first-come-first-serve bus-allocation scheme using ticket assignments. The Bell System Technical Journal, 60(7):1257–1269, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-7-1257.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-7-1257.pdf.

Smith:1981:VBA

[SA81b] J. O. Smith and J. B. Allen. Variable bandwidth adaptive delta modulation. The Bell System Technical Journal, 60 (5):719–737, May/June 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-5-719.
pdf;   http://www.alcatel-

lucent.com/bstj/vol60-1981/
articles/bstj60-5-719.pdf.

Sandberg:1985:APR

[SA85] I. W. Sandberg and J. B. Allen.

Almost-periodic response determination for models of the basilar membrane. AT&T Technical Journal, 64(8):1775–1786, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Saleh:1980:CFR

[Sal80]  A. A. M. Saleh.  Compu-

tation of the frequency response of a class of symmetric N-way power dividers. The Bell System Technical Journal, 59(8):1493–1512, October 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-8-1493.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-8-1493.pdf.

Saleh:1982:MAM

[Sal82]  A. A. M. Saleh.  Matrix

analysis of mildly nonlinear, multiple-input, multiple-output systems with memory. The Bell System Technical Journal, 61(9):2221–2243, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2221.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-9-2221.pdf.

Salz:1983:SPD

[Sal83]  J. Salz.  On the start-up

problem in digital echo cancelers. The Bell System Technical Journal, 62(6):1353–1364, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-6-1353.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-6-1353.pdf.

Salz:1985:CLC

[Sal85a]  Jack Salz.  Coherent light-

wave communications. AT&T Technical Journal, 64(10):2153– 2209, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Salz:1985:DTC

[Sal85b] Jack Salz. Digital transmission over cross-coupled linear channels. AT&T Technical Journal, 64(6 part 1):1147–1159, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Sandberg:1980:DND

[San80] I. W. Sandberg. Diffeomorphisms and Newton-direction algorithms. The Bell System Technical Journal, 59(9): 1721–1733, November 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-9-1721.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-9-1721.pdf.

Sandberg:1981:CRN

[San81a] I. W. Sandberg. Criteria for the response of nonlinear systems to be L-asymptotically periodic. The Bell System Technical Journal, 60(10):2359–2371, December 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol60/bstj60-10-2359.pdf;
http://www.alcatel-lucent.
com/bstj/vol60-1981/articles/
bstj60-10-2359.pdf.

Sandberg:1981:NDA

[San81b] I. W. Sandberg. On Newtondirection algorithms and diffeomorphisms. The Bell System Technical Journal, 60(3): 339–346, March 1981. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol60/bstj60-3-339.pdf;
http://www.alcatel-lucent.
com/bstj/vol60-1981/articles/
bstj60-3-339.pdf.

Sandberg:1982:ENS

[San82a] I. W. Sandberg. Expansions for nonlinear systems. The Bell System Technical Journal, 61(2):159–199, February 1982. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-2-159.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-2-159.pdf.

Sandberg:1982:VET

[San82b] I. W. Sandberg. Volterra expansions for time-varying nonlinear systems. The Bell System Technical Journal, 61 (2):201–225, February 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-2-201.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-2-201.pdf.

Sandberg:1985:CGE

[San85a] I. W. Sandberg. Criteria for the global existence of functional expansions for input/output maps. AT&T Technical Journal, 64(7): 1639–1658, September 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Sandberg:1985:NIO

[San85b] Irwin W. Sandberg. Nonlinear input-output maps and approximate representations. AT&T Technical Journal, 64(8):1967– 1983, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Sanferrare:1987:TLS

[San87] Robert J. Sanferrare. Terrestrial lightwave systems. AT&T Technical Journal, 66(1):95–107, January 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Streeter:1983:HFB

[SAT83] L. A. Streeter, J. M. Ackroff, and G. A. Taylor. Human factors and behavioral science: On abbreviating command names. The Bell System Technical Journal, 62(6):1807–1826, July/ August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1807.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1807.pdf.

Stone:1980:BJB

[SB80] J. Stone and C. A. Burrus. B.S.T.J. briefs: Reduction of the 1.38-μ m water peak in optical fibers by deuterium-hydrogen exchange. The Bell System Technical Journal, 59(8):1541–1548, October 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-8-1541.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-8-1541.pdf.

Schenker:1982:ARS

[SB82] L. Schenker and S. J. Barbera. Automated repair service bureau: Preface. The Bell System Technical Journal, 61(6):1095–1096, July/ August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-6-1095.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-6-1095.pdf.

Steele:1983:SRS

[SB83] R. Steele and F. Benjamin. Sample reduction and subsequent adaptive interpolation of speech signals. The Bell System Technical Journal, 62(6): 1365–1398, July/August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1365.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1365.pdf.

Sulewski:1982:DBL

[SBD82] P. Sulewski, D. J. Bishop, and R. C. Dynes. A description of the Bell Laboratories scanned acoustic microscope. The Bell System Technical Journal, 61(9):2167–2183, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2167.

pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2167.pdf.

Staehler:1983:TSP

[SC83] R. E. Staehler and J. I.

Cochrane. Traffic service position system no. 1B: Overview and objectives. The Bell System Technical Journal, 62(3): 755–764, March 1983. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol62/bstj62-3-755.pdf;
http://www.alcatel-lucent.
com/bstj/vol62-1983/articles/
bstj62-3-755.pdf.

Scanlon:1983:PAD

[Sca83] J. M. Scanlon. The 3B20D processor & DMERT operating system: Prologue. The Bell System Technical Journal, 62(1):167–169, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-1-167.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-1-167.pdf.

Schiavone:1981:PPR

[Sch81] J. A. Schiavone. Prediction of positive refractivity gradients for line-of-sight microwave radio paths. The Bell System Technical Journal, 60(6):803–822, July/ August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-6-803.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-6-803.pdf.

Schenker:1983:TND

[Sch83a] L. Schenker. Total network data system: Introduction. The Bell System Technical Journal, 62(7):2123–2125, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-7-2123.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-7-2123.pdf.

Schmidt:1983:CLD

[Sch83b] R. L. Schmidt. A comparison of line difference predictions for time-frequency multiplexing of television signals. The Bell System Technical Journal, 62(7):1955–1975, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-7-1955.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-7-1955.pdf.

Saul:1985:RIP

[SCS85] R. H. Saul, F. S. Chen, and P. W. Shumate, Jr. Reliability of InGaAs photodiodes for SL applications. AT&T Technical Journal, 64(3):861–882, 1985.

CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Stone:1984:OAR

[SCWB84] Julian Stone, Andrew R. Chraplyvy, Jay M. Wiesenfeld, and Charles A. Burrus Jr. Overtone absorption and Raman spectra of H_{2} and D_{2} in silica optical fibers. AT&T Bell Laboratories Technical Journal, 63(6 part 1):991–1000, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Semplak:1980:HRA

[Sem80] R. A. Semplak. Horn-reflector antenna — eliminating weathercover reflections. The Bell System Technical Journal, 59 (8):1333–1342, October 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-8-1333.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-8-1333.pdf.

Semmelman:1981:DSP

[Sem81]  C. L. Semmelman.  Dig-

ital signal processor: Design of the assembler. The Bell System Technical Journal, 60(7):1483–1497, September 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-7-1483.
pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-7-1483.pdf.

Semplak:1983:HMS

[Sem83a] R. A. Semplak. Hybrid-mode, shielded, offset parabolic antenna. The Bell System Technical Journal, 62(1):111–120, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-1-111.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-1-111.pdf.

Semplak:1983:MSN

[Sem83b] R. A. Semplak. Measurements of selective near-in sidelobe reduction of a pyramidal, horn-reflector antenna. The Bell System Technical Journal, 62(3):595–605, March 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-3-595.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-3-595.pdf.

Semplak:1985:GMT

[Sem85] R. A. Semplak. 100-GHz measurements of two astigmatic launchers. AT&T Technical Journal, 64(5):1019–1031, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Sethi:1980:DCR

[SF80] D. P. S. Sethi and H. T. Freedman. Decision criteria for rehabilitation of the distribution network. The Bell System Technical Journal, 59(6):881–895, July/ August 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-6-881.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-6-881.pdf.

Smithgall:1981:HSM

[SF81] D. H. Smithgall and R. E.

Frazee. High-speed measurement and control of fibercoating concentricity. The Bell System Technical Journal, 60(9):2065–2080, November 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-9-2065.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-9-2065.pdf.

Smith:1982:SAP

[SF82] R. G. Smith and S. R. Forrest. Sensitivity of avalanche photodetector receivers for longwavelength optical communications. The Bell System Technical Journal, 61(10):2929– 2946, December 1982. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/

                                        Vol61/bstj61-10-2929.pdf;
                                        http://www.alcatel-lucent.
                                        com/bstj/vol61-1982/articles/
                                        bstj61-10-2929.pdf.

Sharma:1981:MPE

[Sha81] D. K. Sharma. McDonald’s problem — an example of using Dijkstra’s programming method. The Bell System Technical Journal, 60(9):2157–2165, November 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-9-2157.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-9-2157.pdf.

Shapira:1985:SAR

[Sha85] Joseph Shapira. Simple analytical representation of antenna spatial radiation patterns with application to the pyramidal horn-reflector antenna. AT&T Technical Journal, 64(7):1601– 1624, September 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Sharp:1988:DCO

[Sha88] R. L. Sharp. Design of a certifiable one-way data-flow device. AT&T Technical Journal, 67(3):44–52, May 1988. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Sheng:1988:VPL

[She88] Diane D. Sheng. Virtual privateline performance and customer cost impacts. AT&T Technical Journal, 67(6):47–68, November 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Shichor:1982:FRE

[Shi82] E. Shichor. Fast recursive estimation using the lattice structure. The Bell System Technical Journal, 61(1):97–115, January 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-1-97.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-1-97.pdf.

Siller:1984:EIW

[Sil84] C. A. Siller, Jr. Experimental investigation of wide-angle sidelobe suppression in a pyramidal horn-reflector antenna. AT&T Bell Laboratories Technical Journal, 63(4):531–544, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Sengupta:1985:CRT

[SJ85] B. Sengupta and D. L. Jagerman. A conditional response time of the M/M/1 processorsharing queue. AT&T Technical Journal, 64(2 part 1):409– 421, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Stelte:1984:ATI

[SKP84] David J. Stelte, Henry J. Kafka, and W. Joseph Paule. AT&T Technologies implementation of Local Area Data Transport — a hardware and software overview. AT&T Bell Laboratories Technical Journal, 63(6 part 2):1135– 1190, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Smith:1982:PLD

[Smi82] P. W. Smith. On the physical limits of digital optical switching and logic elements. The Bell System Technical Journal, 61(8):1975–1993, October 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-8-1975.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-8-1975.pdf.

Smith:1983:MSS

[Smi83] D. R. Smith. A model for specialservice circuit activity. The Bell System Technical Journal, 62 (10):2911–2934, December 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol62/bstj62-10-2911.pdf;
http://www.alcatel-lucent.
com/bstj/vol62-1983/articles/
bstj62-10-2911.pdf.

Singhal:1989:CDO

[SMNV89] K. Singhal, C. C. McAndrew, S. R. Nassif, and V. Visvanathan. The CENTER design optimization system. AT&T Technical Journal, 68(3):77–92, 1989. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Stuller:1980:ITC

[SNR80] J. A. Stuller, A. N. Netravali, and J. D. Robbins. Interframe television coding using gain and displacement compensation. The Bell System Technical Journal, 59(7):1227–1240, September 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-7-1227.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-7-1227.pdf.

Sondhi:1983:RPS

[Son83]  M. M. Sondhi.  Random

processes with specified spectral density and first-order probability density. The Bell System Technical Journal, 62(3):679–701, March 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-3-679.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-3-679.pdf.

Shah:1989:IVD

[SP89] Nitin J. Shah and Shin-Shem Pei. III–V device technologies for electronic applications. AT&T Technical Journal, 68(1):19–28, January/February 1989. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Soong:1987:VQA

[SRJR87] Frank K. Soong, E. Rosenberg, Biing-Hwang Juang, and Lawrence R. Rabiner. Vector quantization approach to speaker recognition. AT&T Technical Journal, 66(2):14–26, 1987. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Saleh:1983:ALP

[SS83] A. A. M. Saleh and J. Salz. Adaptive linearization of power amplifiers in digital radio systems. The Bell System Technical Journal, 62(4):1019–1033, April 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-4-1019.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-4-1019.pdf.

Sears:1984:PQH

[SS84] F. M. Sears and J. R. Simpson. Polarization quality of highbirefringence single-mode fibers. AT&T Bell Laboratories Technical Journal, 63(2):365–371, 1984.

CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Shepp:1986:CTP

[SS86] Lawrence A. Shepp and Sushanta Srivastava. Computer tomography of PKM and AKM exit cones. AT&T Technical Journal, 65(1):78–88, January/February 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Sondhi:1981:IQN

[SSR81] M. M. Sondhi, C. E. Schmidt, and L. R. Rabiner. Improving the quality of a noisy speech signal. The Bell System Technical Journal, 60(8):1847–1859, October 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-8-1847.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-8-1847.pdf.

Shepp:1980:PMA

[SSW80] L. A. Shepp, D. Slepian, and A. D. Wyner. On prediction of moving-average processes. The Bell System Technical Journal, 59(3):367–415, March 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-3-367.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-3-367.pdf.

Steel:1984:TEC

[SSW84] Raymond Steel, Carl-Erik Sundberg, and Wei Choong Wong. Transmission errors in companded PCM over Gaussian and Rayleigh fading channels. AT&T Bell Laboratories Technical Journal, 63(6 part 1):955–990, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Sampson:1982:SPC

[ST82] S. F. Sampson and D. W. Tietz. Stored program controlled network: Data base administration system — architecture and data base design. The Bell System Technical Journal, 61(7):1779–1797, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-7-1779.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-7-1779.pdf.

Steele:1982:PSN

[Ste82]  R. Steele.  Peak signal-to-

noise formulas for multistage delta modulation with RCshaped Gaussian input signals. The Bell System Technical Journal, 61(3):347–362, March 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-3-347.
                                        pdf;   http://www.alcatel-

lucent.com/bstj/vol61-1982/
articles/bstj61-3-347.pdf.

Strakhov:1983:AMW

[Str83]  N. A. Strakhov.  Applica-

tion of the minimum-weight spanning-tree algorithm to assignment of communication facilities. The Bell System Technical Journal, 62(8):2581–2602, October 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-8-2581.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-8-2581.pdf.

Stroustrup:1984:DAC

[Str84] Bjarne Stroustrup. Data abstraction in C. AT&T Bell Laboratories Technical Journal, 63(8 part 2):1701–1732, October 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Strip:1988:TRM

[Str88] D. R. Strip. Technology for robotic mechanical assembly: force-directed insertions. AT&T Technical Journal, 67(2):23–34, March/April 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Sumner:1983:HFB

[Sum83] E. E. Sumner. Human factors and behavioral science: Introduction. The Bell System Technical Journal, 62(6): 1561–1568, July/August 1983.

CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1561.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-6-1561.pdf.

Sundberg:1982:EPP

[Sun82]  C. E. Sundberg.  Error

probability of partial-response continuous-phase modulation with coherent MSK-type receiver, diversity, and slow Rayleigh fading in Gaussian noise. The Bell System Technical Journal, 61(8):1933–1963, October 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-8-1933.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-8-1933.pdf.

Sundberg:1983:ACC

[Sun83a] C. E. Sundberg. Alternative cell configurations for digital mobile radio systems. The Bell System Technical Journal, 62(7):2037–2065, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-7-2037.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-7-2037.pdf.

Sundberg:1983:CPM

[Sun83b] C. E. Sundberg. On continuous phase modulation in cellular digital mobile radio systems. The Bell System Technical Journal, 62(7):2067–2089, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-7-2067.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-7-2067.pdf.

Surette:1986:QAN

[Sur86] Gerald J. Surette. Quality in AT&T network systems. AT&T Technical Journal, 65(2):21–29, March/April 1986. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Steele:1981:STS

[SV81] R. Steele and D. Vitello.

Simultaneous transmission of speech and data using codebreaking techniques. The Bell System Technical Journal, 60(9):2081–2105, November 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-9-2081.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-9-2081.pdf.

Salz:1980:SSS

[SW80] J. Salz and J. J. Werner. Spectral shaping by simultaneous amplitude and frequency modulation. The Bell System Technical Journal, 59(4):557–570, April 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-4-557.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-4-557.pdf.

Smith:1981:RSE

[SW81] D. R. Smith and W. Whitt. Resource sharing for efficiency in traffic systems. The Bell System Technical Journal, 60(1):39–55, January 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-1-39.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-1-39.pdf.

Sheinbein:1982:SPC

[SW82] D. Sheinbein and R. P. Weber. Stored program controlled network: 800 service using SPC network capability. The Bell System Technical Journal, 61(7):1737–1744, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-7-1737.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-7-1737.pdf.

Swartz:1983:SBS

[SW83] R. G. Swartz and B. A. Wooley. Stabilized biasing of semiconductor lasers. The Bell System Technical Journal, 62(7):1923–1936, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-7-1923.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-7-1923.pdf.

Swartwout:1985:DLL

[Swa85] D. Swartwout. Datastream — a language for large files. AT&T Technical Journal, 64(9):2037– 2060, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Steele:1984:SDD

[SWS84a] Raymond Steele, Wai Choong Wong, and Carl-Erik W. Sundberg. Soft decision demodulation to reduce the effect of transmission errors in logarithmic PCM transmitted over Rayleigh fading channels. AT&T Bell Laboratories Technical Journal, 63 (10 part 1):2193–2213, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Sundberg:1984:WSC

[SWS84b] C.-E. Sundberg, W. C. Wong, and R. Steele. Weighting strategies for companded PCM transmitted over Rayleigh fading and Gaussian channels. AT&T Bell Laboratories Technical Journal, 63(4):587–626, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Schwartz:1982:DFM

[SY82] S. C. Schwartz and Y. S. Yeh. On the distribution function and moments of power sums with log-normal components. The Bell System Technical Journal, 61(7):1441–1462, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-7-1441.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-7-1441.pdf.

Szelag:1980:TDS

[Sze80]  C. R. Szelag.  Trunk de-

mand servicing in the presence of measurement uncertainty. The Bell System Technical Journal, 59(6):845–860, July/ August 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-6-845.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-6-845.pdf.

Szelag:1982:STF

[Sze82]  C. R. Szelag.  A short-

term forecasting algorithm for trunk demand servicing. The Bell System Technical Journal, 61(1):67–96, January 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-1-67.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-1-67.pdf.

Titus:1982:BJB

[TA82] J. M. Titus and H. W. Arnold.

B.S.T.J. briefs: Low-elevation-angle propagation effects on COMSTAR satellite signals. The Bell System Technical Journal, 61(7):1567–1572, September 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-7-1567.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-7-1567.pdf.

Taylor:1984:VSC

[Tay84] Geoffrey W. Taylor. Velocitysaturated characteristics of short-channel MOSFETs. AT&T Bell Laboratories Technical Journal, 63(7):1325–1404, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Toy:1983:PAD

[TG83] W. N. Toy and L. E. Gallaher.

The 3B20D processor & DMERT operating system: Overview and architecture of the 3B20D processor. The Bell System Technical Journal, 62(1):181–190, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-1-181.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-1-181.pdf.

Thompson:1980:DWP

[Tho80] T. J. Thompson. Designer’s Workbench: Providing a production environment. The Bell System Technical Journal, 59(9):1811–1825, November 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-9-1811.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-9-1811.pdf.

Thomas:1987:EBL

[Tho87] D. G. Thomas. Extending the boundaries of lightwave technology. AT&T Technical Journal, 66(1):2–4, January 1987. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Timor:1980:IDS

[Tim80] U. Timor. Improved decoding scheme for frequency-hopped multilevel FSK system. The Bell System Technical Journal, 59 (10):1839–1855, December 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

                                        tronic). URL http://bstj.
                                        bell-labs.com/BSTJ/images/
                                        Vol59/bstj59-10-1839.pdf;
                                        http://www.alcatel-lucent.

com/bstj/vol59-1980/articles/
bstj59-10-1839.pdf.

Timor:1981:MDF

[Tim81] Uzi Timor. Multistage decoding of frequency-hopped FSK system. The Bell System Technical Journal, 60(4):471–483, April 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-4-471.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-4-471.pdf.

Timor:1982:MFH

[Tim82] Uzi Timor. Multitone frequencyhopped MFSK system for mobile radio. The Bell System Technical Journal, 61(10):3007– 3017, December 1982. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol61/bstj61-10-3007.pdf;
http://www.alcatel-lucent.
com/bstj/vol61-1982/articles/
bstj61-10-3007.pdf.

Tolendino:1988:STS

[TNG88] L. F. Tolendino, S. D. Nelson, and S. A. Gossage. Sandia’s terminal switching network and fiber optic loop. AT&T Technical Journal, 67(3):37–43, May 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Tolleth:1984:SSM

[Tol84] David W. Tolleth. System sparing for minicomputer-based operations systems. AT&T Bell Laboratories Technical Journal, 63(6 part 2):1029–1047, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Tortorella:1981:CCT

[Tor81]  M. Tortorella.  Cutoff calls

and telephone equipment reliability. The Bell System Technical Journal, 60(8):1861–1889, October 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-8-1861.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-8-1861.pdf.

Tow:1983:PDC

[Tow83] J. Tow. Practical design considerations for coupled-single-amplifier-biquad active bandpass filters. The Bell System Technical Journal, 62(6): 1399–1413, July/August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-6-1399.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-6-1399.pdf.

Toy:1987:DVT

[Toy87] Wing N. Toy. Dual versus triplication reliability estimations.

AT&T Technical Journal, 66(6): 15–20, November 1987. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Tribolet:1982:IMI

[TRW82] J. M. Tribolet, L. R. Rabiner, and J. G. Wilpon. An improved model for isolated word recognition. The Bell System Technical Journal, 61(9):2289–2312, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2289.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2289.pdf.

Turin:1985:UBV

[Tur85] William Turin. Union bounds on Viterbi algorithm performance. AT&T Technical Journal, 64 (10):2375–2385, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Temkin:1983:ILO

[TZD^{+}83] H. Temkin, C. L. Zipfel, M. A.

DiGiuseppe, A. K. Chin, V. G. Keramidas, and R. H. Saul. In-GaAsP LEDs for 1.3-μ m optical transmission. The Bell System Technical Journal, 62(1):1– 24, January 1983. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol62/bstj62-1-
1.pdf;   http://www.alcatel-

                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-1-1.pdf.

Uhrhane:1982:ARS

[Uhr82] F. J. Uhrhane. Automated repair service bureau: Mechanized loop testing strategies and techniques. The Bell System Technical Journal, 61(6):1209–1234, July/ August 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-6-1209.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol61-1982/
                                        articles/bstj61-6-1209.pdf.

Ungar:1980:ELP

[Ung80] S. G. Ungar. Effects of lightning punctures on the coreshield voltage of buried cable. The Bell System Technical Journal, 59(3):333–366, March 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-3-333.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-3-333.pdf.

Vardi:1981:AOS

[Var81] Y. Vardi. Absenteeism of operators: A statistical study with managerial applications. The Bell System Technical Journal, 60(1):13–38, January 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/

images/Vol60/bstj60-1-13.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-1-13.pdf.

Vesonder:1988:RBP

[Ves88] Gregg T. Vesonder. Rule-based programming in the Unix system. AT&T Technical Journal, 67(1):69–80, January 1988. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Vigants:1981:MRO

[Vig81]  A. Vigants.  Microwave ra-

dio obstruction fading. The Bell System Technical Journal, 60(6):785–801, July/August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-6-785.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-6-785.pdf.

Wang:1988:PSM

[WA88] Jin-Der Wang and Ender Ayanoglu.

Priority statistical multiplexer design for SNA*/SDLC access to a virtual-circuit packet network. AT&T Technical Journal, 67(6): 69–86, November 1988. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Walker:1983:PRS

[Wal83] E. H. Walker. Penetration of radio signals into buildings in the cellular radio environment.

The Bell System Technical Journal, 62(9):2719–2734, November 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-9-2719.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-9-2719.pdf.

Wang:1983:ERM

[Wan83] Y. Y. Wang. Experimental results of 20-Mb/s FSK digital transmission on 4-GHz (TD) radio. The Bell System Technical Journal, 62(5):1209–1231, May/ June 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-5-1209.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-5-1209.pdf.

Warren:1980:DSP

[War80] R. E. Warren. Developing space planning aids for central office equipment systems. The Bell System Technical Journal, 59 (10):1965–2000, December 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol59/bstj59-10-1965.pdf;
http://www.alcatel-lucent.
com/bstj/vol59-1980/articles/
bstj59-10-1965.pdf.

W:1985:ADP

[WC85] Fuhrmann S. W. and R. B.

Cooper. Application of decomposition principle in M/G/1 vacation model to two continuum cyclic queueing models — especially token-ring LANs. AT&T Technical Journal, 64(5):1091– 1099, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

White:1981:HMC

[WCS81] L. D. White, R. W. Coons, and R. C. Strum. A 200-Hz to 30-MHz computer-operated impedance/admittance bridge (COZY). The Bell System Technical Journal, 60(3):405–443, March 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol60/bstj60-3-405.
pdf;   http://www.alcatel-
lucent.com/bstj/vol60-1981/
articles/bstj60-3-405.pdf.

Weinberger:1982:DSM

[Wei82] P. J. Weinberger. Database systems: Making UNIX operating systems safe for databases. The Bell System Technical Journal, 61(9):2407–2422, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2407.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2407.pdf.

Weinberger:1984:CDI

[Wei84] Peter J. Weinberger. Cheap dynamic instruction counting. AT&T Bell Laboratories Technical Journal, 63(8 part 2): 1815–1826, October 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Welsh:1987:LDL

[Wel87]  Frederic S. Welsh.  Light-

wave data links and interfaces. AT&T Technical Journal, 66(1): 65–72, January 1987. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Werner:1985:ECI

[Wer85] J. J. Werner. Effects of channel impairments on the performance of an in-band data-driven echo canceler. AT&T Technical Journal, 64(1 part 1):91–113, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Weste:1981:MIS

[Wes81] N. H. E. Weste. MULGA — an interactive symbolic layout system for the design of integrated circuits. The Bell System Technical Journal, 60(6):823–857, July/ August 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-6-823.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-6-823.pdf.

Wang:1982:DSD

[WH82] C. C. Wang and C. P. Huang. Database systems: Database administration system — architecture and design issues. The Bell System Technical Journal, 61(9):2439–2458, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-9-2439.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-9-2439.pdf.

Whitt:1983:CBD

[Whi83a] W. Whitt. Comparing batch delays and customer delays. The Bell System Technical Journal, 62(7):2001–2009, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-7-2001.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-7-2001.pdf.

Whitt:1983:PQN

[Whi83b] W. Whitt. Performance of the queueing network analyzer. The Bell System Technical Journal, 62(9):2817–2843, November 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-9-2817.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-9-2817.pdf.

Whitt:1983:QNA

[Whi83c] W. Whitt. The queueing network analyzer. The Bell System Technical Journal, 62(9):2779–2815, November 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-9-2779.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-9-2779.pdf.

Whitt:1984:HTA

[Whi84a] W. Whitt. Heavy-traffic approximations for service systems with blocking. AT&T Bell Laboratories Technical Journal, 63(5):689–708, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Whitt:1984:AQE

[Whi84b] W. Whitt. On approximations for queues, I: Extremal distributions. AT&T Bell Laboratories Technical Journal, 63(1):115– 138, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Whitt:1984:AQI

[Whi84c] W. Whitt. On approximations for queues. III. Mixtures of exponential distributions. AT&T Bell Laboratories Technical Journal, 63(1):163–175, 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Whitt:1984:OCM

[Whi84d] Ward Whitt. Open and closed models for networks of queues.

AT&T Bell Laboratories Technical Journal, 63(9):1911–1979, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Whitt:1985:BWS

[Whi85] Ward Whitt. Blocking when service is required from several facilities simultaneously. AT&T Technical Journal, 64(8):1807– 1856, 1985. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Willett:1982:DRS

[Wil82] R. J. Willett. Design of recovery strategies for a fault-tolerant no. 4 electronic switching system. The Bell System Technical Journal, 61(10):3019–3040, December 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

bell-labs.com/BSTJ/images/
Vol61/bstj61-10-3019.pdf;
http://www.alcatel-lucent.
com/bstj/vol61-1982/articles/
bstj61-10-3019.pdf.

Wilpon:1985:SAA

[Wil85] Jay G. Wilpon. A study on the ability to automatically recognize telephone-quality speech from large customer populations. AT&T Technical Journal, 64(2 part 1):423–451, 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Witsenhausen:1980:BJB

[Wit80] H. S. Witsenhausen. B.S.T.J. briefs: On source networks with minimal breakdown degradation. The Bell System Technical Journal, 59(6):1083–1087, July/ August 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-6-1083.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-6-1083.pdf.

Witsenhausen:1981:MWC

[Wit81] H. S. Witsenhausen. Minimizing the worst-case distortion in channel splitting. The Bell System Technical Journal, 60(8):1979–1983, October 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-8-1979.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-8-1979.pdf.

Witsenhausen:1984:CSS

[Wit84] Hans S. Witsenhausen. On the capacity of sticky storage devices. AT&T Bell Laboratories Technical Journal, 63(7):1311– 1323, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

White:1983:MAL

[WM83] I. A. White and S. C. Mettler. Modal analysis of loss and mode mixing in multimode parabolic index splices. The Bell System Technical Journal, 62(5):1189–1207, May/ June 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-5-1189.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-5-1189.pdf.

Weiss:1988:TAD

[WM88] Alan A. Weiss and Debasis Mitra. Transient analysis of a data network with a processorsharing switch. AT&T Technical Journal, 67(5):4–16, September/October 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Worrall:1982:VSS

[Wor82] D. P. Worrall. 1A voice storage system: New custom calling services. The Bell System Technical Journal, 61(5):821–839, May/ June 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-5-821.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-5-821.pdf.

Windhausen:1987:BAN

[WPG87] Richard A. Windhausen, Nimesh D.

Parikh, and P. K. Govind. Basic access network termination chip for ISDN 4-wire loops. AT&T Technical Journal, 66(6):

35–50, November 1987. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Wilpon:1983:RID

[WR83] J. G. Wilpon and L. R. Rabiner. On the recognition of isolated digits from a large telephone customer population. The Bell System Technical Journal, 62(7):1977–2000, September 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol62/bstj62-7-1977.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol62-1983/
                                        articles/bstj62-7-1977.pdf.

Wilpon:1984:IWD

[WRM84] J. G. Wilpon, L. R. Rabiner, and T. Martin. An improved word-detection algorithm for telephone-quality speech incorporating both syntactic and semantic constraints. AT&T Bell Laboratories Technical Journal, 63(3):479–498, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Wong:1982:TDP

[WSX82] W. C. Wong, R. Steele, and C. S. Xydeas. Transmitting data on the phase of speech signals. The Bell System Technical Journal, 61(10):2947–2970, December 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http://bstj.

                                        bell-labs.com/BSTJ/images/

Vol61/bstj61-10-2947.pdf;
http://www.alcatel-lucent.
com/bstj/vol61-1982/articles/
bstj61-10-2947.pdf.

Warner:1980:TDC

[WW80] D. D. Warner and C. L.

Wilson. Two dimensional concentration dependent diffusion. The Bell System Technical Journal, 59(1):1–41, January 1980. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol59/bstj59-1-
1.pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-1-1.pdf.

Witsenhausen:1981:SCM

[WW81] H. S. Witsenhausen and A. D.

Wyner. Source coding for multiple descriptions II: A binary source. The Bell System Technical Journal, 60(10): 2281–2292, December 1981. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol60/bstj60-10-2281.pdf;
http://www.alcatel-lucent.
com/bstj/vol60-1981/articles/
bstj60-10-2281.pdf.

Wolf:1984:CWO

[WWKZ84] Jack K. Wolf, Aaron D. Wyner, Janos Korner, and Jacob Ziv. Coding for a write-once memory. AT&T Bell Laboratories Technical Journal, 63(6 part 2):1089– 1112, 1984. CODEN ABLJER.

ISSN 0748-612X (print), 2376-7162 (electronic).

Wolf:1980:SCM

[WWZ80] J. K. Wolf, A. D. Wyner, and J. Ziv. Source coding for multiple descriptions. The Bell System Technical Journal, 59(8):1417–1426, October 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-8-1417.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-8-1417.pdf.

Wynn:1981:BMD

[Wyn81]  W. D. Wynn.  A bub-

ble memory differential detector. The Bell System Technical Journal, 60(4):485–500, April 1981. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol60/bstj60-4-485.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol60-1981/
                                        articles/bstj60-4-485.pdf.

Yanofchick:1985:DMS

[Yan85] Raymond J. Yanofchick. T — a data management system. AT&T Technical Journal, 64(9):2119–2129, 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Yavelberg:1982:HPE

[Yav82] I. S. Yavelberg. Human performance engineering considerations for very large computerbased systems: The end user. The Bell System Technical Journal, 61(5):765–797, May/June 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-5-765.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-5-765.pdf.

Yeh:1985:NAS

[YG85] Y. S. Yeh and L. J. Greenstein.

New approach to space diversity combining in microwave digital radio. AT&T Technical Journal, 64(4):885–905, 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Yacobellis:1983:PAD

[YMNW83] R. H. Yacobellis, J. H. Miller, B. G. Niedfeldt, and S. S. Weber. The 3B20D processor & DMERT operating system: Field administration subsystems. The Bell System Technical Journal, 62(1):323–339, January 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-1-323.
pdf;   http://www.alcatel-
lucent.com/bstj/vol62-1983/
articles/bstj62-1-323.pdf.

Yoo:1985:AAN

[Yoo85] S. W. Yoo. An algebraic approach to a nonproduct form network. AT&T Technical Journal, 64(10):2505–2523, 1985. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Youssef:1984:AFT

[You84] M. N. Youssef. On the accuracy of forecasting telephone usage demand. AT&T Bell Laboratories Technical Journal, 63 (5):819–849, 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Yue:1980:FHM

[Yue80a] O. Yue. Frequency-hopping, multiple-access, phase-shift-keying system performance in a Rayleigh fading environment. The Bell System Technical Journal, 59(6):861–879, July/August 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

                                        //bstj.bell-labs.com/BSTJ/
                                        images/Vol59/bstj59-6-861.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol59-1980/
                                        articles/bstj59-6-861.pdf.

Yue:1980:SPA

[Yue80b] O. Yue. Saddle-point approximation for M-ary phaseshift keying with adjacent satellite interference. The Bell System Technical Journal, 59(7):1139–1152, September 1980. CODEN BSTJAN.

ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-7-1139.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-7-1139.pdf.

Yum:1980:MUS

[Yum80]  T. Yum.  Measuring the

utilization of a synchronous data link: An application of busy-period analysis. The Bell System Technical Journal, 59(5):731–744, May/June 1980. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:

//bstj.bell-labs.com/BSTJ/
images/Vol59/bstj59-5-731.
pdf;   http://www.alcatel-
lucent.com/bstj/vol59-1980/
articles/bstj59-5-731.pdf.